A Complete Bibliography of Publications in IEEE Transactions on Software Engineering (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 May 2021 Version 1.01

Title word cross-reference

I [LS94]. N [BKL90]. t [IKWZ90]. x [HP91].

-composition [LS94]. -Kernel [HP91]. -resilient [IKWZ90]. -version [BKL90].

0.1 [LLG98].

1 [BDF96].

3D [Ell99].

80x86 [SPL96].

95 [WB97].

A* [LR95]. abilities [DMW^{+}95]. ability [FW93b]. Abstract [CL95, IW95, SWO97]. abstraction [Bis90, CNR90, GSM^{+}95, HKL^{+}98, SWO97]. Abstractions [SDK^{+}95, CMS96, EHO94, GCB^{+}98, MR98]. ABYSS [WC90]. access [DS90, GS93, HO90, KM99, VMC90]. Accessing [RP93]. Accountability [Kai96]. accuracy [Jor95]. accurate [HR94a, KIR^{+}99, LBJ^{+}95]. Achieving [Hel93, KM90a, TT95, TN97, ACG93]. ACL2 [You97]. acquisition [RW91]. action [KR95, LKBK91]. actions [TC95, WB97, WR96]. activity [QSvMG96]. Ada [Kar90a, MO90, YLT93, AE92, Cor96b, Dil90, GN97, GMP90, Hem90, HW89, KSB89, KB90, KB91, LL97, Mos90, TCO91, WB97]. Ada95 [Che97]. Ada95-like [Che97].

Adaptive [Fer99, GH91, KS94b, FTTV96, HRPL^{+}95]. Addendum [Sto94]. Additivity [BMB97, PDB^{+}97]. adequacy [PZ91, Zhu96]. ADTEST [GN97]. advanced [AB94]. AETG [CDFP97]. after [Ebr97]. against [Cha90a]. agent [HR94b]. agents [DFP98]. aggregation [SMS99]. aid [CER90]. aided [BKWZ94, TMN^{+}99, WS92]. AIN [LLG98]. Albrecht [RLL93]. algebra [PP96, TVK90]. algebraic [BDM96]. algorithm [AL95, AS97, Amm92, BB90, Gra91, HG99, KS91, Lan99, RCGF97, RLP90, SG91, SLL93, Vos91, Wan96, War96, YLN88]. algorithm-based [BB90]. Algorithmic [KSH^{+}92]. Algorithms [FL97, QSvMG96, BR90, FTTV96, FD94, IM93, MK93, MRT91, Pik06, QHK91, RvH93, Rus99, VAA99, VF92, WM97, War96, WS97]. Algres [CCTCR92]. alias [HR96]. all-DU-paths [PZ95]. all-edges [PZ95]. all-uses [PZ95]. Allocating [Kau90, Tow86, Tow90]. allocation [Hou90, HT93, ZA91]. alternating [Suz90]. alternative [Hu97, PKJH99]. alternatives [SW93, WRH98]. ambitious [CCF94]. Amdahl [KH92b]. among [PZ95]. Amulet [MMM^{+}97b]. analogies [SS97]. analogy [MS99]. analyses [Fea98, KL94, Tia95]. Analysis [BKL90, CS92, CBMW93, CDY92, KH92a, PZ91, SMS99, YD92, AEHB96, AR96, AJ93, ABC^{+}91, BM94, BBT92, BR90, BTW95, CS94, CCDD98, CB93, CGK98, CB99, CK94a, CDK98, CCF94, CBS99, CVW98, Cor96b, DZ90, DS97b, DC96, DF98, Ebr97, FW93b, GAF91, HW90, HS93, Han95, HRS98, HKL94, HG99, HL93, HvB94, How90, HW94, IM93, IKP95, IW95, JLHM91, Joh95, KdBMdS91, KB91, KAS93, Kem96, Kem99, LRF95, LCK98, Lev90, LM95, LBJ^{+}95, LS90b, Liu98, LB94, LKA^{+}95, NK91, NvS90, NM94, OO90, PC92, PJ97, RT93, RLP90, RT90, Ser96, SPL96, SR99, SY93, Suz90, SK91, TR91, TWJ92, TYC95, VF92, Ves94, WS92, WK90, WS90, YLT93, YD93, Zhu96, vML99]. analyzers [Jar98, KB90]. Analyzing [ACD98, DIM95, RH96, SB91, SHH91, WJ91, AE92, CNW90, Ebe92, JD96, KP91, Kit98]. animation [Ell99]. Anonymous [JR99]. aperiodic [BC99, RCGF97]. Apollo [Pek92]. Application [HS90, BH98, Fea98, HNUI94, KdBMdS91, LMM96, TMC99, TYOJ91, TYC95]. applications [AAA^{+}90, BG98, Bou94, CCTCR92, CTV^{+}98, GSM^{+}95, Hou90, Kar90b, LOHS98a, LOHS^{+}98b, MVD96, MK92a, RH91, SS91, SMYP99, UGB93, WB98]. Apprentice [RW91]. approach [AEHB96, Ada96, BLS91, BJ91, BG91, BG98, BBT92, CCdC92, CCdC93, Che96, Chu91, CI93, CDFP97, CY90, DZ90, Ell99, GL92, GS90b, HBRY91, IKP95, KIR^{+}99, KS99a, KP91, LKBK91, LC92, LSB94, Liu98, MBK91, MCN92, OB92, PW92, PW93, Puc92, Ros95, SI91, SO94, SSM90]. approaches [SF95]. appropriateness [Sea93]. Approximate [CCJS94, Joh95, MW90, Ser96]. arbitrary [SGL97]. Architectural [GN95, Buh98, TMA^{+}96]. Architecture [RW94, CTV^{+}98, DC95, GD93, HRPL^{+}95, HP91, JZ98, Le 98, LV95, LKA^{+}95, MQR95, RS96, RMSF91, SR92, SDK^{+}95, SPL96, WB98, WC90]. Architecture-directed [RW94]. architecture-independence [GD93]. architectures [DR93, IW95, ORSvH95]. argument [PT98]. Argus [GSW92]. arrivals [CW90]. artificial [HNUI94, SR93, WS92]. ASN.1 [NY90]. ASN.1-C [NY90]. aspect [MWB99]. aspect-oriented [MWB99]. aspects [SD98]. assertions [CBMW93, Ros95, SM94]. assess [MS99, PSTV97, PT98]. Assessing [PJ97, VV93, MWB99]. assessment [ABK94, AW95, BS96]. Assignment [PSA97, LMM96, LS95]. assistance [RW91]. associations [PLR94]. ASTRAL [CPKM94, CPGK97]. Asymptotic [SK91]. asynchronous [CGP99, MK93]. asynchronously [CDS99]. atomic [TC95, WB97]. atomicity [CMSW93]. attacker [JO97]. attacks [Aba97]. attainment [MYK94]. attention [LL97]. attribute [MK93]. auction [FR96]. audit [HL93]. authentication [Aba97, Sch98]. Author [Ano97b, Ano98a]. automata [NSY92, PA91]. Automated [ABC^{+}91, ACDW94, BH98, HS91a, Kor90, TVK90, DM90, PW92, RW91]. Automatic [AHH96, BM94, CGP99, HvB94, LP98a, LB91, MH91, AW95, DO91, Gir93, HG99, KB95, NTFT91, OO90, PSTS91]. Automatically [WGS94, MBK91]. Automating [BKWZ94, CGP^{+}90, GL92]. availability [Gai91, HM95, LKBK91, LK92, TT95]. available [FC99]. Avalon [Win90]. Avalon/C [Win90]. avionics [DS97b]. avoid [TM98]. avoidance [Lan99]. Aware [Ano98b, BG98]. awareness [WB98]. axioms [CS91, Pik06].

B [BR95]. balancing [AG91, Kun91, LR92, Rom91]. bang [RLL93]. banishing [DW90]. banker [Lan99]. base [RS93]. based [AEHB96, AK98a, ABBC99, AG93b, BB90, BKWZ94, BM94, BLY93, BG98, BMB96, BMB97, BMB99, CS94, CCdC92, CCdC93, CRCK98, CT98, Che97, CBMW93, CI93, CDFP97, DO91, EH91, FTTV96, FvBK^{+}91, GHS95, Gir93, HJR91, Har92, HL96, HWR99, IKP95, JO97, KB91, KM97, KS91, LV97, LV95, LDvB94, LvBP94, LS99, Lus94, MT96, MTMS92, MRF^{+}98, MMM97a, Mor90a, NJ91, PDB^{+}97, RGB93, Shi93, SI91, SVK97, SC96, SMMM98, SSM90, Tai96, TMA^{+}96, Tia95, TLP95, TT95, Voa92, WM95, Wan96, WH94, YLN88, YLT93, ZA91, Zus97]. basis [Fen94, TWJ92]. batched [LM90]. Bayesian [BC90, CS94, CBS99]. BDL [BA99]. behavior [GGJ^{+}90, HW90, JO97, KIT93, LS91, SPL98, WP94]. behavioral [CBMW93, SD98]. benchmarking [MS97b]. benefits [MS99, PSTV97, PT98]. Bernoulli [LM95]. between [CY94, LSHZ94, NKF94, SD98, SMS99, Zhu96]. Bidding [Cha90a]. bifunctional [LJS^{+}90]. binary [CM96]. bit [Suz90]. blackboard [VS95]. Block [GS93, Puc92]. blocking [YD93]. blocks [TMT93]. Boolean [WGS94]. Botting [BC96]. bottlenecking [NWPM95]. bound [AM90]. bounded [NM94]. Bounding [SGL97]. bounds [ACDW94, Bur96, CCS91, FM94, WM95, YPV91]. box [Edw97]. Boyer [Gol90]. branch [FW93a, FW93c]. branches [Kau90, Tow86, Tow90]. bridging [FKV91, LSHZ94]. buffer [Dan94, GS93, LMM96]. Building [BSL99]. Burh [YLT93].

C [CNR90, CGK98, NY90, PLR94, WM97, Win90, YD95]. cache [CS92, CRL99, LM95, LMM96]. Caching [KM92]. CAD [BH98]. calibration [Ebr99, GHS95]. Call [FHLS99, Hal95, CCHK90]. Capacity [RJT93, Mut92]. capture [VV93]. capture-recapture [VV93]. capturing [RD92]. case [AL95, BG91, BHT^{+}93, Buh98, FEG92, GGJ^{+}90, GSW92, HvB94, LBJ^{+}95, MH92, NK91, NAT93, SW93, SKN96, TVK90, WRH98, BKWZ94, RLL93]. case-study [MH92]. cases [LCK98]. causal [AM98, LP98b]. cause [NK91]. cause-effect [NK91]. cautious [IKK90]. Cecil [OO90]. central [LR92]. centralized [CDIY90]. Certification [WR94a]. CES [GSW92]. chain [WT94]. chains [Bou94]. change [CNW90, GJB96, KM90b]. changes [BW94, Mad92, MW90]. changing [DM94, SK91]. channel [Bar95]. channels [CKA93, KP91, Sto94, TGC90]. characterization [Bou94, CGL94, HWR99, MRF^{+}98, SS95a, YCHL92]. characterizing [GGJ^{+}90]. check [CGR95]. Checker [BCY01, FG97, Hol97]. checking [AG93b, Bar95, CAB^{+}98, HKL^{+}98, MS97a, PT98, SY93]. checklist [Che96]. checkpointing [GDT92, NvS90, RS93]. checkpoints [DKKP91, HNR99]. checks [LCKS90]. chemical [IW95]. Cheung [Hir90]. Chidamber [HM96]. Cicero [HR98]. circuits [AJ93]. circularity [Muh91]. Clarifying [PZ93]. class [HvB94, McL96]. classification [CBC^{+}92]. Client [KM99, NWPM95]. Client-access [KM99]. client-server [NWPM95]. Clock [Aud97]. clocks [ACG93, Aud97]. Closure [AG93a, QHK91]. closures [Wan96, YLN88]. clustered [GS93]. clustering [KB95]. CNC [Mor90b]. code [CCDD98, CGK98, DM94, Dev95, Edw97, FPV98, GH95, ML95, PP94, PSTV97]. coding [vML99]. Cognitive [SV92]. coherence [CRL99]. coherency [CDY92]. cohesion [BO94, BK98]. collaborative [GSW92]. colored [Gae96]. combinatorial [CDFP97, LW90]. combinators [CD99]. come [ZJ96]. Comment [AS97]. Comments [BCY01, BBC96, BC96, BMB97, CSCK95, Con90, Fal92, Foo93, Gri94, Hir90, How90, Kar90a, Kau90, KPF97, MO90, MBB^{+}97, PKJH99, PDB^{+}97, RGRF01, YSW^{+}01, YLT93]. commerce [Kai96]. commercial [TLP95]. Commitment [MTMS92]. Commitment-based [MTMS92]. common [RS93, KM97]. Communicating [Sha92, BV95, CDS99, EJ90, LvBP94, PSA97]. Communication [DB98, SB98, Aba97, AM90, Bar90, BDV91, HR94b, MIS97, Ran93, WSY91]. communications [HW89, Kar90a]. community [NG90]. compact [DC96]. Comparative [NvS90, PJ97]. Comparing [PVB95, You97, SO94]. comparison [BR95, DR93, FW93a, JLB93, RLL93, SL91, SW93, YS93]. Compartmented [BPWC90]. competing [Bou94]. competitors [Cha90a]. compilation [Eng99]. compiler [Cha90b, LJS^{+}90, NY90]. compilers [SS95a]. Compiling [GH95, NSY92, Eng99]. complete [IKWZ90]. completely [LM90]. Completeness [HL96]. completion [SGL97]. Complex [KM92, Buh98, DK94, HLN^{+}90, SMYP99]. Complexity [TZ95, Amm92, BU91, CS91, GK91, Har92, LJS91, SM94]. compliance [EFM^{+}99]. Component [AK98a, TMA^{+}96, SMS99]. component- [TMA^{+}96]. components [BBH93, FP94, Fre91, HW91, MMM97a, WR94a]. Composite [Muh91, FH92]. Composition [BG97, JZ98, LS94, McL96]. Compositional [BV95, FG97, KLB93, MR98]. compositions [HT96]. Compound [Sah92]. Compound-Poisson [Sah92]. Comprehending [ADS99]. comprehension [VV96, vML99]. Computation [Kor97, CCJS94, HR96]. Computational [GQ96, Kar90b, LW90, MS92, WHH^{+}92]. compute [Ves94]. Computer [WS92, BKWZ94, BM94, HL93, LS90c, Tai96, TMN^{+}99, UGB93]. computer-aided [BKWZ94, TMN^{+}99]. computers [CMS98]. Computing [Ano98b, FM94, AT93, ACG93, AG93a, CD99, JR99, KSH^{+}92, MR98, Pek92, SS95b, WB98, WW91]. concept [CBC^{+}92, CL95, DJC94, KNE92, SR99]. concepts [PZ93]. concurrency [CDY92, HD91, IKK90, RC90, YD92, YD93]. concurrency-coherency [CDY92]. concurrent [ABC^{+}91, ACDW94, BDM96, CT98, CB93, Che98, Cor96a, DA91, GPS96, Gol90, HGR92, Nes93, NM94, Sta93, TCO91, Tay92, TLK92, TC95, Woo95]. concurrent-software [NM94]. conditional [SY93]. conditions [MT96]. confidence [HT90]. confidentiality [Jac91]. configurable [GSM^{+}95]. conflict [SMS99]. Conflicts [LS99, vLDL98]. conformance [PF90]. conjunctions [HMRS98]. connectionist [KWM92]. Connectors [WF98]. conservative [LW97]. considerations [KL94]. Consistency [HNR99, Fea98, HL96, KK99, MS97a, TN97, You97]. consistent [CCS91]. constrained [ABC^{+}91]. Constraint [DO91, Gir93, CBMW93, GH95, GS99, OO90, TYC95]. Constraint-based [DO91, Gir93]. constraint-oriented [GS99]. constraints [CT98, Hua90, XP93a, XP93b]. Constructing [CCHK90, MBK91]. construction [Muh91, NTFT91]. Constructive [HR98]. containing [Kau90, Tow86, Tow90]. contention [YD92]. contentions [CSL91]. context [SPL98]. contingency [Chu91]. continuous [DL91, Foo93]. contribution [Di 93]. Control [Yad90, AHSR93, AG94a, Amm92, BM94, BA99, CB94, CDY92, Dan94, DS97b, EH91, HO90, HD91, HG99, IKK90, Jac95, JLHM91, LJS91, LHHR94, MK92a, PW93, YD92, YD93]. control-flow [Amm92, LJS91]. control-intensive [AG94a]. Controlled [LRLH91, MS99, PT98]. controllers [WK90]. convergence [AG93a]. Conversion [Nov95a]. cooperation [BDF96, SC98]. cooperative [DT94, PS93]. Coordinating [CTV^{+}98]. copies [AT93, HH97]. Copying [HW91]. Correct [MQR95, BC90]. Correction [CCdC93, Das93, LOHS98a, Tow90, XP93a, ACG93]. correlated [NG90]. correlation [TMT93]. Correspondence [BC96, Ebe92]. Corrigenda [HS97]. Corrigendum [Gir93]. Cost [CVW98, Bin97, CBS99, Ebr99, Gut95, HZO98, LP98b, MBM94, PN97, PSTV97, RW97, Wey90, ZA91, ZEWH95]. cost-benefits [PSTV97]. Cost-effective [CVW98]. cost-effectiveness [RW97]. costs [AM90]. count [CCS91]. counterexample [JD96]. counting [JLB93]. coupling [BDW99]. coverage [RW97]. covers [BM94]. Covert [KP91, TGC90]. Creation [Nov95b]. criteria [Con90, Hof90, PZ95, RT93, Zhu96]. Critical [KSB89, MO90, BV95, BR90, BF93, DMW^{+}95, GMMP91, GDT92, Jac95, KB90, LW97, RvH93]. critique [FN99]. cross [HR94b]. cross-RPC [HR94b]. cryptographic [AN96]. CSCW [SC98]. CSDL [DT94]. CSP [Cam90, KR93, LR97, Moo90, Sch98, SNH93, TC95]. CSPL [Che97]. cut [DW90]. Cyclomatic [GK91].

D [Con90, Kau90, XP93a]. Data

[BB93, CG92, GS90a, Mos90, BDS93, Bis90, BBT92, BR90, CCTCR92, CS94, CB93, CGK98, CB99, CNW90, CL95, DO91, DM90, EHO94, FW93a, GN97, Gir93, GS93, GCB^{+}98, HLR92, HvB94, Kor90, LRF95, Lev90, MCB90, Mil90, Nov95b, PZ91, RT93, RLP90, Sch93, SWO97, Smi91, Sta95, War96, WH94, Wey90, Wey93, WGS94, YD92]. Data-driven [GS90a]. data-flow [HLR92]. data-intensive [CCTCR92]. Data-oriented [CG92]. database [CI93, DCN93, Fal92, GD93, IKK90, KH92a, LMM96, Mul90, NJ91, SS91, Sin90, Wan96, WMS^{+}92, YLN88, YCHL92]. databases [CY90, HD91, PP96, WSY91]. dataflow [CK94a, Fra92, KSH^{+}92, RH91]. datasets [Kit98]. dataspace [RC90]. DBMS [DR93]. dead [CGK98]. deadlines [CW90, FL97, SLL93, XP90, Xu93]. deadlock [Cor96a, KB91, KS91, Lan99, YLT93]. deadlocks [Bar90, KS94c, Muh91]. Debugging [GGH90, GBB90, TCO91, KB95, PC90, Sin91, TFCB90, VD95]. Dec [CCdC93]. decay [Hel93]. declarative [vLW98]. decomposition [GS99, Hua90, Moo90, PW92]. decompositional [LSB94]. deductive [MW92]. def [PLR94]. def-use [PLR94]. defect [CBC^{+}92, FN99, Lev90]. defects [AE92, KK99]. deferred [CB99, SLL93, YD93]. Defining [BMB99, DL91, Foo93]. definite [Sta95]. definition [CMBC93, LV95, Yad90]. delay [vG91]. deliberations [RD92]. delivered [FHLS98]. demand [ML95]. DeMarco [RLL93]. DeMillo [Gir93]. denial [YG90]. density [GK91]. Dependability [LB94, AAA^{+}90, BS96, CMS98, KIR^{+}99, KM90a, LI95, MS97b]. dependences [PC90]. dependencies [DS90]. dependency [BDV91, Mos90, SD98]. dependent [BD91, HDL90]. Derivation [War96, ACDW94, HvB94, MH91, PF90, RGB93, Sta93, WS92]. Describing [Le 98]. descriptions [Kun91, MH91]. Design [HPR^{+}97, Jar98, SVK97, WEHL94, AK98a, BLS91, BS91a, BL94, BB90, BBM96, BK98, BMB99, CGZ91, CRCK98, CK94b, CNW90, CSCK95, CDFP97, CHB92, CY90, CL95, DT94, DF98, Ebr97, EH91, FH92, FGNR92, FR96, GCG90, Gri94, HW91, HNUI94, JD96, KR93, KH92b, KB95, KS93, LSB94, LB91, MDC92, NY90, ORSvH95, Sat92, Sen97, SW93, SO94, SO98, SC98, TMC99, TC95, WRH98, Yad90]. design-level [BK98]. designers [FGNR92]. Designing [AK98b, CCTCR92, HR94b]. designs [AE92, VV93]. detect [HKL^{+}98, LR97, TM98]. detected [CY96]. Detecting [LW92, DMW^{+}95, FW93b, GS90b]. detection [ABK94, BB90, CGK98, Cor96a, HL93, HMRS98, IKP95, KS91, KS94c, LCKS90, LLG98, MK92b, PVB95, PZC^{+}96, RT93, WH94]. detector [JD96]. Determining [Sin91]. Deterministic [AJ93, SB99, TCO91]. developers [BSV96]. Developing [BBH93, PJ91]. development [BHT^{+}93, BSV96, BG98, CRCK98, DB98, GHM98, HLN^{+}90, KM90a, MTMS92, MBM94, MVD96, MA93, MWB99, MMM^{+}97b, NAT93, PK91, PN97, PSTV97, RD92, RF92, RGFR98, RGRF01, RP99, Smi90, SF95, SMYP99, TBH91, VF92, YSW^{+}01, ZEWH95, vG91]. deviations [Cug98]. device [TMC99]. diagnosis [WLS97]. diagrams [BBC96, BC95, DM90, Fra92, PSTS91]. dictionary [BDS93]. different [Kun91, NvS90, Nov95b]. difficult [Fin96]. Dijkstra [NBK95]. dimensional [DCN93, WMS^{+}92]. directed [GKNV93, RW94]. directions [MMM95]. disciplined [DZ90]. Discrete [CNT99, BL94, De 90, Gae96, WK90]. Discrete-event [CNT99, BL94, Gae96]. discussion [SB98]. disk [CS92]. dispatcher [LR92]. Distributed [AB94, CMS96, JZ98, AT93, AG91, ACG93, BS91a, BKT92, Bar90, CK94a, CSL91, CDIY90, CDY92, De 90, DJC94, GSM^{+}95, GSW92, HW90, HD91, HM95, HNR99, HPR^{+}97, Hou90, HBRY91, HMRS98, IKWZ90, KM90a, KS91, KS94c, KB95, LS90b, LB94, LS99, MYK94, Min91, Mul90, PC92, PS93, PSA97, RS93, Sat92, SGCT90, SS95b, SMYP99, TT95, Tri96, TN97, VD95, WHH^{+}92, WW91]. distribution [CGP99, TYOJ91, Vos91]. distributions [Gut95]. do [ZJ96]. document [Ebr97]. documentation [NAT93, PMI94, PP98]. does [HT90]. Domain [TMC99, ABBC99, CRL99, FGNR92, HRPL^{+}95, KS99b, SM98, Tia95, Pek92]. domain-oriented [FGNR92]. Domain-specific [TMC99, ABBC99, CRL99, HRPL^{+}95, KS99b]. drawing [BDL95, DM90, GKNV93]. driven [AG93b, GS90a, vLDL98]. drivers [TMC99]. DSA [CMS96]. DSP [RH91]. DSPNs [HG99]. during [BHT^{+}93, RD92, VV96]. Dynamic [SZ92, WM97, Kor97, KM90b, LR92, RCGF97, RT90, Voa92, VV96]. dynamic-priority [RCGF97]. dynamically [SVK97].

Early [RWG90, MK92a]. economic [CCH96]. economy [WHH^{+}92]. edges [PZ95]. editing [GSW92]. Editor [Raj92]. Editorial [Ano97c, CS99, CG96, Gar98, Jaz98, LJ98, Mea96, NJ96, Son97, GN99, HM99, RW99]. editors [Shi93]. effect [ABK94, Bis90, NK91]. Effective [BTW95, CVW98, CY90, MMM^{+}97b]. effectiveness [FW93a, RW97]. effects [DH92, MW90, WS90, Zha91, ZEWH95]. efficiency [LW90, TT95]. Efficient [Gae96, HL91, HMRS98, KS94c, LM90, MRT91, QHK91, BL94, BB90, CCF94, SB99, WW91]. effort [GBB90, Jor95, MK92a, PN97, SS97, SF95]. election [FC99, IKWZ90]. electronic [AB96, Kai96]. Elements [JD96]. elicitation [HPW98]. elimination [LRF95, SL91]. Embedded [HT91, AHH96, Ell99]. emerging [MWB99]. Empirical [JV99, RH98, AR96, ADS99, CBS99, FP94, KP92, KS99a, LL97, LCKS90, MA93, OL94, SB98, Sea99, SL91, SV92, SR93, TMN^{+}99, Wey90, vG91]. enabling [PC92]. enactment [Cug98]. Encapsulation [GD93, ZEWH95]. end [FL97, TBH91]. end-to-end [FL97]. Engineering [BMB97, JV99, KAS93, SO98, AN96, BKWZ94, BBT92, BMB96, BTW95, CCdC92, CCdC93, CCF94, CBS99, EDRM90, FN98, LCK98, LOHS98a, LOHS^{+}98b, OB92, Par93, PDB^{+}97, RD92, RMSF91, Sea99, SW93, SM98, SMMM98, TMN^{+}99, Woo95, Zus97, vLDL98, Ano95]. enhance [BCD97, CL95]. entities [Buh98]. entity [BDS93, MM90]. entity-relationship [MM90]. entropy [Har92]. entropy-based [Har92]. enumerative [PS93]. Environment [Mad92, BFG93, BDF96, BKWZ94, BLW90, BC99, CCTCR92, CDM95, Che97, DCN93, HJR91, HLN^{+}90, HM93, KIT93, MS91, MS92, MA93, Mut92, MMM^{+}97b, PJ91, RH91, RUPT90, RLL93, SMYP99, SK91, TMN^{+}99, VF92, WB98, YD95, YD92]. environments [CP91, FTTV96, FGNR92, GHM98, HBRY91, PK91, YCHL92]. EPOS [JC93]. EQL [WM95]. Ergodicity [CCS91]. error [ABK94, BB90, FL97, LP98b, LCKS90, LW92, NK91, NG90, SB91, SLvH92]. error-prone [SB91]. errors [Ebr97, How90, LR97]. ESML [RM93]. ESML-extended [RM93]. Estella [CBMW93]. estimate [CCH96]. estimates [TBH91]. Estimating [MMN^{+}92, Mut92, SD90, SS97, Chu91, LP98b, ML95, MS99, SF95]. Estimation [Haa99, Ada96, BC90, GBB90, GS93, Gut95, HS97, Lar90, LJ90, MBM94, MK92a, PN97, TYOJ91, TDN93]. Estimeetings [TBH91]. European [MVD96]. EVA [MKH97]. evaluate [Zha91]. Evaluating [CNW90, Cor96a, FHLS98, Hu97, MWB99, PKJH99, SKN96, ACJ^{+}96, Ebe92, MA93, Sch99, Sea93, WRH98]. evaluation [BKZ93, BS91b, BLW90, CGZ91, CMS98, CHBC93, ECK^{+}91, GDG92, HL91, HCN98, KIR^{+}99, LKBK91, LP98a, LJ90, MYK94, MK93, OL94, ODK99, SG90, SC98, TZ95, YZM97, Zav91, ZQ91]. event [AG93b, BL94, CNT99, De 90, Gae96, KR95, LV95]. event-action [KR95]. event-based [LV95]. event-driven [AG93b]. Evolution [RS96, SPL98, BFG93, JC93, KS99a, LX93, Luq90, Mad92, RUPT90]. evolving [KM90b, MKH97]. Exact [LM95]. examination [MvMS93]. example [BW94, Sta93]. Exception [Gov93, CG92]. exchange [KR93]. exclusion [XP90, Xu93]. executable [DFPT90, Hem90, LS91, Nes93, NP92]. Executing [RH91]. execution [CP91, CS94, CPDGM91, DH92, GD93, HS90, SB99, SO98, SGL97, TCO91, TLP95, YPV91, YMST99]. Existence [SD98, MT96]. expected [CY96, DIM95]. Experience [Jor95, MRF^{+}98, SO94, ACJ^{+}96, CMSW93, RWG90, Wey93]. Experiences [ELC^{+}98, Das92, Das93]. experiment [BPV94, BKL90, KM90a, MS99, PVB95, PSTV97, PT98]. Experimental [BS91b, BLW90, AHSR93, CMS98, ECK^{+}91, FW93a, PJ97, YS93]. Experimentally [GGJ^{+}90]. Experimenting [ODK99]. experiments [BSL99]. expert [CS94]. Explicit [Aba97]. exploratory [CDK98]. exporting [WB98]. exposure [MvMS93]. expressing [NKF94]. expression [ABC^{+}91]. expressions [BG91, HvB94, LM90, Lus94]. expressiveness [BCD97]. extended [Fra92, Lan99, MM90, NSY92, PMdO99, RM93]. Extending [HO90, SR98, SY93]. extensible [CCDD98, GD93]. extensions [DS98]. Extracting [LV97].

Factors [RGFR98, YSW^{+}01, RGRF01, YS93]. failure [BC90, CS94, FF96, Gut95, ML95, MMN^{+}92, Sch93, Voa92]. failure-based [Voa92]. failures [CY96, MMN^{+}92, NG90]. fallible [AHSR93]. families [BSL99]. fast [Ran93]. Fault [AAA^{+}90, AG93a, AK98b, BG91, DMW^{+}95, FW93b, HT93, KIR^{+}99, KIT93, MvMS93, Mor90a, MK92b, OA96, ORSvH95, Pik06, PJ91, RT93, Rus99, SGCT90, SL91, Tai96, TC95, WH94, WW91, WS90]. fault-based [Mor90a, Tai96]. fault-detecting [DMW^{+}95, FW93b]. fault-detection [WH94]. fault-prone [MK92b, OA96]. fault-tolerance [AK98b, SGCT90]. fault-tolerant [AG93a, BG91, HT93, ORSvH95, Pik06, PJ91, Rus99, WW91]. faults [BKL90, KIT93]. Feature [LJ98, JD96, JZ98, KK98, LLG98, TM98]. FieldBus [CDM95]. file [Gai91, RP93, Sat92, TN97]. files [LM90, RP93]. filter [Gra91]. fine [RH91, KIT93]. fine-grained [RH91]. finite [FvBK^{+}91, GS93, LvBP94]. finite-state [LvBP94]. firing [CCS91]. firm [BC99]. fit [SV92]. Fixed [Ves94, BTW95, HKL94, KAS93, MT96]. Fixed-priority [Ves94, HKL94]. flat [DA91]. flexible [Cug98, HYT91, Jar98, MKH97]. Flow [Dan94, Amm92, BM94, BR90, EG90, FG97, FW93a, HLR92, KP91, LJS91, LV97, LRF95, RLP90, Wey90, Wey93]. fluid [CNT99]. flush [CKA93, Sto94]. fly [DS98]. foolproofing [NAT93]. form [ABBC99, Edw97, FN91]. form-based [ABBC99]. Formal [Che98, CGR95, DS97b, GCG90, GMP90, IW95, ND90, ORSvH95, RGB93, RvH93, Sta93, Suz90, BLS91, BBC96, BC95, BW94, Che96, CPKM94, DJC94, ELC^{+}98, Fea98, Fin96, Fra92, FW93b, FKV91, GS99, GPS96, GJB96, Hem90, LOHS98a, LOHS^{+}98b, MH91, Pik06, PC90, RS94, Rus99, War96, Zhu96]. formalism [DFPT90, GMMP91]. Formally [WLS97]. formed [FM94]. foundation [AG93a]. foundations [HL93]. fourth [HG99]. fourth-order [HG99]. Fragmenting [SI91]. frame [BMR95]. frames [BR95]. Framework [KPF97, ACJ^{+}96, BG98, BDW99, CPKM94, Dev95, DM90, GJB96, HGR92, KPF95, LLG98, MYK94, MBB^{+}97, NKF94, PP94, Smi92, SC96, VS95]. franca [Tay92]. front [TBH91]. front-end [TBH91]. Function [AR96, LJ90, JLB93, KP92, MBM94, RLL93]. Functional [Mil90, BO94, Gov93, MCB90]. functions [DL91, Foo93, Hu97, LV97, PC92, PKJH99]. fundamental [PZ93]. Fundamentals [MW92].

G [YLT93]. galaxy [BB91]. games [DIM95]. gap [FKV91, LSHZ94]. general [Dil90, HGR92, KR95, McL96, QSvMG96].

Generalization [FN91]. Generalized [CMBC93, GDG92, KH92b, KS94c, LvBP94]. generated [PP98]. generating [BU91, HWR99, Vos91, WGS94]. Generation [SLvH92, AW95, BM94, CGP^{+}90, DO91, GN97, Gir93, HKR90, Kor90, OO90, PSTS91, QSvMG96, TMC99, TVK90]. generators [BG97]. generic [Nov97, WM97]. GenVoca [BG97]. Ginger2 [TMN^{+}99]. given [Vos91]. global [BSV96, KS94b, VD95]. goal [OB92, RP99, RSA98, vLDL98]. goal-driven [vLDL98]. good [Bur96]. grained [RH91]. grammar [WW91]. Grammars [MC92, Hem90, Le 98]. graph [BDL95, EG90, LV97, Le 98, Luq90, NM94, War96]. graphical [ACS90, MS97a, WMS^{+}92]. graphs [AM90, CCJS94, GKNV93, Mos90, Ser96]. Gries [NBK95]. group [Rei96]. Growth [Ano96b, LKBK91, TYOJ91]. GSPNs [Buc99, Kem96, Kem99]. guarantee [KR93]. guaranteed [SHH91]. Guaranteeing [Bur96, GHS95]. guarantees [MIS97]. GUARDIAN [LI95]. Guest [Ano97c, CS99, CG96, Gar98, GN99, HM99, Jaz98, LJ98, Mea96, NJ96, RW99, Son97, Raj92]. GUI [TMA^{+}96]. guide [SO98]. guided [Bin97]. Guiding [RSA98].

handlers [Gov93]. handling [CG92, FGH^{+}94, SLvH92]. hard [AS97, HKL94, SZ92, SG91, SHH91, XP93a, XP93b]. hard-real-time [SHH91, XP93a, XP93b]. Hashing [LRLH91]. Heap [Gra91]. Heap-filter [Gra91]. heterogeneous [CP91, HBRY91, SK91]. heuristic [Kun91]. HI [HTI90]. HI-VISUAL [HTI90]. Hierarchical [Buc99, HM95, KIR^{+}99, NM94, HL96, Muh91, SGCT90]. hierarchically [CB94]. hierarchy [DMW^{+}95]. High [KMR91, BBH93, BMB99, GMMP91, LC92, RMSF91, YD92]. high-level [BMB99, GMMP91]. high-risk [BBH93]. higher [ACG93, Cam90]. highlights [BPWC90]. highly [FC99]. history [NK91, NAT93]. Hoffman [Con90]. homogeneous [GAF91, Rom91]. hookup [McC90]. hopCP [AG94a]. horizontally [SI91]. human [YS93]. hybrid [CDIY90, HW90, TWJ92, YD92]. Hypercharts [PMdO99]. hypercube [BB90]. hypergeometric [TYOJ91]. hypermedia [PMdO99]. hypertext [Ost95].

I/O [GSM^{+}95]. IC [AG94a]. iconic [DCN93, HTI90]. Identification [VV96, TGC90]. Identifying [MM90, SR99, BBH93, TC95]. Identity [KM92]. IEEE [Ano95, Ano96a, Ano97a]. II [LJS^{+}90]. image [DCN93]. impact [RGFR98, RGRF01, YSW^{+}01]. imperfect [ABK94]. Implementation [KM92, AL95, CDOT97, Das92, Das93, FR96, NY90, RGRF01, TMC99]. implementations [SPL96]. implemented [ACD98, GSW92]. Implementing [CMSW93, SG90, WB97, HP91, LR95, RGFR98, YSW^{+}01]. implications [CMBC93, DC96, KK99, PC90]. imposition [Min91]. improve [Ebr99]. improved [GS99, Wan96, Yad90, YLN88]. improvement [BHT^{+}93, Che96, FF96, MRF^{+}98, WR94b]. improvements [FW93c, Kar90b]. Improving [BSV96, KP92, BFL^{+}95, ECK^{+}91, FN98]. in-place [CVW98]. In-process [CHBC93, CBC^{+}92]. including [SW93]. Inconsistency [FGH^{+}94, GHM98, RP99, SSV99]. inconsistent [Pik06]. Incremental [BB91, HKR90, Shi93, BR90, MK93, RLP90]. independence [Bou94, GD93]. Independent [Tri96]. Index [Ano95, Ano96a, Ano97a, Ano97b, Ano98a]. indexes [BP90]. indicators [BBM96]. indices [Fal92, FM94]. individual [HS91b]. induction [LSY94]. industrial [BFL^{+}95, CGR95, GPS96, KM97, LSHZ94, LOHS98a, LOHS^{+}98b, MVD96]. industry [Sta93]. infeasibility [BF93]. inference [FTTV96]. Inferring [vLW98]. influence [Gut99, Kun91, Sat92]. influences [HW91]. Informal [FKV91, HW94, MH91]. Information [HBRY91, AHSR93, AB94, BLS91, Cha90b, CNR90, FG97, HJR91, HR96, Lus94, MBK91, PW92, RW97, RMSF91, SL92, WW90]. inheritance [Edw97]. injection [AAA^{+}90, KIT93]. inlining [DH92]. input [FL97, Tia95]. inputs [Gra91]. insider [Zav91]. inspection [CHBC93, Che96, Ebr97, SB98]. inspections [PVB95, PSTV97]. inspire [HT90]. instantiated [QHK91]. instrumentation [WRH98]. integrated [BS91a, FGNR92]. Integrating [Tia95, VS95]. Integration [DFKM98, LCK98, SR93]. integrity [SD98]. Intel [SPL96]. intelligence [HNUI94]. intelligent [HRPL^{+}95, NBK95, TM98]. intensive [AG94a, CCTCR92, War96]. interaction [DFP98, KK98, LLG98]. Interactions [LJ98, TM98]. interactive [Ell99, SLvH92, VAA99, You97]. interconnected [JR99]. interconnection [RS94]. Interface [Eng99, CER90, CL95, MMM^{+}97b, NAT93, Sea93, SMS99]. interfaces [Con90, Eng99, HO90, Hof90, LS94, MS97a, NHWS92]. Internet [RP93]. interoperation [GQ96]. interpretable [BBH93]. interpretation [RM93]. Interprocedural [PLR94, HM93]. Interprocess [BDV91]. interval [Sin91]. intralevel [CPKM94]. Introducing [CHB92, DK94]. Introduction [Ano97c, CS99, CG96, DS97a, FH97, FT98, JKS98, Jaz98, McH96, Mea96, NJ96, RG98, Son97, GN99, HM99, Raj92, RW99, Gar98]. intrusion [IKP95, JO97, PZC^{+}96]. Invariant [KS91]. Invariant-based [KS91]. Invariants [BR95]. investigation [AHSR93, Wan96]. Issue [Ano98b, DS97a, JKS98, Raj92, RG98]. issues [BG91, GQ96, HNR99, MMM95]. iteration [SV92, Sta95]. iterative [BR90, MT96]. iterators [Lam90, WEHL94].

J [Foo93, Gir93, XP93a, YLT93]. Jade [RP93]. job [LR92]. jobs [SLL93, SGL97]. join [Gra91]. joining [Gra91]. judgment [AHSR93].

Karam [YLT93]. KAT [LKBK91].

Kemerer [HM96]. kernel

[DFP98, KZB^{+}91, HP91]. key [SD98]. KIDS [Smi90]. KLAIM [DFP98]. Knowledge [FH92, SR92, AEHB96, BSL99, EDRM90, HJR91, JFH92, LKBK91, RF92, SI91, TWJ92]. knowledge-based [AEHB96, HJR91, SI91].

L [XP93a]. L.0 [Nes93]. Language [Gri94, MDC92, ACS90, ABBC99, BL94, BKT92, BA99, BCD97, BdS96, CRL99, Cha90b, CGP^{+}90, DFP98, DT94, Ell99, GGH90, HLR92, HC90, JM94, KS94a, KS99b, KLB93, LR95, LS91, LJS^{+}90, LV95, MCB90, NS91, Nes93, OO90, RM93, Shi93, SSM90, WMS^{+}92]. language-based [Shi93]. languages [Eng99, FKV91, Gov93, RS94, TMC99, TVC94]. Larch [Che98, GGH90, Win90]. large [CAB^{+}98, DM94, DK94, GCB^{+}98, Lev90, MRF^{+}98, PSTV97, TBH91, TLP95, Tri96, VV96]. large-scale [Tri96]. large-system [MRF^{+}98]. late [vG91]. latency [WS90, YZM97]. law [KH92b]. layered [GL92]. layering [ZEWH95]. Layers [RS95, RF92]. Layout [Sea93]. LBC [LR92]. leader [FC99]. learned [MWB99]. Learning [TM98, KSH^{+}92, SF95]. least [LM95]. length [CDM95]. Lessons [NOTS98, MWB99]. level [BK98, BMB99, CMBC93, GMMP91, LC92, PLR94, QSvMG96, Ran93]. libraries [MBK91]. library [NTFT91]. life [ABK94, BF93]. life-critical [BF93]. lightweight [ELC^{+}98, Fea98]. like [Che97, NS91]. limited [Dan94]. line [DJC94, ESS92, GJB96, WRH98, WLS97]. Linear [SM94, HL91, Liu98, Ves94, Vos91]. lingua [Tay92]. Lipton [Foo93]. Lisp [KM97]. LISPACK [IM93]. list [Ano99]. Literate [Ost95]. liveness [SY93]. LL [Shi93]. Load [LRLH91, AG91, AW95, BDV91, HWR99, Kun91, LR92, Rom91, SS95b]. load-balancing [LR92]. local [FC99, HMRS98]. locality [CDY92]. location [KS94b, TT95]. location-based [TT95]. locking [EG90, HD91, RT90, TR91, YD93]. logarithmic [CS94]. logging [AM98]. logic [AL95, Cam90, CCdC92, CCdC93, Han95, KB91, KM97, Nes93, Par93, SR98, Tay92, YLT93]. logic-based [CCdC92, CCdC93, KB91, YLT93]. logic-synthesis [AL95]. logical [Aud97, FMM94]. Logres [CCTCR92]. look [ND90]. Loop [SG95, Bis90]. loops [AEHB96, Kau90, Tow86, Tow90]. LOTOS [HvB94, PF90]. Lower [AM90]. LRU [GS93]. luck [DIM95]. lumpable [FM94]. LUSTRE [HLR92].

M [YLT93]. Machine [SF95, IW95, KSH^{+}92]. machines [KMR91, LvBP94, Sha92]. maintaining [KH92a, LMR92]. Maintenance [WH92, CB93, CER90, GL91, GK91, HM93, Jor95, PC90, Raj92, Sch99, VV96]. Maisie [BL94]. Making [BW94, DF98]. management [BB93, CGZ91, GHM98, HBRY91, KM90b, LMM96, LS99, WF93]. manager [ND90]. Managerial [CDK98]. Managing [EFM^{+}99, LJ98, RP99, SSV99, WSY91, vLDL98]. manipulation [Gri94, MDC92, NJ91]. manufacturing [HYT91]. MAP [Mor90b]. maps [Buh98]. marked [CCJS94, Ser96]. marking [War96]. Markov [Bou94, CM96, SNH93, WT94]. masking [AK98b]. masses [Fin96]. massively [AG91]. Mathematical [Fin96, Di 93]. Mawl [ABBC99]. maximally [Qin93]. mean [Ser96]. meaning [GN95]. meaning-preserving [GN95]. means [HLR92]. measure [Har92, TZ95]. Measurement [BMB97, KPF97, WS90, AR96, BMB96, BDW99, Das92, Das93, Fen94, HM96, KP92, KPF95, MBB^{+}97, Nov95a, PDB^{+}97, TLP95, WH94, YZM97, Zha91, Zus97]. Measurements [HW89, Kar90a, CBC^{+}92]. Measures [SS95b, BS96, BMB99, CS91, Ebe92, KMBR92, LJS91, MRF^{+}98]. Measuring [BO94, BK98, KK99, MS97b, Sch99, LC92]. mechanical [BH98]. Mechanically [Gol90]. mechanism [TFCB90]. Mechanizing [Cam90]. mediator [SKN96]. medium [Gra91, VMC90]. medium-size [Gra91]. Meet [Qin93]. meetings [PJ97, SB98]. membership [Rei96]. Memory [DS90, Bur96, GGJ^{+}90, GW92, KH92a, Zha91]. merge [Gra91]. Message [AM98, TMA^{+}96]. message-based [TMA^{+}96]. messages [De 90]. meta [NJ91]. meta-rules [NJ91]. metadatabase [HBRY91]. metaphors [RC90]. metaprogramming [LZ95]. Method [RS95, BC90, Che86, De 90, GS99, Hir90, HW94, KdBMdS91, KR93, LRF95, LvBP94, MKH97, PS93, SKN96, YG90]. methodologies [SO94]. Methodology [Sch92, AAA^{+}90, AB94, CDOT97, EH91, IM93, LLG98, LOHS98a, LOHS^{+}98b, MS91, NBK95, PZC^{+}96, RGFR98, RGRF01, YSW^{+}01]. methods [ACJ^{+}96, Cor96a, CGR95, DMW^{+}95, ELC^{+}98, Fea98, FP94, FW93b, FHLS98, GPS96, Haa99, PVB95, Sea99, TC95, VV93]. metric [Sea93]. metrics [BKWZ94, BBM96, CK94b, CDK98, CSCK95, DB98, GBB90, HCN98, HM96, RLL93, Sch92, Sch99]. Microsoft [SMS99]. MIDAS [BS91a]. military [MVD96]. Millo [Foo93]. mine [MH92]. Minimizing [BC99]. Miro [HMT^{+}90]. Misra [De 90]. misuse [HL93]. Mixed [RC90]. mobile [MR98, WB98, WF98]. Mobility [Ano98b, DFP98, FPV98, WR96]. mode [BPWC90]. Modechart [JM94]. Model [Ano96b, BCY01, Bar95, CAB^{+}98, AJ93, AG93b, BFG93, Bos95, BBC96, BC95, CS94, CCH96, CM96, CGK98, Dro95, FF96, HT96, HKL^{+}98, Hol97, IW95, JC93, JO97, KH92a, LP98b, LS90c, LDS^{+}90, Luq90, Mad92, ML95, MC97, Nic90, PN97, PC90, RS94, RT93, RC90, Sah92, Sch93, TYOJ91, VT92, WW90, WT94, Woo95, SMS99]. model-oriented [Nic90]. Modeling [BFL^{+}95, BD91, CB93, HYT91, NG90, SGCT90, Smi91, TMT93, WRH98, DA91, ELC^{+}98, Ell99, HM95, KIR^{+}99, KMBR92, LKBK91, LK92, LS91, Lev90, LB94, PA91, Puc92, RM93, RSA98, SSM90, TLP95, Zha91]. Models [PK91, ADS99, BA93, BBH93, BCLS90, CBS99, Cug98, Ebr99, FMM94, FN99, FvBK^{+}91, HGR92, HZO98, HWR99, Jor95, KWM92, LB91, MT96, MRT91, MK92a, MMM^{+}97b, MS99, NvS90, OB92, Sar95, SNH93, Tia95, VAA99, Ves94]. modern [CMS98, DR93]. modes [FF96]. Modified [SLL93]. Modular [BDM96, EHO94, BA93, LCK98]. modularization [Yad90]. Module [Hou90, Con90, Hof90, HS91a, RS94]. modules [CDS99, LS94, NTFT91, OA96, SR99, WP94]. monitor [HW90]. monitoring [BPV94, HS90, KIT93, LC92, ODK99, TFCB90, WSY91]. monitors [RP99]. monotonic [SLL93, SG95]. MOOD [HCN98]. Moore [Gol90]. motion [GH95]. Motorola [Das93, Das92]. multiagent [CTV^{+}98]. multicast [Dan94]. multicomputer [AG91]. multiframe [MC97]. multigraph [CCHK90]. multilevel [CB99, LS90c, McC90]. multimedia [Ell99]. multiparadigm [ZJ96]. multiperspective [FGH^{+}94]. multiple [GHM98, GAF91, HO90, HL91, Kau90, NKF94, Tow86, Tow90, VS95]. multiple-view [GHM98]. Multiprocessor [Xu93, BS91b, GGJ^{+}90, GST91, LS95, SK91, YZM97, Zha91, ZQ91]. multiprocessors [BB90, CMS96, DS90, GW92]. multitolerant [AK98a]. Multiversion [IKK90, NG90]. mural [FEG92]. mutation [OL94].

navigation [UGB93]. necessary [Fen94]. need [SWO97]. needed [Sar95]. Negotiated [NHWS92]. negotiation [SMS99]. nested [Aud97]. net [CMBC93, CGL94, DC96, FMM94, GMMP91, MT96, Smi92, WK90]. net-based [MT96]. nets [BDM96, BG91, BD91, Bou94, BV95, CCS91, CMBC93, CNT99, DS98, Fer99, FN91, FM94, Gae96, GDG92, Haa99, HYT91, HT91, LCK98, LSY94, Liu98, NM94, PC92, Suz90, SSM90, TYC95]. Network [Ano98b, BG98, BDV91, FN91, Gai91, HGR92, HP91, KML96, Pek92, PA91, Rom91, SL92]. Network-Aware [Ano98b]. networks [CGP99, IKWZ90, LS90c, NWPM95, QSvMG96, TM98, WSY91]. neural [HGR92, KSH^{+}92, MS92, Smi92, SL92, WS92]. No [Sar95, MMN^{+}92, RT90]. no-waiting [RT90]. nondeterministic [LvBP94]. nonfunctional [MCN92]. noninterference [TFCB90]. nonlinear [TVC94]. nonmasking [AK98b]. nonrandomness [Fel90]. nonregenerative [Haa99]. nontraditional [SS91]. normalization [Amm92]. notation [Fin96, LS90a]. notations [BR95]. note [HH97, Pik06]. null [De 90]. NUMA [ZQ91]. number [AM90, CY96, Ebr97]. numbers [Vos91]. Numerical [Kem96].

O [GSM^{+}95]. Object [KM92, LX93, SMS99, ADS99, BBM96, BA99, BBC96, BC95, BMB99, BDW99, CK94b, CDK98, CSCK95, CHB92, DK94, HCN98, HS97, HNUI94, Lar90, LMR92, MCB90, MM90, SPL98, SD98, WH92]. object-based [BMB99]. Object-oriented [LX93, BBM96, BDW99, CDK98, CHB92, HCN98, HS97, HNUI94, Lar90, LMR92, MCB90, WH92]. Objectcharts [CHB92]. objectives [Hel93]. Objects [KM92, HO90, IK96, Mil90, Mos92, SVK97, Win90]. obligations [CPKM94]. Observer [DJC94]. office [DZ90]. Offutt [Gir93]. on-line [DJC94, GJB96, WRH98, WLS97]. On-the-fly [DS98]. online [HYT91]. ontological [WW90]. open [Min91]. operation [Mor90b]. operational [LW97, ODK99, vLW98]. operations [ZJ96]. opinion [CS94]. opportunism [Sen97]. Optimal [BP90, Gut95, IKWZ90, Mul90, PS93, AM98, BU91, DKKP91, Fal92, GDT92, KH92b, RCGF97, Sch93, Sin91]. optimistic [AM98, HD91]. Optimization [BA93, BJ91, Pek92, Sar95, SS91]. optimized [BBH93]. optimizing [SS95a]. OR-tree [LW90]. oracles [PP98]. Orca [BKT92]. order [Cam90, GPS96, HG99]. organization [SB98]. organizing [Yad90]. orientation [DK94]. oriented [BBM96, BDW99, BLW90, CK94b, CDK98, CSCK95, CHB92, CG92, ESS92, FGNR92, GS99, GDT92, HCN98, HS97, HNUI94, Lar90, LMR92, LX93, MS91, MCB90, MWB99, MCN92, Nic90, OB92, WH92]. Orthogonal [CBC^{+}92, RS96]. OSI [Kar90a, HW89]. OSI-style [Kar90a, HW89]. output [BKWZ94]. overhead [HW89, Kar90a]. owned [Mut92].

packaging [CP91]. page [WF93]. PAISLey [Zav91]. papers [FHLS99]. paradigm [JR99, TT95, WEHL94]. paradigms [VS95]. Parallel [Pek92, AG91, BKT92, Bar90, BB93, BLW90, Bur96, GS90a, GDG92, IM93, JR99, KH92b, LRF95, LW90, LSB94, PC92, PA91, Qin93, SD90, SB99, YPV91, ZQ91]. parallel/distributed [PC92]. parallelism [Aud97, GD93, GS90b]. parallelization [WS97]. parameters [HvB94, TYOJ91]. Parametric [BDL95]. Parnas [XP93a]. parsers [HKR90, SLvH92]. Parsing [TVC94, BB91, CDOT97, SD90, Shi93]. partial [GPS96]. partial-order [GPS96]. partially [ACD98]. partially-implemented [ACD98]. Partition [Gut99, HT90, CY94, Han95, TDN93, WJ91]. partitioning [CI93, CY90]. partly [LM90]. Pascal [NS91]. Pascal-like [NS91]. Passive [ACG93]. Passive-space [ACG93]. path [BM94, Hal95]. paths [PZ95]. pathwise [Hua90]. pattern [BBT92]. patterns [Dev95, PP94]. payment [AB96]. per-object [BA99]. Performance [CB99, DR93, GDG92, GAF91, Liu98, Mor90b, RT90, SS95a, TR91, YD93, Zha91, ZQ91, ACG93, EG90, FM94, GW92, HW90, HNUI94, HWR99, IM93, KSH^{+}92, KMR91, Ran93, RMSF91, SW93, Woo95]. periodic [CS92, GHS95, PSA97, SLL93]. Persistent [KM92, Mos92]. perspective [HM96]. pessimistic [AM98]. Petri [BD91, Bou94, BV95, CCS91, CMBC93, CGL94, CNT99, DS98, DC96, FMM94, Fer99, FN91, Gae96, GMMP91, GDG92, Haa99, HYT91, HT91, LCK98, LSY94, Liu98, NM94, PC92, Suz90, SSM90, TYC95, WK90]. phase [TR91]. philosophers [KM90b]. physical [CY90]. Pictorial [SL92]. pictures [Lam95]. PIE [Voa92]. pipelined [GW92]. place [CVW98]. places [WR96]. Planning [HZO98, GCB^{+}98]. Playground [GSM^{+}95]. point [JLB93, KP92, MT96, RLL93]. pointers [CB93, Kar90b, PLR94]. points [AR96, LJ90, MBM94, MRT91]. Poisson [CS94, Sah92]. policies [CB93, KS94b]. policy [Bos95, CS92, LR92, LS99, RT90]. policy-based [LS99]. port [SVK97]. port-based [SVK97]. possibilistic [McL96]. possibly [BC90]. postconditions [BR95]. potential [SS95b]. PQL [Jar98]. practical [Das92, Das93, KL94, Ros95, SWO97]. practice [AN96, Bar95, LSHZ94]. Pre [CDM95, AS97, SG91]. Pre-run-time [CDM95, AS97, SG91]. precedence [AM90, XP90, Xu93]. Precise [PMI94]. Predicate [Par93, BG91, Kar90b, ROS98, Tai96]. predicate/transition [BG91]. predicates [HMRS98, VD95]. predict [RW97]. Predicting [HS91b, OA96]. Prediction [KWM92, FN99, Jor95, KdBMdS91, ZQ91]. Predictive [KMBR92]. preemptive [CW90, RCGF97]. presence [WR96]. presentation [Jac91, JFH92]. preserving [GN95]. preventing [YG90]. prevention [Bar90]. price [ZA91]. primary [HT93]. primary-site [HT93]. primitives [BDS93]. priority [BTW95, HKL94, KAS93, LS95, RCGF97, Ves94, WM95]. Prism [MS91, SKN96, Mad92]. Prism-methodology [MS91]. privately [Mut92]. probabilities [ML95]. probability [MMN^{+}92, Rom91]. problem [BMR95, HS93, KK98, KM90b, LMM96]. problems [EH91, TDN93]. procedure [BMR95, CCHK90, Kit98, PT98]. procedures [Nov97]. process [ADS99, BFG93, BFL^{+}95, BHT^{+}93, BPV94, CHBC93, CB94, CM96, Che97, CBC^{+}92, CGL94, Cug98, JC93, JLHM91, JO97, KM90a, KK99, KB95, LHHR94, LJ90, MS91, MK92a, MCN92, Sch99, Sen97, SSV99, SO98, TBH91]. process-control [CB94, JLHM91, LHHR94]. process-oriented [MS91, MCN92]. processes [AR96, CNW90, CVW98, EJ90, GHS95, HT91, SO98, VAA99, VV96, XP90, Xu93]. processing [CDIY90, CDY92, DKKP91, KH92b, ZQ91]. processor [Bar95, GST91, Kau90, Tow86, Tow90]. processors [AM90, CGP99, GAF91, LR95, LBJ^{+}95]. product [Dro95, FN91]. production [CCdC92, CCdC93, GS90a, Hu97, PKJH99]. productivity [BSV96, GK91, HS91b, MVD96]. profiles [Hal95]. Profiling [RLP90]. Program [Han95, KNE92, ACG93, BC90, CCF94, DH92, Dev95, Eng99, GL91, Gri94, GN95, HT90, HM93, HC90, HW94, Jar98, Kor97, LV97, LC92, MW92, MDC92, MW90, NTFT91, PP94, PP96, PP98, PC90, Smi90, VF92, WR94b]. Programmer [RWG90, GSM^{+}95]. Programming [HLR92, NS91, BKT92, Bis90, BLW90, CCTCR92, GH91, Gov93, HTI90, How90, JR99, LJS^{+}90, Liu98, MKH97, MR98, MWB99, Ost95, RC90, Ros95, Tay92]. programs [Aud97, Bar90, BM94, Bur96, CT98, CB94, Che98, Cor96b, Dil90, GH95, GPS96, Gol90, GMP90, Hua90, KSB89, Kau90, Kor97, LMR92, LSB94, LB94, MO90, Mos90, Muh91, MK92b, PMI94, Qin93, RGB93, SB99, Sta93, SHH91, TCO91, Tai96, TLK92, Tow86, Tow90, TVK90, VD95, WM95, WF98, WH92, Yad90, YD95, YPV91]. project [AHSR93, SS97, TBH91]. Projecting [AE92]. projection [Che86, Hir90]. projects [LSHZ94]. projects-bridging [LSHZ94]. prolegomena [ORSvH95]. Prolog [DW90, DA91, HS91a, Kar90b, NJ91]. Prolog-based [NJ91]. prone [MK92b, OA96, SB91]. Proof [CKA93, Sto94, CPKM94]. Properties [BMB97, LJS91, Dil90, FMM94, FG97, Jac91, McL96, PZ91, PDB^{+}97, VMC90]. Property [BMB96, BMB97, PDB^{+}97, Zus97]. Property-based [BMB96, BMB97, PDB^{+}97, Zus97]. Protean [BCD97]. protection [WC90]. protocol [Che86, Hir90, HR98, LR97, Ran93, Rei96, Suz90, SSM90]. protocols [AN96, Aba97, CRL99, CDY92, HT96, HP91, Kai96, KM99, MYK94, Min91, Nes93, Sch98, VMC90]. prototype [BPWC90]. Prototyping [BPV94, LSHZ94, BLY93, CCTCR92, KLB93]. Provable [FW93c]. prover [Che98, Gol90, KM97]. Proving [FMM94]. Prudent [AN96]. pump [KML96, MH92]. purpose [KR95]. PVS [ORSvH95, ROS98].

QBD* [ACS90]. QBE [YS93]. QDA [HW94]. Qualitative [Sea99]. Quality [FF96, BBM96, Dro95, KMBR92, MIS97, MA93, WH94, ZEWH95]. quality-of-service [MIS97]. quantification [WMS^{+}92]. quantifying [BF93]. quantitative [JO97, ODK99]. quasi [FM94]. quasi-lumpable [FM94]. queries [QHK91, TVK90]. Query [SS91, ACS90, GD93, PP96, WMS^{+}92]. Querying [NP92, DCN93]. queue [LS95]. queueing [FN91]. queues [Joh95]. quorum [MYK94].

R [Foo93, Gir93, YLT93, KdBMdS91]. race [KB90]. races [KSB89, MO90]. random [CY94, CM96, CY96, CW90, Gut99, HS90, SL92, TDN93, Vos91]. randomly [SK91]. Rapid [Fea98, WF93]. Rapide [LKA^{+}95]. rate [BC90, CW90, Hel93, SLL93]. rate-monotonic [SLL93]. ratio [MvMS93]. reachability [CGK98, NM94]. reactive [ACJ^{+}96, BA99, CGP99, HLN^{+}90]. reader [Joh95]. reader/writer [Joh95]. Real [Ano97d, AS97, ACD98, BS91b, BTW95, BF93, BC99, CBMW93, CSL91, FMM94, FL97, GL92, GH95, GHS95, HS90, HLR92, HS93, HKL94, HPW98, HNUI94, Hou90, JLHM91, JM94, KS94a, KLB93, KS93, LS90b, LS95, MC97, NSY92, PSA97, Qin93, RRH93, SZ92, Sha92, SG91, SVK97, SMYP99, SHH91, TFCB90, TYC95, WS90, XP93a, XP93b, YMST99]. Real-time [Ano97d, AS97, ACD98, BS91b, BTW95, BF93, BC99, CBMW93, CSL91, FMM94, FL97, GL92, GH95, GHS95, HS90, HLR92, HS93, HKL94, HNUI94, Hou90, JLHM91, JM94, KS94a, KLB93, KS93, LS90b, LS95, MC97, NSY92, PSA97, Qin93, RRH93, SZ92, Sha92, SG91, SVK97, SMYP99, TFCB90, TYC95, WS90, YMST99]. reality [CGR95]. realizations [EHO94]. realtime [CPGK97]. Reasoning [WR96, Cha90b, FH92, MW90, RF92, SR92]. reasons [vG91]. Recalibrating [BCLS90]. recapture [VV93]. recently [LM95]. recognition [BBT92, KNE92]. reconfigurable [SVK97]. Reconfiguration [VAA99]. record [Chu91]. recovery [HPR^{+}97, KB95, LW92, MRT91, NvS90, NG90, Puc92, RS93, TMT93, Tri96, WF93]. recursion [ACS90, SV92]. recursions [HL91]. redistributing [GS90b]. reduce [CDM95]. Reducing [De 90, KK99]. reduction [Bin97, BBH93]. redundancy [BJ91, ECK^{+}91]. Reference [Ano96b, CTV^{+}98]. references [GW92]. refinement [GH95, Hal95, MH92, MMM97a, MQR95, PZ91, RW94]. Refining [BMB97, PDB^{+}97]. Reflections [LZ95]. Regeneration [AT93, HH97]. Region [GS90b, LRF95]. regional [CDY92]. regression [Bin97, MS99, RW97, RH96, RH98]. regular [BG91, KS99b, Lus94]. rejection [CW90]. related [FN98]. relation [Zhu96]. relational [CI93, CY90, Fal92, LS90a, MM90, NJ91, TVK90, Wan96, YLN88, YCHL92]. relations [MC92, SPL98, SI91, SWO97, XP90, Xu93]. relationship [BDS93, CY94, MM90]. relationships [NK91, NKF94, PZ95]. RELAY [RT93]. release [SGL97, XP90, Xu93]. relevant [Smi91]. Reliability [Lev90, Ada96, ABK94, AB96, BJ91, BA93, BCLS90, BF93, ECK^{+}91, FHLS98, HZO98, KdBMdS91, KL94, KWM92, KP92, LKBK91, LK92, Sah92, Sar95, Sch93, Sch99, Tia95, TLP95, TYOJ91, TDN93, ZA91]. rely [KR93]. remaining [Ebr97]. remarks [DMW^{+}95]. remote [JR99, Mor90b, SG90]. rendezvous [NWPM95]. reorganization [CB99]. replacement [CCH96]. replacements [MK93]. replay [TFCB90]. replicated [KM99, PVB95, Sin90]. replication [TT95]. Reply [KPF97, Zus97, Con90, DMW^{+}95, Hir90, YSW^{+}01]. report [ACJ^{+}96]. Repository [BKZ93, BKWZ94]. repository-based [BKWZ94]. Representation [Edw97, JFH92, DC96, FH92, FP94, HM93, NJ91, RF92, SR92, TWJ92]. representations [Nov95b, QSvMG96, SWO97]. Representing [MCN92, OB92]. requirement [DFPT90, TWJ92]. Requirements [CRCK98, HPW98, LF91, LHHR94, AG93b, BW94, DS97b, ELC^{+}98, FKV91, GHS95, HRS98, HL96, HKL^{+}98, JLHM91, JFH92, LCK98, LB91, Moo90, MCN92, NKF94, PVB95, Qin93, RD92, Ran93, RRH93, RW91, RP99, SM98, SMMM98, vLW98, vLDL98, RW91]. research [MMM95]. resilient [IKWZ90]. resolution [KS94c, LF91]. Resource [HT93, BB93, CSL91, GHS95, SGL97]. resource-based [GHS95]. resources [HBRY91, Qin93]. Response [BC96, BMB97, RGRF01, WM95, BC99, CSL91, Gai91]. Response-time [WM95]. restricted [HvB94]. restructuring [GN95, GCB^{+}98]. resulting [AW95]. results [KH92b, PJ97]. retrieval [Cha90b, MBK91, SL92]. retrieving [MMM97a]. retrospective [KZB^{+}91]. Reusability [Di 93, BLW90]. reusability-oriented [BLW90]. reusable [FP94, HW91, IK96, LV97]. reuse [BKZ93, BKWZ94, CL95, FF96, HR96, LL97, NHWS92, Nov95b, Nov97, RS96, Sen97]. Reusing [MMM95]. reveals [MMN^{+}92]. reverse [CCdC92, CCdC93]. review [PJ97]. reviewer [Ano99]. revisited [Aba97]. reward [MT96, QSvMG96]. rewriting [AG94b, CCH96, WP94]. rigorous [BGRS91, RM93]. rings [FD94, LSY94]. RISC [LBJ^{+}95]. risk [BBH93, Sch99]. robustness [MS97b]. role [Sen97]. rollback [RS93]. Rough [WW91]. RPC [HR94b]. RSF [DFPT90]. RTSS’96 [Ano97d]. Rule [EH91, CBMW93, IKP95, RGB93, WM95]. Rule-based [EH91, CBMW93, IKP95, RGB93, WM95]. rules [CKA93, LW97, NJ91, Sto94]. run [AS97, CDM95, SG91, TM98]. run-time [TM98]. Rushby [Pik06].

safe [Edw97, RH98, YMST99]. safety [Dil90, FN98, HRS98, HKL^{+}98, Jac95, LW97]. safety-critical [Jac95]. SARA [LB91]. satisfying [XP93a, XP93b]. save [LDvB94]. Scalar [GW92]. scale [PSTV97, Sat92, Tri96, VV96]. scanning [BB91]. scenario [CRCK98, SMMM98]. scenario-based [CRCK98, SMMM98]. scenarios [DFKM98, DF98, RSA98, SC98, vLW98]. scenes [HPW98]. schedulability [SHH91, TYC95]. schedule [CDM95]. scheduler [BS91b, DIM95]. scheduler-luck [DIM95]. schedulers [BTW95, IKK90, KAS93]. schedules [LW92]. Scheduling [XP90, AS97, AM90, BGRS91, CDM95, CW90, FL97, GST91, GS90b, HS90, HKL94, HYT91, Hel93, KS94b, PS93, PSA97, RCGF97, SZ92, SG91, SLL93, WS97, Xu93]. schemas [BDS93, MM90]. scheme [Kun91, Wan96, YLN88, vML99]. schemes [BB90, Sar95, YD92]. Schlumberger [WR94b]. Schnorr [War96]. scientific [Fen94, HR94a]. SDL [LDvB94]. search [Dev95, IK96, LM90, PP94]. searches [LW90]. SeaView [LDS^{+}90]. secondary [BP90]. Section [Ano97c, CS99, CG96, FH97, FT98, Gar98, Jaz98, McH96, Mea96, NJ96, Son97, GN99, HM99, RW99]. secure [FR96, GQ96, HT96, Ran93, Rei96, TGC90]. Security [FG97, Bos95, HMT^{+}90, JO97, KZB^{+}91, LS90c, LDS^{+}90, McC90, ODK99, SPL96, Smi91]. security-relevant [Smi91]. Seesoft [ESS92]. selection [BP90, FvBK^{+}91, LvBP94, MRT91, RT93, RH96, RH98, Sch93, TZ95]. selectivities [Chu91]. self [FD94, LCKS90]. self-stabilizing [FD94]. semantic [MW90, SD98]. Semantically [Fra92]. Semantics [Bin97, FTTV96, BBC96, BC95, KLB93, Smi91]. Semantics-based [FTTV96]. Semaphore [LS95]. Semi [AG91, NTFT91]. Semi-automatic [NTFT91]. Semi-distributed [AG91]. semiautomatic [Smi90]. semijoins [Mul90]. sensitivity [Ves94]. Separate [HR96]. sequences [BU91]. sequencing [CT98, OO90]. sequential [DFKM98]. server [Hem90, NWPM95]. Service [CD99, FC99, FR96, Gai91, Hel93, KK98, MIS97, YG90]. services [ABBC99, JZ98, KM99]. set [BBH93, Fal92, GST91, HCN98]. sets [KS99b]. Seven [RF92]. shared [CMS96, DS90, EHO94, GGH90, RC90, Zha91]. shared-memory [DS90]. sharing [ML95, Mut92, SS95b, SGL97]. shotgun [MRF^{+}98]. SIMD [KMR91]. simple [LS91]. Simulating [WP94]. Simulation [RLL93, BS91a, CNT99, De 90, Fer99, Gae96, HGR92, KIR^{+}99]. simulations [BL94]. single [PLR94]. site [HT93]. size [BKWZ94, Gra91, HS97, Lar90, VT92]. slices [Kor97]. slicing [GL91, LV97].

Smalltalk [Ost95, YD95]. Smooth [Ano96b, DL91, Foo93]. SOFL [LOHS98a, LOHS^{+}98b]. soft [RCGF97]. Software [AHSR93, Ano95, Ano96b, BFG93, BMB97, CPDGM91, Fen94, HS97, JLHM91, JV99, KL94, KPF97, Lar90, LI95, LJ98, LDvB94, MBM94, MVD96, MK92a, NWPM95, Nov97, Sch93, SW93, WH94, YZM97, ZA91, AL95, Ada96, AE92, AG94b, AW95, BFL^{+}95, BKZ93, BKWZ94, BJ91, BA93, BHT^{+}93, BH98, BSV96, BBT92, BBH93, BMB96, BKL90, BCLS90, BF93, CGZ91, CHBC93, CCH96, CAB^{+}98, CS91, CDK98, CNW90, CBS99, CVW98, Cor96a, CER90, DM94, Das92, Das93, DL91, DC95, Di 93, Dro95, DB98, Ebe92, Ebr97, ECK^{+}91, ESS92, FN98, FN99, FGNR92, Foo93, FP94, FF96, FHLS98, Fre91, GL91, GN97, GK91, GBB90, GHM98, GJB96, Gut95, HRS98, HW91, Har92, HCN98, HR94a, HRPL^{+}95, HZO98, HWR99, Hu97, HS91b, IW95, IK96]. software [JD96, Jor95, KdBMdS91, KWM92, KM90a, KS99a, KMBR92, KPF95, Kor90, KMR91, KK99, Le 98, LP98b, LL97, Lev90, LCKS90, LSHZ94, LX93, LW97, LJ90, Luq90, MBK91, MTMS92, ML95, MIS97, MS92, MMM95, MMM97a, MH91, MA93, MBB^{+}97, MS97b, MWB99, MMM^{+}97b, NK91, NOTS98, NG90, NP92, NM94, NHWS92, Nov95b, OA96, OB92, PC92, Par93, PZ91, PZ93, PK91, PKJH99, PN97, PC90, PDB^{+}97, PVB95, PJ97, PSTV97, PSTS91, PJ91, Raj92, RF92, Sah92, Sar95, Sch92, Sea99, Sen97, SR92, SDK^{+}95, SS97, SL91, Sin91, SR93, SO94, SO98, SF95, SVK97, TMA^{+}96, Tia95, TLP95, TYOJ91, TMT93, TMN^{+}99, TFCB90, VV93, VT92, WP94, WC90, WT94, WR94a, WR94b, Woo95, Yad90, Zhu96, Zus97, ZEWH95, vG91, vML99, Ano96a, Ano97a]. solution [CDS99, DS98]. solutions [FN91]. solvers [HS93]. solving [EH91, Sar95]. Some [DMW^{+}95, LW97, VMC90, AAA^{+}90, KH92b, PZ93, TDN93]. sorting [SM94]. source [CCDD98, Dev95, PP94]. Space [SB99, ACG93, MVD96]. SPADE [BFG93, BDF96]. SPADE-1 [BDF96]. SPARE [VF92]. Spawn [WHH^{+}92]. Special [Ano97c, Ano98b, CS99, CG96, FH97, FT98, Gar98, Jaz98, McH96, Mea96, NJ96, Son97, DS97a, GN99, HM99, JKS98, LL97, Raj92, RW99, RG98, Sar95]. specialization [CPDGM91, Nov97]. specialized [BA99]. specific [ABBC99, CRL99, HRPL^{+}95, KS99b, TMC99]. Specification [AG94a, BLS91, BG91, Bos95, CPGK97, HS93, HDL90, Lam90, LKA^{+}95, PC92, ACJ^{+}96, BLY93, BCD97, CT98, CL95, EJ90, Fin96, Fra92, FKV91, GCG90, GS99, HMT^{+}90, HR98, IW95, JM94, KR93, KS94a, LHHR94, Moo90, ND90, NKF94, PMdO99, Smi92, SC96, SSM90, TWJ92, WEHL94, WGS94, WMS^{+}92, WK90, YG90, ZJ96]. specification-based [BLY93, CT98, SC96]. specificational [LC92]. specifications [BMR95, BW94, CAB^{+}98, DFPT90, FMM94, FGH^{+}94, GGH90, HKL^{+}98, LDvB94, MH91, Nic90, NSY92, NTFT91, NP92, PF90, ROS98, TYC95, vLW98]. specified [CBMW93, TVK90]. specify [BDM96, NOTS98, Win90]. Specifying [Hem90, Jac95, Lus94, RRH93, HC90]. spectral [Fel90]. speculative [YMST99]. speed [BSV96]. speedup [SD90]. SPiCE [YD95]. SPIN [BCY01, Hol97]. SQL [YS93]. Stability [Gai91, CB99]. stabilizing [FD94]. standards [EFM^{+}99, FN98]. standby [CB94]. Starvation [KB90]. State [AG93b, Hua90, IKP95, FD94, FvBK^{+}91, HL96, LS90a, LvBP94, QSvMG96, Sha92]. State-based [AG93b, HL96]. state-level [QSvMG96]. statecharts [PMdO99, SR98, CHB92]. stateful [VAA99]. STATEMATE [HLN^{+}90]. statements [SG95]. states [Nic90]. Static [LS90b, Jar98, OO90, WS97]. Statically [YMST99]. Statistical [HL93, Che96, Ebr97, WT94]. statistics [ESS92]. steps [Eng99]. stepsize [HG99]. Stepwise [KS93]. Stochastic [PA91, YPV91, Bou94, CCJS94, CDS99, CMBC93, CGL94, CNT99, DS98, FN91, FM94, GDG92, Haa99, HYT91, HT91, Liu98, MT96, QSvMG96, Ser96]. stop [DM94]. stopping [LW97]. storage [KP91, TGC90, WF93]. Storing [MMM97a]. Strategies [Bar90, RW97, SR93, WJ91]. strategy [ECK^{+}91, FN98]. streams [LM95, NS91]. strength [KM97]. strings [KS99b]. strong [TN97]. Structural [TLK92, GH95, LSY94, SD98]. structure [CDIY90, DM90, SB91, WM95, ZA91]. Structured [CDS99, CB94, LM90, PMI94]. structures [BB93, CB93, CB99, MM90, Mil90, Puc92, QSvMG96, Sta95]. Structuring [BDS93, MIS97, Buc99]. studies [PJ97, RH98, Sea99]. study [AL95, AR96, ADS99, BG91, BHT^{+}93, DH92, FEG92, FP94, GGJ^{+}90, GW92, GSW92, KP92, LL97, LCKS90, MH92, MA93, SB98, SV92, SW93, SR93, SKN96, WRH98, Wey90, vG91]. studying [KS99a, MS92]. style [HW89, JD96, Kar90a, TMA^{+}96]. styles [Le 98]. subdomain [CY96]. subjectivity [BG97]. Subprogram [DH92]. subset [Wan96, YLN88]. subsume [Zhu96]. subtransport [Ran93]. subtransport-level [Ran93]. subtree [MK93]. Subtypes [ROS98]. subtyping [ROS98]. success [Rom91]. Sufficient [MT96]. suite [CK94b, CSCK95, GN97, HM96]. suites [AW95]. superposed [Buc99, Kem96, Kem99]. superposition [LM95]. supervised [KSH^{+}92]. Support [LMR92, MK93, Cug98, GCB^{+}98, HJR91, HO90, LP98a, PMdO99, RUPT90, RF92, SDK^{+}95, WH92, YZM97, vML99]. Supporting [BDF96, FGNR92, IK96, RD92, SMMM98, CGK98]. survey [BSV96, KK98]. swapping [HW91, WEHL94]. switches [OA96]. switching [KdBMdS91]. swizzle [Mos92]. SWSL [KS94a]. symbolic [AHH96, CPDGM91]. Symposium [Ano97d]. synchronization [LS95, Sin90]. synchronous [BdS96, HLR92]. syntactic [DC95]. synthesis [AL95, BH98, Che96, EJ90, HR94b, LB91, MW92, SR92]. synthesizing [Wan96, YLN88]. synthetic [KS94a]. system [AG93b, BG91, BS91b, CP91, CCDD98, CNR90, CDIY90, CDFP97, DA91, DS97b, GSW92, HTI90, HR94b, Jac95, KIR^{+}99, KdBMdS91, KIT93, Kau90, KR95, LI95, LJS^{+}90, LB91, LKA^{+}95, MRF^{+}98, MMM97a, Moo90, NAT93, NBK95, PS93, Qin93, RUPT90, RS93, RP93, RGFR98, Sat92, SB91, Smi90, SC98, SK91, Tow86, Tow90, TN97, TYC95, WRH98, WW90, WW91, XP93a, YSW^{+}01, YD95]. Systematic [Pik06, Rus99, HW94, vML99]. Systems [Ano96b, Ano97d, LJ98, AS97, AT93, AG91, AHH96, AB96, ACJ^{+}96, AK98a, AB94, ABC^{+}91, ACDW94, ACD98, BLS91, BS91a, BKT92, BDM96, BA93, BD91, BDW99, BV95, Buh98, CMSW93, CGP99, CBMW93, CK94a, CSL91, CI93, CDY92, CPGK97, CDOT97, Cug98, DM94, DT94, DK94, DJC94, FMM94, FH92, GN97, GGJ^{+}90, GS90a, GL92, GMMP91, GST91, GDG92, GDT92, GCB^{+}98, HW90, HJR91, HLR92, HKL94, HLN^{+}90, HM95, HYT91, HRPL^{+}95, HS97, HPR^{+}97, HNUI94, Hou90, HW89, HT93, IM93, JLHM91, JM94, Kar90a, KK98, KS94a, KH92b, KS93, LS90a, Lar90, Lev90, LHHR94, LS90b, LS99, Lus94, MKH97, MS92, Min91, Mul90, NWPM95, PLR94, PC92, PW92, PSA97, PA91, PZC^{+}96, RD92, RJT93, RRH93, RM93, RCGF97, RGRF01, RMSF91, RvH93, SG91]. systems [SGCT90, Sin90, SR93, SS95b, TT95, Tri96, TGC90, TC95, WS92, WS90, XP93b, You97, YMST99]. systems-experience [ACJ^{+}96].

T [Hir90]. T.-Y [Hir90]. T9000 [Bar95]. TAME [OB92]. Tandem [LI95]. Task [CSL91, CER90, Jor95, MRT91]. tasking [Cor96b, Dil90]. tasks [CW90, FL97, GDT92, HS90, Hem90, MC97, PS93, PSA97, RCGF97, SZ92]. Teamwork [HJR91]. Teapot [CRL99]. Technique [KM92, CMS98, GKNV93, HD91, RH98, Sin90, Voa92, Yad90, ZJ96]. Techniques [JC93, Bis90, CMSW93, DS98, Ebe92, JLB93, KMBR92, RH96]. technologies [MWB99]. Telecommunications [LJ98, JZ98, KK98]. telephone [KR93, OA96]. Temporal [KB91, YLT93, LSY94, Nes93, SR98, Suz90]. term [AG94b]. Test [FvBK^{+}91, LvBP94, TLP95, AW95, Bin97, BU91, CHBC93, DO91, DMW^{+}95, Fel90, GN97, Gir93, Gut95, HvB94, Kor90, PZ91, PP98, RT93, RH96, RH98, Sch99, TVK90, WGS94, Zhu96]. Test-execution-based [TLP95]. Testability [Fre91, BS96]. Testing [VD95, ABK94, CT98, CY94, CM96, CY96, CDFP97, DM94, FW93a, FW93b, FW93c, FHLS98, GCG90, Gut99, HT90, HS91a, KMR91, LW97, LDvB94, MMN^{+}92, Mor90a, PZ93, PZ95, PC90, PZC^{+}96, RW97, Sin91, SR93, SC96, Tai96, TLK92, TFCB90, TDN93, Wey90, WJ91, Wey93, WT94]. tests [PF90]. textual [MS97a]. their [HT96, TYC95]. them [SDK^{+}95]. theorem [KM97, LS94, LSY94, McC90]. theoretical [KL94]. Theory [Tai96, Cam90, DC95, HM96, LS94, LSHZ94, McL96, Mor90a, SM98]. Thorough [Wan96]. threads [SZ92]. three [DCN93, DR93, Woo95]. three-dimensional [DCN93]. three-view [Woo95]. throughout [KM90a]. throughput [CCS91, CCJS94]. time [AS97, ACG93, AM90, Ano97d, ACDW94, ACD98, BD91, BS91b, BV95, BTW95, BF93, BC99, CS94, CDM95, CBMW93, CNW90, CSL91, DH92, DKKP91, DIM95, FMM94, FL97, Fer99, GL92, GH95, GHS95, GCG90, GMMP91, HS90, HLR92, HS93, HKL94, HNUI94, Hou90, JLHM91, JM94, KS94a, KLB93, KS93, LS90b, LS95, MC97, NSY92, PSA97, Pik06, Qin93, RS93, RRH93, Rus99, SZ92, Sha92, SG91, Sin91, SVK97, SMYP99, SHH91, Tia95, TFCB90, TYC95, TM98, Ves94, WM95, WS90, XP93a, XP93b, YMST99, GAF91]. time-critical [BV95, GMMP91]. time-triggered [Pik06, Rus99]. timed [Fer99, KR93, Liu98, MH92, VMC90]. times [BC99, CCH96, HS90, SGL97, WR96, XP90, Xu93, YPV91]. Timing [Cor96b, HKL94, TYC95, GH95, LBJ^{+}95, XP93a, XP93b]. TLA [Lam95]. TMN [LR97]. token [FD94, VMC90]. tolerance [AK98b, SGCT90, SL91]. tolerant [AG93a, BG91, HT93, ORSvH95, Pik06, PJ91, Rus99, TC95, WW91]. Tolerating [Cug98]. too [Fin96]. Tool [GCB^{+}98, CGZ91, ESS92, FG97, Fra92, GN95, IM93, NOTS98]. tools [CCdC92, CCdC93, MS97a, ODK99, SDK^{+}95, WK90]. toolset [ABC^{+}91]. Total [Ada96, DKKP91]. Towsley [Kau90]. TPL [LJS^{+}90]. Trace [HWR99, Cam90, WP94]. Trace-based [HWR99]. tracing [KIT93]. Tractable [CK94a]. Tradeoffs [BB90, CMSW93, GN95]. trail [HL93]. transaction [BGRS91, CI93, CDIY90, CDY92, GDT92, Lus94, WF93]. transaction-based [CI93, Lus94]. transaction-oriented [GDT92]. transaction-undo [WF93]. Transactions [Ano95, Ano96a, Ano97a, BLS91]. transformation [KNE92, LKBK91, War96]. transformations [BLY93, HC90, KSH^{+}92]. Transient [Kem99, AJ93, HG99]. transition [BG91, IKP95, LS90a, PC92]. transitive [QHK91]. translating [YD95]. transport [Sin90]. transportable [LJS^{+}90]. tree [LM90, LW90, MRT91, Tia95]. tree-based [Tia95]. tree-structured [LM90]. trees [Aud97, KP91, KS99b]. trend [KL94]. Trie [LRLH91]. triggered [Pik06, Rus99]. TROPICO [KdBMdS91]. TROPICO-R [KdBMdS91]. truly [Nes93]. Trust [Ran93]. TuringTool [CER90]. tutoring [NBK95]. twin [WF93]. twin-page [WF93]. Two [FD94, WMS^{+}92, Aba97, CMSW93, PJ97, TR91]. Two-dimensional [WMS^{+}92]. two-phase [TR91]. Two-state [FD94]. type [PT98, SWO97]. types [HDL90, SD98]. typestate [SY93].

unbalanced [Kit98]. unbounded [GDG92]. uncertainty [Gut99]. underlying [CGL94]. Understanding [FPV98]. undo [WF93]. unified [BDW99, GMMP91, HM93, MYK94]. uniform [Jac91]. uniprocessor [ACDW94]. unique [CCS91]. units [Nov95a]. universal [WMS^{+}92]. UNIX [KIT93, Che97]. Unix-based [Che97]. Unreachable [Nic90]. unsafe [LW92]. unstructured [Kor97]. unsuccessful [LM90]. updatable [Mil90]. Update [Sin90, CS92]. usability [LP98a]. usage [CGR95, Hel93]. Use [Buh98, CT98, RS93, BS96, CDK98, CHB92, DF98, GST91, LCK98, LCKS90, PLR94, SC98]. used [LM95]. User [UGB93, CER90, GSM^{+}95, MS97a, MMM^{+}97b, Sea93]. user-configurable [GSM^{+}95]. uses [PZ95]. Using [AG94b, BLY93, BCD97, DA91, GL91, GPS96, HKL^{+}98, KB95, LR97, PP98, RW97, Win90, BD91, BV95, CB93, Che98, CPGK97, Dev95, DKKP91, ELC^{+}98, FF96, GBB90, GDG92, GS93, HDL90, HYT91, HR98, IW95, KWM92, Le 98, LCK98, Liu98, LKA^{+}95, LvBP94, MBM94, MRF^{+}98, Moo90, MCN92, NOTS98, NTFT91, Ost95, PP94, PS93, RT93, RSA98, Sch99, SPL98, SS97, SI91, SR93, SSV99, SL92, SVK97, SY93, Tia95, VV93, VD95, WEHL94, WF93, YD93]. utility [ZA91]. utilize [Qin93].

validating [BMB99, DF98, Sch92]. Validation [KPF97, AG94a, AAA^{+}90, BBM96, BG97, Bos95, BV95, DJC94, GPS96, HPW98, KPF95, LF91, MBB^{+}97]. value [Ser96]. variable [Kar90b, SGL97]. variance [Ada96]. various [Zha91]. VAX [KZB^{+}91]. VDM [BR95, FEG92]. vector [ACG93, CCS91]. verification [AL95, AHH96, BD91, Che98, Che86, EHO94, FG97, GL92, GMP90, HDL90, Hir90, KS91, ORSvH95, Pik06, RvH93, Rus99, SSM90, WM97, You97, YG90]. verified [Moo90, WLS97]. verify [AG94b, SWO97]. Verifying [AL95, Dil90, Sch98, Sta95, Gol90, HLR92, RRH93]. version [BKL90, GJB96, PW93]. vertical [CI93, CY90]. via [ABK94, AK98b, CPDGM91, Cug98, SR99]. video [TMC99]. view [ACG93, Das92, Das93, GHM98, Woo95]. viewpoint [LF91]. viewpoints [SSV99]. views [CL95, KH92a, Nov95b, Nov97, NKF94]. violations [HKL^{+}98]. virtual [AT93, AJ93, Bar95, HH97, JZ98, KH92a]. Visual [EDRM90, MS97a, Cha90b, CDOT97, CGP^{+}90, FTTV96, HMT^{+}90, KP91, HTI90, RWG90]. visualization [Ebe92]. visualizing [ESS92]. VMM [KZB^{+}91]. Vol [Ano95, Ano97a, Ano96a]. voting [LCKS90, RJT93]. vs [Gut99].

Waite [War96]. waiting [RT90]. ware [LK92]. Warm [CB94]. warp [Fer99, GAF91]. weak [OL94]. Web [CD99, MRF^{+}98]. Web-based [MRF^{+}98]. well [FM94, PMI94]. well-formed [FM94]. well-structured [PMI94]. Weyuker [CS91]. Where [ZJ96]. White [Edw97]. widget [Sea93]. within [Das92, Das93, Kau90, Tow86, Tow90, VS95]. WorkBench [RWG90]. Working [Mos92, GST91, HLN^{+}90]. workload [KS94a, Kun91, WS90, YCHL92]. workstation [BPWC90, Mut92]. workstations [JR99]. world [HPW98]. worst [LBJ^{+}95]. Wp [LvBP94]. Wp-method [LvBP94]. write [CS92]. writer [Joh95]. writing [CRL99]. WWW [CTV^{+}98].

X [LK92]. X-ware [LK92]. Xception [CMS98].

Yeast [KR95].

Z [Jac95, NOTS98].

References

Arlat:1990:FID

[AAA^{+}90] J. Arlat, M. Aguera, L. Amat, Y. Crouzet, J. C. Fabre, J. C. Laprie, E. Martins, and D. Powell. Fault injection for dependability validation: a methodology and some applications. IEEE Transactions on Software Engineering, 16(2):166–182, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                                   ieeexplore.ieee.org/stamp/
                                                   stamp.jsp?arnumber=44380.

Aue:1994:DIS

[AB94] A. Aue and M. Breu. Distributed information systems: an advanced methodology. IEEE Transactions on Software Engineering, 20(8):594–605, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310669.

Anderson:1996:REP

[AB96] R. J. Anderson and S. J.

Bezuidenhoudt. On the reliability of electronic payment systems. IEEE Transactions on Software Engineering, 22(5):294–301, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=502222.

Abadi:1997:ECR

[Aba97] M. Abadi. Explicit communication revisited: two new attacks on authentication protocols. IEEE Transactions on Software Engineering, 23(3):185–186, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585505.

Atkins:1999:MDS

[ABBC99] D. L. Atkins, T. Ball, G. Bruns, and K. Cox. Mawl: a domainspecific language for formbased services. IEEE Transactions on Software Engineering, 25(3):334–346, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798323.

Avrunin:1991:AAC

[ABC^{+}91] G. S. Avrunin, U. A. Buy, J. C. Corbett, L. K. Dillon, and J. C. Wileden. Automated analysis of concurrent systems with the constrained expression toolset. IEEE Transactions on Software Engineering, 17(11):1204–1222, November 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106975.

Ammann:1994:EIE

[ABK94] P. E. Ammann, S. S. Brilliant, and J. C. Knight. The effect of imperfect error detection on reliability assessment via life testing. IEEE Transactions on Software Engineering, 20(2):142–148, February 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=265635.

Avrunin:1998:API

[ACD98] G. S. Avrunin, J. C. Corbett, and L. K. Dillon. Analyzing partially-implemented realtime systems. IEEE Transactions on Software Engineering, 24(8):602–614, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707696.

Avrunin:1994:ADT

[ACDW94] G. S. Avrunin, J. C. Corbett, L. K. Dillon, and J. C. Wileden. Automated derivation of time bounds in uniprocessor concurrent systems. IEEE Transactions on Software Engineering, 20(9):708–719, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=317429.

Ahuja:1993:PST

[ACG93] M. Ahuja, T. Carlson, and A. Gahlot. Passive-space and time view: vector clocks for achieving higher performance, program correction, and distributed computing. IEEE Transactions on Software Engineering, 19(9):845–855, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241768.

Ardis:1996:FES

[ACJ^{+}96] M. A. Ardis, J. A. Chaves, L. J.

Jagadeesan, P. Mataga, C. Puchol, M. G. Staskauskas, and J. Von Olnhausen. A framework for evaluating specification methods for reactive systemsexperience report. IEEE Transactions on Software Engineering, 22(6):378–389, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=508312.

Angelaccio:1990:QGQ

[ACS90] M. Angelaccio, T. Catarci, and G. Santucci. QBD*: a graphical query language with recursion. IEEE Transactions on Software Engineering, 16(10):1150–1163, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=60295.

Adams:1996:TVA

[Ada96]  T. Adams.  Total variance

approach to software reliability estimation. IEEE Transactions on Software Engineering, 22(9):687–688, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=541438.

Agarwal:1999:COP

[ADS99] R. Agarwal, P. De, and A. P. Sinha. Comprehending object and process models: an empirical study. IEEE Transactions on Software Engineering, 25(4):541–556, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=799953.

Agresti:1992:PSD

[AE92] W. W. Agresti and W. M.

Evanco. Projecting software defects from analyzing Ada designs. IEEE Transactions on Software Engineering, 18(11):988–997, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177368.

Abd-El-Hafiz:1996:KBA

[AEHB96] S. K. Abd-El-Hafiz and V. R.

Basili. A knowledge-based approach to the analysis of loops. IEEE Transactions on Software Engineering, 22(5):339–360, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=502226.

Ahmad:1991:SDL

[AG91] I. Ahmad and A. Ghafoor.

Semi-distributed load balancing for massively parallel multicomputer systems. IEEE Transactions on Software Engineering, 17(10):987–1004, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99188.

Arora:1993:CCF

[AG93a] A. Arora and M. Gouda.

Closure and convergence: a foundation of fault-tolerant computing. IEEE Transactions on Software Engineering, 19(11):1015–1027, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256850.

Atlee:1993:SBM

[AG93b] J. M. Atlee and J. Gannon. State-based model checking of event-driven system requirements. IEEE Transactions on Software Engineering, 19(1):24–40, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=210305.

Akella:1994:SVC

[AG94a] V. Akella and G. Gopalakrishnan. Specification and validation of control-intensive IC’s in hopCP. IEEE Transactions on Software Engineering, 20(6):405–423, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=295890.

Antoy:1994:UTR

[AG94b] S. Antoy and J. Gannon. Using term rewriting to verify software. IEEE Transactions on Software Engineering, 20(4):259–274, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277574.

Alur:1996:ASV

[AHH96] R. Alur, T. A. Henzinger, and Pei-Hsin Ho. Automatic symbolic verification of embedded systems. IEEE Transactions on Software Engineering, 22(3):181–201, March 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=489079.

Abdel-Hamid:1993:SPC

[AHSR93] T. K. Abdel-Hamid, K. Sengupta, and D. Ronan. Software project control: an experimental investigation of judgment with fallible information. IEEE Transactions on Software Engineering, 19(6):603–612, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232025.

Agrawala:1993:DMT

[AJ93] A. K. Agrawala and B. N.

Jain. Deterministic model and transient analysis of virtual circuits. IEEE Transactions on Software Engineering, 19(2):187–197, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214835.

Arora:1998:CBD

[AK98a] A. Arora and S. S. Kulkarni. Component based design of multitolerant systems. IEEE Transactions on Software Engineering, 24(1):63–78, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=663998.

Arora:1998:DMF

[AK98b] A. Arora and S. S. Kulkarni. Designing masking fault-tolerance via nonmasking fault-tolerance. IEEE Transactions on Software Engineering, 24(6):435–450, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=689401.

Aagaard:1995:VLS

[AL95] M. Aagaard and M. Leeser. Verifying a logic-synthesis algorithm and implementation: a case study in software verification. IEEE Transactions on Software Engineering, 21(10):822–833, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=469458.

Al-Mouhamed:1990:LBN

[AM90] M. A. Al-Mouhamed. Lower bound on the number of processors and time for scheduling precedence graphs with communication costs. IEEE Transactions on Software Engineering, 16(12):1390–1401, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62447.

Alvisi:1998:MLP

[AM98] L. Alvisi and K. Marzullo.

Message logging: pessimistic, optimistic, causal, and optimal. IEEE Transactions on Software Engineering, 24 (2):149–159, February 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=666828.

Ammarguellat:1992:CFN

[Amm92] Z. Ammarguellat. A controlflow normalization algorithm and its complexity. IEEE Transactions on Software Engineering, 18(3):237–251, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126773.

Abadi:1996:PEP

[AN96] M. Abadi and R. Needham. Prudent engineering practice for cryptographic protocols. IEEE Transactions on Software Engineering, 22(1):6–15, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=481513.

Anonymous:1995:IIT

[Ano95] Anonymous. 1995 index IEEE Transactions on Software Engineering vol. 21. IEEE Transactions on Software Engineering, 21(12):1–??, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489077.

Anonymous:1996:IIT

[Ano96a]  Anonymous.  1996 index

IEEE Transactions on Software vol. 22. IEEE Transactions on Software Engineering, 22(12):912–919, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553641.

Anonymous:1996:RMS

[Ano96b] Anonymous. Reference model for smooth growth of software systems (003)5402022. IEEE Transactions on Software Engineering, 22(8):–??, August 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=536959.

Anonymous:1997:IIT

[Ano97a]  Anonymous.  1997 index

IEEE Transactions on Software vol. 23. IEEE Transactions on Software Engineering, 23(12):800–807, December 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637394.

Anonymous:1997:AI

[Ano97b] Anonymous. Author index.

IEEE Transactions on Software Engineering, 23(1):60–61, January 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=581330.

Anonymous:1997:GEI

[Ano97c]  Anonymous.  Guest edito-

rial: Introduction to the special section. IEEE Transactions on Software Engineering, 23(2):65–66, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585496.

Anonymous:1997:RTS

[Ano97d] Anonymous. Real-time systems symposium (rtss’96). IEEE Transactions on Software Engineering, 23(10):0:1, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637142.

Anonymous:1998:AI

[Ano98a] Anonymous. Author index. IEEE Transactions on Software Engineering, 24(1):93–95, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=664000.

Anonymous:1998:SIM

[Ano98b]  Anonymous.  Special issue

— mobility and network-aware computing. IEEE Transactions on Software Engineering, 24(5):0:1, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=685254.

Anonymous:1999:RL

[Ano99]  Anonymous.  1998 reviewer

list. IEEE Transactions on Software Engineering, 25(1):141– 144, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=748923.

Abran:1996:FPA

[AR96] A. Abran and P. N. Robillard. Function points analysis: an empirical study of its measurement processes. IEEE Transactions on Software Engineering, 22(12):895–910, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553638.

Abdelzaher:1997:CPR

[AS97] T. F. Abdelzaher and K. G.

Shin. Comment on “A pre-run-time scheduling algorithm for hard real-time systems”. IEEE Transactions on Software Engineering, 23(9):599–600, September 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=629495.

See [SG91].

Adam:1993:RVC

[AT93] N. R. Adam and R. Tewari. Regeneration with virtual copies for distributed computing systems. IEEE Transactions on Software Engineering, 19(6): 594–602, June 1993. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
232024.

Audenaert:1997:CTL

[Aud97] K. Audenaert. Clock trees: logical clocks for programs with nested parallelism. IEEE Transactions on Software Engineering, 23(10):646–658, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637147.

Avritzer:1995:AGL

[AW95] A. Avritzer and E. R. Weyuker. The automatic generation of load test suites and the assessment of the resulting software. IEEE Transactions on Software Engineering, 21 (9):705–716, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=464549.

Berman:1993:OMR

[BA93] O. Berman and N. Ashrafi. Optimization models for reliability of modular software systems. IEEE Transactions on Software Engineering, 19(11):1119–1123, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=256858.

Bertrand:1999:BSL

[BA99] F. Bertrand and M. Augeraud. BDL: a specialized language for per-object reactive control. IEEE Transactions on Software Engineering, 25(3):347–362, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798324.

Barbosa:1990:SPC

[Bar90] V. C. Barbosa. Strategies for the prevention of communication deadlocks in distributed parallel programs. IEEE Transactions on Software Engineering, 16(11):1311–1316, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60319.

Barrett:1995:MCP

[Bar95] G. Barrett. Model checking in practice: the T9000 virtual channel processor. IEEE Transactions on Software Engineering, 21(2):69–78, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345823.

Balasubramanian:1990:TDE

[BB90] V. Balasubramanian and P. Banerjee. Tradeoffs in the design of efficient algorithm-based error detection schemes for hypercube multiprocessors. IEEE Transactions on Software Engineering, 16(2):183–196, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44381.

Beetem:1991:ISP

[BB91] J. F. Beetem and A. F.

Beetem. Incremental scanning and parsing with galaxy. IEEE Transactions on Software Engineering, 17(7):641–651, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83901.

Biswas:1993:DSP

[BB93] J. Biswas and J. C. Browne. Data structures for parallel resource management. IEEE Transactions on Software Engineering, 19(7):672–686, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238568.

Botting:1996:CFS

[BBC96] R. J. Botting, R. H. Bourdeau, and B. H. C. Cheng. Comments on “A formal semantics for object model diagrams”. IEEE Transactions on Software Engineering, 22(12):911–??, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553639.

See [BC95] and response [BC96].

Briand:1993:DIM

[BBH93] L. C. Briand, V. R. Brasili, and C. J. Hetmanski. Developing interpretable models with optimized set reduction for identifying high-risk software components. IEEE Transactions on Software Engineering, 19(11):1028–1044, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256851.

Basili:1996:VOO

[BBM96] V. R. Basili, L. C. Briand, and W. L. Melo. A validation of object-oriented design metrics as quality indicators. IEEE Transactions on Software Engineering, 22(10):751–761, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=544352.

Briand:1992:PRA

[BBT92] L. C. Briand, V. R. Basili, and W. M. Thomas. A pattern recognition approach for software engineering data analysis. IEEE Transactions on Software Engineering, 18(11):931–942, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177363.

Becker:1990:BEM

[BC90] G. Becker and L. Camarinopoulos. A Bayesian estimation method for the failure rate of a possibly correct program. IEEE Transactions on Software Engineering, 16(11):1307–1310, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60318.

Bourdeau:1995:FSO

[BC95] R. H. Bourdeau and B. H. C. Cheng. A formal semantics for object model diagrams. IEEE Transactions on Software Engineering, 21(10):799–821, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=469459.

See comments [BBC96].

Bourdeau:1996:CRB

[BC96] R. H. Bourdeau and B. H. C. Cheng. Correspondence: Response to Botting’s comments. IEEE Transactions on Software Engineering, 22(12):911, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553640.

See [BBC96].

Buttazzo:1999:MAR

[BC99] G. C. Buttazzo and M. Caccamo. Minimizing aperiodic response times in a firm real-time environment. IEEE Transactions on Software Engineering, 25(1):22–32, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=748916.

Bloom:1997:UPL

[BCD97] B. Bloom, A. Cheng, and A. Dsouza. Using a Protean language to enhance expressiveness in specification. IEEE Transactions on Software Engineering, 23(4):224–234, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588539.

Brocklehurst:1990:RSR

[BCLS90] S. Brocklehurst, P. Y. Chan, B. Littlewood, and J. Snell. Recalibrating software reliability models. IEEE Transactions on Software Engineering, 16(4):458–470, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54297.

Bang:2001:CMC

[BCY01] Ki-Seok Bang, Jin-Young Choi, and Chuck Yoo. Comments on “The Model Checker SPIN”. IEEE Transactions on Software Engineering, 27(6):573–576, June 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=926177.

See [Hol97].

Berthomieu:1991:MVT

[BD91] B. Berthomieu and M. Diaz.

Modeling and verification of time dependent systems using time Petri nets. IEEE Transactions on Software Engineering, 17(3):259–273, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=75415.

Bandinelli:1996:SCS

[BDF96] S. Bandinelli, E. Di Nitto, and A. Fuggetta. Supporting cooperation in the SPADE-1 environment. IEEE Transactions on Software Engineering, 22(12):841–865, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553634.

Bertolazzi:1995:PGD

[BDL95] P. Bertolazzi, G. Di Battista, and G. Liotta. Parametric graph drawing. IEEE Transactions on Software Engineering, 21(8):662–673, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403790.

Battiston:1996:MAN

[BDM96] E. Battiston, F. De Cindio, and G. Mauri. Modular algebraic nets to specify concurrent systems. IEEE Transactions on Software Engineering, 22(10):689–705, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=544348.

Batini:1993:SPD

[BDS93] C. Batini, G. Di Battista, and G. Santucci. Structuring primitives for a dictionary of entity relationship data schemas. IEEE Transactions on Software Engineering, 19(4):344–365, April 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=223803.

Boussinot:1996:SSL

[BdS96] F. Boussinot and R. de Simone.

The SL synchronous language. IEEE Transactions on Software Engineering, 22(4):256–266, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=491649.

Braccini:1991:ICD

[BDV91] A. Braccini, A. Del Bimbo, and E. Vicario. Interprocess communication dependency on network load. IEEE Transactions on Software Engineering, 17(4):357–369, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90435.

Briand:1999:UFC

[BDW99] L. C. Briand, J. W. Daly, and J. K. Wust. A unified framework for coupling measurement in object-oriented systems. IEEE Transactions on Software Engineering, 25 (1):91–121, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=748920.

Butler:1993:IQR

[BF93] R. W. Butler and G. B. Finelli. The infeasibility of quantifying the reliability of lifecritical real-time software. IEEE Transactions on Software Engineering, 19(1):3–12, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=210303.

Bandinelli:1993:SPM

[BFG93] S. C. Bandinelli, A. Fuggetta, and C. Ghezzi. Software process model evolution in the SPADE environment. IEEE Transactions on Software Engineering, 19(12):1128–1144, December 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=249659.

Bandinelli:1995:MII

[BFL^{+}95] S. Bandinelli, A. Fuggetta, L. Lavazza, M. Loi, and G. P. Picco. Modeling and improving an industrial software process. IEEE Transactions on Software Engineering, 21(5):440–454, May 1995. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=387473.

Belli:1991:SFT

[BG91] F. Belli and K. E. Grosspietsch. Specification of faulttolerant system issues by predicate/transition nets and regular expressions — approach and case study. IEEE Transactions on Software Engineering, 17(6):513–526, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87278.

Batory:1997:CVS

[BG97] D. Batory and B. J. Geraci.

Composition validation and subjectivity in GenVoca generators. IEEE Transactions on Software Engineering, 23(2):67–82, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585497.

Bolliger:1998:FBA

[BG98] J. Bolliger and T. Gross. A framework based approach to the development of network aware applications. IEEE Transactions on Software Engineering, 24(5):376–390, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=685260.

Breitbart:1991:RTS

[BGRS91] Y. Breitbart, D. Georgakopoulos, M. Rusinkiewicz, and A. Silberschatz. On rigorous transaction scheduling. IEEE Transactions on Software Engineering, 17(9):954–960, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92915.

Bhansali:1998:ASS

[BH98] S. Bhansali and T. J. Hoar. Automated software synthesis: an application in mechanical CAD. IEEE Transactions on Software Engineering, 24(10):848–862, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=729684.

Bhandari:1993:CSS

[BHT^{+}93] I. Bhandari, M. Halliday, E. Tarver, D. Brown, J. Chaar, and R. Chillarege. A case study of software process improvement during development. IEEE Transactions on Software Engineering, 19(12):1157–1170, December 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=249661.

Binkley:1997:SGR

[Bin97] D. Binkley. Semantics guided regression test cost reduction. IEEE Transactions on Software Engineering, 23(8):498–516, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=624306.

Bishop:1990:EDA

[Bis90] J. M. Bishop. The effect of data abstraction on loop programming techniques. IEEE Transactions on Software Engineering, 16(4):389–402, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54291.

Belli:1991:ARO

[BJ91] F. Belli and P. Jedrzejowicz. An approach to the reliability optimization of software with redundancy. IEEE Transactions on Software Engineering, 17(3):310–312, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=75419.

Bieman:1998:MDL

[BK98] J. M. Bieman and Byung-Kyoo Kang. Measuring designlevel cohesion. IEEE Transactions on Software Engineering, 24(2):111–124, February 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=666825.

Brilliant:1990:AFE

[BKL90] S. S. Brilliant, J. C. Knight, and N. G. Leveson. Analysis of faults in an N-version software experiment. IEEE Transactions on Software Engineering, 16(2):238–247, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44387.

Bal:1992:OLP

[BKT92] H. E. Bal, M. Frans Kaashoek, and Andrew S. Tanenbaum. Orca: a language for parallel programming of distributed systems. IEEE Transactions on Software Engineering, 18(3):190–205, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126768.

Banker:1994:AOS

[BKWZ94] R. D. Banker, R. J. Kauffman, C. Wright, and D. Zweig. Automating output size and reuse metrics in a repository-based computer-aided software engineering (CASE) environment. IEEE Transactions on Software Engineering, 20(3):169–187, March 1994. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=268919.

Banker:1993:RES

[BKZ93] R. D. Banker, R. J. Kauffman, and D. Zweig. Repository evaluation of software reuse. IEEE Transactions on Software Engineering, 19(4):379–389, April 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=223805.

Bagrodia:1994:MLD

[BL94] R. L. Bagrodia and Wen-Toh Liao. Maisie: a language for the design of efficient discreteevent simulations. IEEE Transactions on Software Engineering, 20(4):225–238, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277572.

Babin:1991:SDT

[BLS91] G. Babin, F. Lustman, and P. Shoval. Specification and design of transactions in information systems: a formal approach. IEEE Transactions on Software Engineering, 17(8):814–829, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83916.

Browne:1990:EER

[BLW90] J. C. Browne, T. Lee, and J. Werth. Experimental evaluation of a reusability-oriented parallel programming environment. IEEE Transactions on Software Engineering, 16(2):111–120, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=44375.

Berzins:1993:UTS

[BLY93] V. Berzins, Luqi, and A. Yehudai. Using transformations in specification-based prototyping. IEEE Transactions on Software Engineering, 19(5):436–452, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=232011.

Bertolino:1994:AGP

[BM94] A. Bertolino and M. Marre. Automatic generation of path covers based on the control flow analysis of computer programs. IEEE Transactions on Software Engineering, 20(12):885–899, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368137.

Briand:1996:PBS

[BMB96] L. C. Briand, S. Morasca, and V. R. Basili. Propertybased software engineering measurement. IEEE Transactions on Software Engineering, 22(1):68–86, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=481535.

See comments [PDB^{+}97, Zus97] and response [BMB97].

Briand:1997:RCP

[BMB97] L. C. Briand, S. Morasca, and V. R. Basili. Response to: Comments on “Propertybased Software Engineering Measurement: Refining The Additivity Properties”. IEEE Transactions on Software Engineering, 23(3):196–197, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585509.

See [BMB96, PDB^{+}97, Zus97].

Briand:1999:DVM

[BMB99] L. C. Briand, S. Morasca, and V. R. Basili. Defining and validating measures for object-based high-level design. IEEE Transactions on Software Engineering, 25(5): 722–743, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=815329.

Borgida:1995:FPP

[BMR95] A. Borgida, J. Mylopoulos, and R. Reiter. On the frame problem in procedure specifications. IEEE Transactions on Software Engineering, 21(10):785–798, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=469460.

Bieman:1994:MFC

[BO94] J. M. Bieman and L. M. Ott. Measuring functional cohesion. IEEE Transactions on Software Engineering, 20(8):644–657, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=310673.

Boswell:1995:SVS

[Bos95] A. Boswell. Specification and validation of a security policy model. IEEE Transactions on Software Engineering, 21(2):63–68, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=345822.

Boucherie:1994:CIC

[Bou94]  R. J. Boucherie.  A char-

acterization of independence for competing Markov chains with applications to stochastic Petri nets. IEEE Transactions on Software Engineering, 20(7):536–544, July 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=297942.

Barcucci:1990:OSS

[BP90] E. Barcucci and R. Pinzani. Optimal selection of secondary indexes. IEEE Transactions on Software Engineering, 16(1):32–38, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44361.

Bradac:1994:PPM

[BPV94] M. G. Bradac, D. E. Perry, and L. G. Votta. Prototyping a process monitoring experiment. IEEE Transactions on Software Engineering, 20(10):774–784, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=328994.

Berger:1990:CMW

[BPWC90] J. L. Berger, J. Picciotto, J. P. L. Woodward, and P. T. Cummings. Compartmented mode workstation: prototype highlights. IEEE Transactions on Software Engineering, 16(6):608–618, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=55089.

Burke:1990:CAI

[BR90] M. G. Burke and B. G. Ryder. A critical analysis of incremental iterative data flow analysis algorithms. IEEE Transactions on Software Engineering, 16(7):723–728, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=56098.

Bicarregui:1995:IFP

[BR95] J. Bicarregui and B. Ritchie. Invariants, frames and postconditions: a comparison of the VDM and B notations. IEEE Transactions on Software Engineering, 21(2):79–89, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=345824.

Bagrodia:1991:MID

[BS91a] R. L. Bagrodia and C. C. Shen. MIDAS: integrated design and simulation of distributed systems. IEEE Transactions on Software Engineering, 17(10):1042–1058, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=99192.

Blake:1991:EER

[BS91b] B. A. Blake and K. Schwan.

Experimental evaluation of a real-time scheduler for a multiprocessor system. IEEE Transactions on Software Engineering, 17(1):34–44, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67577.

Bertolino:1996:UTM

[BS96] A. Bertolino and L. Strigini. On the use of testability measures for dependability assessment. IEEE Transactions on Software Engineering, 22(2):97–108, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=485220.

Basili:1999:BKT

[BSL99] V. R. Basili, F. Shull, and F. Lanubile. Building knowledge through families of experiments. IEEE Transactions on Software Engineering, 25(4):456–473, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799939.

Blackburn:1996:ISP

[BSV96] J. D. Blackburn, G. D. Scudder, and L. N. Van Wassenhove. Improving speed and productivity of software development: a global survey of software developers. IEEE Transactions on Software Engineering, 22(12):875–885, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553636.

Burns:1995:EAE

[BTW95] A. Burns, K. Tindell, and A. Wellings. Effective analysis for engineering real-time fixed priority schedulers. IEEE Transactions on Software Engineering, 21(5):475–480, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=387477.

Boyd:1991:CGO

[BU91] S. C. Boyd and H. Ural. On the complexity of generating optimal test sequences. IEEE Transactions on Software Engineering, 17(9):976–978, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92918.

Buchholz:1999:HSS

[Buc99]  P. Buchholz.  Hierarchi-

cal structuring of superposed GSPNs. IEEE Transactions on Software Engineering, 25 (2):166–181, March/April 1999.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761443.

Buhr:1998:UCM

[Buh98] R. J. A. Buhr. Use case maps as architectural entities for complex systems. IEEE Transactions on Software Engineering, 24(12):1131–1155, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738343.

Burton:1996:GGM

[Bur96] F. W. Burton. Guaranteeing good memory bounds for parallel programs. IEEE Transactions on Software Engineering, 22(10):762–773, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=544353.

Bucci:1995:CVT

[BV95] G. Bucci and E. Vicario. Compositional validation of timecritical systems using communicating time Petri nets. IEEE Transactions on Software Engineering, 21(12):969–992, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=489073.

Bustard:1994:MCF

[BW94] D. W. Bustard and A. C. Winstanley. Making changes to formal specifications: requirements and an example. IEEE Transactions on Software Engineering, 20(8):562–568, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=310666.

Chan:1998:MCL

[CAB^{+}98] W. Chan, R. J. Anderson, P. Beame, S. Burns, F. Modugno, D. Notkin, and J. D. Reese. Model checking large software specifications. IEEE Transactions on Software Engineering, 24(7):498–520, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=708566.

Camilleri:1990:MCT

[Cam90] A. J. Camilleri. Mechanizing CSP trace theory in higher order logic. IEEE Transactions on Software Engineering, 16(9):993–1004, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58786.

Chen:1993:MAC

[CB93] Ing-Ray Chen and S. A. Banawan. Modeling and analysis of concurrent maintenance policies for data structures using pointers. IEEE Transactions on Software Engineering, 19(9):902–911, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241772.

Chen:1994:WSH

[CB94] Ing-Ray Chen and F. B. Bastani. Warm standby in hierarchically structured processcontrol programs. IEEE Transactions on Software Engineering, 20(8):658–663, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310674.

Chen:1999:PSA

[CB99] I. R. Chen and S. A. Banawan. Performance and stability analysis of multilevel data structures with deferred reorganization. IEEE Transactions on Software Engineering, 25(5): 690–700, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=815327.

Chillarege:1992:ODC

[CBC^{+}92] R. Chillarege, I. S. Bhandari, J. K. Chaar, M. J. Halliday, D. S. Moebus, B. K. Ray, and M. Y. Wong. Orthogonal defect classification — a concept for in-process measurements. IEEE Transactions on Software Engineering, 18(11):943–956, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=177364.

Cheng:1993:ART

[CBMW93] A. M. K. Cheng, J. C. Browne, A. K. Mok, and Rwo-Hsi Wang. Analysis of real-time rule-based systems with behavioral constraint assertions specified in Estella. IEEE Transactions on Software Engineering, 19(9):863–885, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=241770.

Chulani:1999:BAE

[CBS99] S. Chulani, B. Boehm, and B. Steece. Bayesian analysis of empirical software engineering cost models. IEEE Transactions on Software Engineering, 25(4):573–583, July/ August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=799958.

Canfora:1992:LBA

[CCdC92] G. Canfora, A. Cimitile, and U. de Carlini. A logic-based approach to reverse engineering tools production. IEEE Transactions on Software Engineering, 18(12):1053–1064, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=184760.

See correction [CCdC93].

Canfora:1993:CAL

[CCdC93] G. Canfora, A. Cimitile, and U. de Carlini. Correction to “A logic-based approach to reverse engineering tools production (Dec 92 1053–1064). IEEE Transactions on Software Engineering, 19(6):640–??, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232029.

See [CCdC92].

Canfora:1998:ESS

[CCDD98] G. Canfora, A. Cimitile, U. De Carlini, and A. De Lucia. An extensible system for source code analysis. IEEE Transactions on Software Engineering, 24(9):721–740, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=713328.

Choi:1994:EEA

[CCF94] Jong-Deok Choi, R. Cytron, and J. Ferrante. On the efficient engineering of ambitious program analysis. IEEE Transactions on Software Engineering, 20(2):105–114, February 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=265631.

Chan:1996:EME

[CCH96] Taizan Chan, Siu Leung Chung, and Teck Hua Ho. An economic model to estimate software rewriting and replacement times. IEEE Transactions on Software Engineering, 22(8):580–598, August 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=536958.

Callahan:1990:CPC

[CCHK90] D. Callahan, A. Carle, M. W. Hall, and K. Kennedy. Constructing the procedure call multigraph. IEEE Transactions on Software Engineering, 16(4):483–487, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=54302.

Campos:1994:ATC

[CCJS94] J. Campos, J. M. Colom, H. Jungnitz, and M. Silva. Approximate throughput computation of stochastic marked graphs. IEEE Transactions on Software Engineering, 20 (7):526–535, July 1994. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
297941.

Campos:1991:ETB

[CCS91] J. Campos, G. Chiola, and M. Silva. Ergodicity and throughput bounds of Petri nets with unique consistent firing count vector. IEEE Transactions on Software Engineering, 17(2):117–125, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67593.

Cacace:1992:DPD

[CCTCR92] F. Cacace, S. Ceri, L. Tanca, and S. Crespi-Reghizzi. Designing and prototyping dataintensive applications in the Logres and Algres programming environment. IEEE Transactions on Software Engineering, 18(6):534–546, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142875.

Cardelli:1999:SCW

[CD99] L. Cardelli and R. Davies.

Service combinators for Web computing. IEEE Transactions on Software Engineering, 25(3):309–316, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798321.

Cohen:1997:ASA

[CDFP97] D. M. Cohen, S. R. Dalal, M. L. Fredman, and G. C. Patton. The AETG system: an approach to testing based on combinatorial design. IEEE Transactions on Software Engineering, 23(7):437–444, July 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=605761.

Ciciani:1990:HDC

[CDIY90] B. Ciciani, D. M. Dias, B. R. Iyer, and P. S. Yu. A hybrid distributed centralized system structure for transaction processing. IEEE Transactions on Software Engineering, 16(8):791–806, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=57619.

Chidamber:1998:MUM

[CDK98] S. R. Chidamber, D. P. Darcy, and C. F. Kemerer. Managerial use of metrics for object-oriented software: an exploratory analysis. IEEE Transactions on Software Engineering, 24(8):629–639, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707698.

Cavalieri:1995:PRT

[CDM95] S. Cavalieri, A. Di Stefano, and O. Mirabella. Pre-run-time scheduling to reduce schedule length in the FieldBus environment. IEEE Transactions on Software Engineering, 21(11):865–880, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=473215.

Costagliola:1997:PMI

[CDOT97] G. Costagliola, A. De Lucia, S. Orefice, and G. Tortora. A parsing methodology for the implementation of visual systems. IEEE Transactions on Software Engineering, 23(12):777–799, December 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637392.

Campos:1999:SSA

[CDS99] J. Campos, S. Donatelli, and M. Silva. Structured solution of asynchronously communicating stochastic modules. IEEE Transactions on Software Engineering, 25(2):147–165, March/ April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761442.

Ciciani:1992:ACC

[CDY92] B. Ciciani, D. M. Dias, and P. S. Yu. Analysis of concurrency-coherency control protocols for distributed transaction processing systems with regional locality. IEEE Transactions on Software Engineering, 18(10):899–914, October 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=163606.

Cordy:1990:TUI

[CER90] J. R. Cordy, N. L. Eliot, and M. G. Robertson. TuringTool: a user interface to aid in the software maintenance task. IEEE Transactions on Software Engineering, 16(3):294–301, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=48937.

Cui:1992:DOE

[CG92] Q. Cui and J. Gannon. Dataoriented exception handling. IEEE Transactions on Software Engineering, 18(5):393–401, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=135772.

Ciardo:1996:GEI

[CG96] G. Ciardo and C. Ghezzi. Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 22(9):601–??, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541431.

Chen:1998:CDM

[CGK98] Yih-Fam Chen, E. R. Gansner, and E. Koutsofios. A C++ data model supporting reachability analysis and dead code detection. IEEE Transactions on Software Engineering, 24(9):682–694, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=713323.

Ciardo:1994:CSP

[CGL94] G. Ciardo, R. German, and C. Lindemann. A characterization of the stochastic process underlying a stochastic Petri net. IEEE Transactions on Software Engineering, 20(7):506–515, July 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=297939.

Crimi:1990:AVL

[CGP^{+}90] C. Crimi, A. Guercio, G. Pacini, G. Tortora, and M. Tucci. Automating visual language generation. IEEE Transactions on Software Engineering, 16(10):1122–1135, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=60293.

Caspi:1999:ADR

[CGP99] P. Caspi, A. Girault, and D. Pilaud. Automatic distribution of reactive systems for asynchronous networks of processors. IEEE Transactions on Software Engineering, 25(3):416–427, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798329.

Craigen:1995:FMR

[CGR95] D. Craigen, S. Gerhart, and T. Ralston. Formal methods reality check: industrial usage. IEEE Transactions on Software Engineering, 21(2):90–98, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=345825.

Cardenas-Garcia:1991:MTE

[CGZ91] S. Cardenas-Garcia and M. V. Zelkowitz. A management tool for evaluation of software design. IEEE Transactions on Software Engineering, 17(9):961–971, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92916.

Chang:1990:BAC

[Cha90a] C. K. Chang. Bidding against competitors. IEEE Transactions on Software Engineering, 16(1):100–104, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44368.

Chang:1990:VLC

[Cha90b] S. K. Chang. A visual language compiler for information retrieval by visual reasoning. IEEE Transactions on Software Engineering, 16 (10):1136–1149, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60294.

Coleman:1992:IOH

[CHB92] D. Coleman, F. Hayes, and S. Bear. Introducing Objectcharts or how to use Statecharts in object-oriented design. IEEE Transactions on Software Engineering, 18(1):8–18, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120312.

Chaar:1993:PES

[CHBC93] J. K. Chaar, M. J. Halliday, I. S. Bhandari, and R. Chillarege. In-process evaluation for software inspection and test. IEEE Transactions on Software Engineering, 19(11):1055–1070, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256853.

Cheung:1986:PMP

[Che86] T. Y. Cheung. On the projection method for protocol verification. IEEE Transactions on Software Engineering, SE-12(11):1088–1089, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312998.

See [Hir90].

Chernak:1996:SAI

[Che96] Y. Chernak. A statistical approach to the inspection checklist formal synthesis and improvement. IEEE Transactions on Software Engineering, 22(12):866–874, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553635.

Chen:1997:CAL

[Che97] Jen-Yen Jason Chen. CSPL: an Ada95-like, Unix-based process environment. IEEE Transactions on Software Engineering, 23(3):171–184, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585504.

Chetali:1998:FVC

[Che98] B. Chetali. Formal verification of concurrent programs using the larch prover. IEEE Transactions on Software Engineering, 24(1):46–62, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=663997.

Chu:1991:CAE

[Chu91] P. C. Chu. A contingency approach to estimating record selectivities. IEEE Transactions on Software Engineering, 17(6):544–552, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87280.

Chu:1993:TBA

[CI93] W. W. Chu and I. T. Ieong.

A transaction-based approach to vertical partitioning for relational database systems. IEEE Transactions on Software Engineering, 19(8):804–812, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238583.

Cheung:1994:TDA

[CK94a] Shing Chi Cheung and J. Kramer. Tractable dataflow analysis for distributed systems. IEEE Transactions on Software Engineering, 20(8):579–593, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=310668.

Chidamber:1994:MSO

[CK94b] S. R. Chidamber and C. F. Kemerer. A metrics suite for object oriented design. IEEE Transactions on Software Engineering, 20(6):476–493, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=295895.

See comments [CSCK95].

Camp:1993:PRF

[CKA93] T. Camp, P. Kearns, and M. Ahuja. Proof rules for flush channels. IEEE Transactions on Software Engineering, 19(4):366–378, April 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=223804.

See addendum [Sto94].

Cowan:1995:ADV

[CL95] D. D. Cowan and C. J. P. Lucena. Abstract data views: an interface specification concept to enhance design for reuse. IEEE Transactions on Software Engineering, 21(3):229–243, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372150.

Chen:1996:BMP

[CM96] Sanping Chen and S. Mills. A binary Markov process model for random testing. IEEE Transactions on Software Engineering, 22(3):218–223, March 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=489081.

Chiola:1993:GSP

[CMBC93] G. Chiola, M. A. Marsan, G. Balbo, and G. Conte. Generalized stochastic Petri nets: a definition at the net level and its implications. IEEE Transactions on Software Engineering, 19(2):89–107, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214828.

Clemencon:1996:DSA

[CMS96] C. Clemencon, B. Mukherjee, and K. Schwan. Distributed shared abstractions (DSA) on multiprocessors. IEEE Transactions on Software Engineering, 22(2):132–152, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=485223.

Carreira:1998:XTE

[CMS98] J. Carreira, H. Madeira, and J. G. Silva. Xception: a technique for the experimental evaluation of dependability in modern computers. IEEE Transactions on Software Engineering, 24(2):125–136, February 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=666826.

Cabrera:1993:IAT

[CMSW93] L. F. Cabrera, J. A. McPherson, P. M. Schwarz, and J. C. Wyllie. Implementing atomicity in two systems: techniques, tradeoffs, and experience. IEEE Transactions on Software Engineering, 19(10):950–961, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=245737.

Chen:1990:CIA

[CNR90] Y. F. Chen, M. Y. Nishimoto, and C. V. Ramamoorthy. The C information abstraction system. IEEE Transactions on Software Engineering, 16(3):325–334, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48940.

Ciardo:1999:DES

[CNT99] G. Ciardo, D. M. Nicol, and K. S. Trivedi. Discrete-event simulation of fluid stochastic Petri nets. IEEE Transactions on Software Engineering, 25(2):207–217, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761446.

Chmura:1990:ESD

[CNW90] L. J. Chmura, A. F. Norcio, and T. J. Wicinski. Evaluating software design processes by analyzing change data over time. IEEE Transactions on Software Engineering, 16(7):729–740, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=56099.

Constantine:1990:CRC

[Con90]  L. L. Constantine.  Com-

ments, with reply, on “On criteria for module interfaces” by D. Hoffman. IEEE Transactions on Software Engineering, 16(12):1440–??, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62453.

See [Hof90].

Corbett:1996:EDD

[Cor96a]  J. C. Corbett.  Evaluating

deadlock detection methods for concurrent software. IEEE Transactions on Software Engineering, 22(3):161–180, March 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489078.

Corbett:1996:TAA

[Cor96b] J. C. Corbett. Timing analysis of Ada tasking programs. IEEE Transactions on Software Engineering, 22(7):461–483, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=538604.

Callahan:1991:PSH

[CP91] J. R. Callahan and J. M. Purtilo. A packaging system for heterogeneous execution environments. IEEE Transactions on Software Engineering, 17(6):626–635, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=87286.

Coen-Porisini:1991:SSS

[CPDGM91] A. Coen-Porisini, F. De Paoli, C. Ghezzi, and D. Mandrioli. Software specialization via symbolic execution. IEEE Transactions on Software Engineering, 17(9):884–899, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92907.

Coen-Porisini:1997:SRS

[CPGK97] A. Coen-Porisini, C. Ghezzi, and R. A. Kemmerer. Specification of realtime systems using ASTRAL. IEEE Transactions on Software Engineering, 23(9):572–598, September 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=629494.

Coen-Porisini:1994:FFA

[CPKM94] A. Coen-Porisini, R. A. Kemmerer, and D. Mandrioli. A formal framework for ASTRAL intralevel proof obligations. IEEE Transactions on Software Engineering, 20(8):548–561, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310665.

Carroll:1998:RDS

[CRCK98] J. M. Carroll, M. B. Rosson, G. Chin, and J. Koenemann. Requirements development in scenario-based design. IEEE Transactions on Software Engineering, 24(12):1156–1170, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=738344.

Chandra:1999:TDS

[CRL99] S. Chandra, B. Richards, and J. R. Larus. Teapot: a domainspecific language for writing cache coherence protocols. IEEE Transactions on Software Engineering, 25(3):317–333, May/ June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798322.

Cherniavsky:1991:WAS

[CS91] J. C. Cherniavsky and C. H. Smith. On Weyuker’s axioms for software complexity measures. IEEE Transactions on Software Engineering, 17(6):636–638, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=87287.

Carson:1992:APU

[CS92] S. D. Carson and S. Setia. Analysis of the periodic update write policy for disk cache. IEEE Transactions on Software Engineering, 18(1):44–54, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=120315.

Campodonico:1994:BAL

[CS94] S. Campodonico and N. D. Singpurwalla. A Bayesian analysis of the logarithmic– Poisson execution time model based on expert opinion and failure data. IEEE Transactions on Software Engineering, 20(9):677–683, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=317426.

Chiola:1999:GEI

[CS99] G. Chiola and W. H. Sanders.

Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 25(2):145–146, March/ April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761441.

Churcher:1995:CMS

[CSCK95] N. I. Churcher, M. J. Shepperd, S. Chidamber, and C. F. Kemerer. Comments on “A metrics suite for object oriented design”. IEEE Transactions on Software Engineering, 21(3):263–265, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372153.

See [CK94b].

Chu:1991:TRT

[CSL91] W. W. Chu, C. M. Sit, and K. K. Leung. Task response time for real-time distributed systems with resource contentions. IEEE Transactions on Software Engineering, 17(10):1076–1092, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=99195.

Carver:1998:USC

[CT98] R. H. Carver and Kuo-Chung Tai. Use of sequencing constraints for specification-based testing of concurrent programs. IEEE Transactions on Software Engineering, 24(6):471–490, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=689403.

Ciancarini:1998:CMA

[CTV^{+}98] P. Ciancarini, R. Tolksdorf, F. Vitali, D. Rossi, and A. Knoche. Coordinating multiagent applications on the WWW: a reference architecture. IEEE Transactions on Software Engineering, 24(5):362–375, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=685259.

Cugola:1998:TDP

[Cug98] G. Cugola. Tolerating deviations in process support systems via flexible enactment of process models. IEEE Transactions on Software Engineering, 24(11):982–1001, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=730546.

Cook:1998:CEA

[CVW98] J. E. Cook, L. G. Votta, and A. L. Wolf. Cost-effective analysis of in-place software processes. IEEE Transactions on Software Engineering, 24(8):650–663, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707700.

Craig:1990:RRT

[CW90] D. W. Craig and C. M. Woodside. The rejection rate for tasks with random arrivals, deadlines, and preemptive scheduling. IEEE Transactions on Software Engineering, 16(10):1198–1208, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60299.

Cornell:1990:EAV

[CY90] D. W. Cornell and P. S. Yu.

An effective approach to vertical partitioning for physical design of relational databases. IEEE Transactions on Software Engineering, 16(2):248–258, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=44388.

Chen:1994:RBP

[CY94] T. Y. Chen and Y. T. Yu. On the relationship between partition and random testing. IEEE Transactions on Software Engineering, 20(12):977–980, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368132.

Chen:1996:ENF

[CY96] T. Y. Chen and Y. T. Yu. On the expected number of failures detected by subdomain testing and random testing. IEEE Transactions on Software Engineering, 22(2):109–119, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=485221.

Dotan:1991:UFC

[DA91] Y. Dotan and B. Arazi. Using flat concurrent Prolog in system modeling. IEEE Transactions on Software Engineering, 17(6):493–512, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=87277.

Danzig:1994:FCL

[Dan94] P. B. Danzig. Flow control for limited buffer multicast.

IEEE Transactions on Software Engineering, 20(1):1–12, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=263751.

Daskalantonakis:1992:PVS

[Das92] M. K. Daskalantonakis. A practical view of software measurement and implementation experiences within Motorola. IEEE Transactions on Software Engineering, 18 (11):998–1010, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177369.

See correction [Das93].

Daskalantonakis:1993:CPV

[Das93] M. K. Daskalantonakis. Correction to “A practical view of software measurement and implementation experiences within Motorola”. IEEE Transactions on Software Engineering, 19(2):199–200, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214837.

See [Das92].

Dutoit:1998:CMS

[DB98] A. H. Dutoit and B. Bruegge.

Communication metrics for software development. IEEE Transactions on Software Engineering, 24(8):615–628, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=707697.

Dean:1995:STS

[DC95] T. R. Dean and J. R. Cordy. A syntactic theory of software architecture. IEEE Transactions on Software Engineering, 21(4):302–313, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=385969.

Dwyer:1996:CPN

[DC96] M. B. Dwyer and L. A. Clarke. A compact Petri net representation and its implications for analysis. IEEE Transactions on Software Engineering, 22(11):794–811, November 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553699.

DelBimbo:1993:TDI

[DCN93] A. Del Bimbo, M. Campanai, and P. Nesi. A three-dimensional iconic environment for image database querying. IEEE Transactions on Software Engineering, 19(10):997–1011, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=245741.

DeVries:1990:RNM

[De 90] R. C. De Vries. Reducing null messages in Misra’s distributed discrete event simulation method. IEEE Transactions on Software Engineering, 16(1):82–91, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44366.

Devanbu:1995:FSC

[Dev95] P. Devanbu. On “A framework for source code search using program patterns”. IEEE Transactions on Software Engineering, 21(12):1009–1010, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=489076.

See [PP94].

Dzida:1998:MUS

[DF98] W. Dzida and R. Freitag. Making use of scenarios for validating analysis and design. IEEE Transactions on Software Engineering, 24(12):1182–1196, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738346.

Desharnais:1998:ISS

[DFKM98] J. Desharnais, M. Frappier, R. Khedri, and A. Mili. Integration of sequential scenarios. IEEE Transactions on Software Engineering, 24 (9):695–708, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=713325.

DeNicola:1998:KKL

[DFP98] R. De Nicola, G. L. Ferrari, and R. Pugliese. KLAIM: a kernel language for agents interaction and mobility. IEEE Transactions on Software Engineering, 24(5):315–330, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=685256.

DeglInnocenti:1990:RFE

[DFPT90] M. Degl’Innocenti, G. L. Ferrari, G. Pacini, and F. Turini. RSF: a formalism for executable requirement specifications. IEEE Transactions on Software Engineering, 16(11): 1235–1246, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60312.

Davidson:1992:SIS

[DH92] J. W. Davidson and A. M. Holler. Subprogram inlining: a study of its effects on program execution time. IEEE Transactions on Software Engineering, 18(2):89–102, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=121752.

DiFelice:1993:RMS

[Di 93] P. Di Felice. Reusability of mathematical software: a contribution. IEEE Transactions on Software Engineering, 19(8):835–843, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238586.

Dillon:1990:VGS

[Dil90] L. K. Dillon. Verifying general safety properties of Ada tasking programs. IEEE Transactions on Software Engineering, 16(1):51–63, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44363.

Dolev:1995:AET

[DIM95] S. Dolev, A. Israeli, and S. Moran. Analyzing expected time by scheduler-luck games. IEEE Transactions on Software Engineering, 21(5):429–439, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=387472.

Diaz:1994:OCF

[DJC94] M. Diaz, G. Juanole, and J. P. Courtiat. Observer — a concept for formal on-line validation of distributed systems. IEEE Transactions on Software Engineering, 20(12):900–913, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368136.

Deubler:1994:IOO

[DK94] H. H. Deubler and M. Koestler. Introducing object orientation into large and complex systems. IEEE Transactions on Software Engineering, 20 (11):840–848, November 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368124.

Dimitrov:1991:OTP

[DKKP91] B. Dimitrov, Z. Khalil, N. Kolev, and P. Petrov. On the optimal total processing time using checkpoints. IEEE Transactions on Software Engineering, 17(5):436–442, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=90446.

DeMillo:1991:DSC

[DL91] R. A. DeMillo and R. J. Lipton.

Defining software by continuous, smooth functions. IEEE Transactions on Software Engineering, 17(4):383–384, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90437.

Ding:1990:FAD

[DM90] C. Ding and P. Mateti. A framework for the automated drawing of data structure diagrams. IEEE Transactions on Software Engineering, 16(5):543–557, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=52777.

Dalal:1994:WST

[DM94] S. R. Dalal and A. A. McIntosh. When to stop testing for large software systems with changing code. IEEE Transactions on Software Engineering, 20(4):318–323, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277579.

DeMillo:1995:SCR

[DMW^{+}95] R. A. DeMillo, A. P. Mathur, W. E. Wong, P. G. Frankl, and E. J. Weyuker. Some critical remarks on a hierarchy of fault-detecting abilities of test methods [and reply]. IEEE Transactions on Software Engineering, 21(10):858–863, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=469455.

DeMillo:1991:CBA

[DO91] R. A. DeMillo and A. J. Offutt. Constraint-based automatic test data generation. IEEE Transactions on Software Engineering, 17(9):900–910, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92910.

See corrigendum [Gir93].

Delis:1993:PCT

[DR93] A. Delis and N. Roussopoulos. Performance comparison of three modern DBMS architectures. IEEE Transactions on Software Engineering, 19(2):120–138, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214830.

Dromey:1995:MSP

[Dro95] R. G. Dromey. A model for software product quality. IEEE Transactions on Software Engineering, 21(2):146–162, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345830.

Dubois:1990:MAD

[DS90] M. Dubois and C. Scheurich.

Memory access dependencies in shared-memory multiprocessors. IEEE Transactions on Software Engineering, 16(6):660–673, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=55094.

Dillon:1997:ISI

[DS97a] L. K. Dillon and S. Sankar. Introduction to the special issue. IEEE Transactions on Software Engineering, 23(5):265–266, May 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=590647.

Dutertre:1997:FRA

[DS97b] B. Dutertre and V. Stavridou.

Formal requirements analysis of an avionics control system. IEEE Transactions on Software Engineering, 23(5):267–278, May 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588520.

Deavours:1998:FST

[DS98] D. D. Deavours and W. H.

Sanders. “On-the-fly” solution techniques for stochastic Petri nets and extensions. IEEE Transactions on Software Engineering, 24(10):889–902, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=729691.

DePaoli:1994:CLC

[DT94] F. De Paoli and F. Tisato. CSDL: a language for cooperative systems design. IEEE Transactions on Software Engineering, 20(8):606–616, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310670.

Debray:1990:TBC

[DW90] S. K. Debray and D. S. Warren. Towards banishing the cut from Prolog. IEEE Transactions on Software Engineering, 16(3):335–349, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48941.

DeAntonellis:1990:DAO

[DZ90] V. De Antonellis and B. Zonta. A disciplined approach to office analysis. IEEE Transactions on Software Engineering, 16(8):822–828, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57621.

Ebert:1992:CVT

[Ebe92] C. Ebert. Correspondence visualization techniques for analyzing and evaluating software measures. IEEE Transactions on Software Engineering, 18(11):1029–1034, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177373.

Ebrahimi:1997:SAN

[Ebr97] N. B. Ebrahimi. On the statistical analysis of the number of errors remaining in a software design document after inspection. IEEE Transactions on Software Engineering, 23(8):529–532, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=624308.

Ebrahimi:1999:HIC

[Ebr99] N. B. Ebrahimi. How to improve the calibration of cost models. IEEE Transactions on Software Engineering, 25 (1):136–140, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=748922.

Eckhardt:1991:EES

[ECK^{+}91] D. E. Eckhardt, A. K. Caglayan, J. C. Knight, L. D. Lee, D. F. McAllister, M. A. Vouk, and J. P. J. Kelly. An experimental evaluation of software redundancy as a strategy for improving reliability. IEEE Transactions on Software Engineering, 17(7):692–702, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83905.

Eisenstadt:1990:VKE

[EDRM90] M. Eisenstadt, J. Domingue, T. Rajan, and E. Motta. Visual knowledge engineering. IEEE Transactions on Software Engineering, 16(10):1164–1177, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60296.

Edwards:1997:RIS

[Edw97]  S. H. Edwards.  Represen-

tation inheritance: a safe form of “white box” code inheritance. IEEE Transactions on Software Engineering, 23(2):83–92, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585498.

Emmerich:1999:MSC

[EFM^{+}99] W. Emmerich, A. Finkelstein, C. Montangero, S. Antonelli, S. Armitage, and R. Stevens. Managing standards compliance. IEEE Transactions on Software Engineering, 25(6): 836–851, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824413.

Eich:1990:PFG

[EG90] M. H. Eich and S. M. Garard.

The performance of flow graph locking. IEEE Transactions on Software Engineering, 16(4):477–483, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54301.

Etessami:1991:RBD

[EH91] F. A. Etessami and G. S. Hura.

Rule-based design methodology for solving control problems. IEEE Transactions on Software Engineering, 17(3):274–282, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=75416.

Ernst:1994:MVD

[EHO94] G. W. Ernst, R. J. Hookway, and W. F. Ogden. Modular verification of data abstractions with shared realizations. IEEE Transactions on Software Engineering, 20(4):288–307, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277576.

Erdogmus:1990:SSC

[EJ90] M. H. Erdogmus and R. Johnston. On the specification and synthesis of communicating processes. IEEE Transactions on Software Engineering, 16(12):1412–1426, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62449.

Easterbrook:1998:EUL

[ELC^{+}98] S. Easterbrook, R. Lutz, R. Covington, J. Kelly, Y. Ampo, and D. Hamilton. Experiences using lightweight formal methods for requirements modeling. IEEE Transactions on Software Engineering, 24(1):4–14, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=663994.

Elliott:1999:EML

[Ell99] C. Elliott. An embedded modeling language approach to interactive 3d and multimedia animation. IEEE Transactions on Software Engineering, 25(3):291–308, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798320.

Engler:1999:ICS

[Eng99] D. R. Engler. Interface compilation: steps toward compiling program interfaces as languages. IEEE Transactions on Software Engineering, 25(3):387–400, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798327.

Eick:1992:STV

[ESS92] S. C. Eick, J. L. Steffen, and E. E. Sumner. Seesoft — a tool for visualizing line oriented software statistics. IEEE Transactions on Software Engineering, 18(11):957–968, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177365.

Falkowski:1992:COS

[Fal92] B. J. Falkowski. Comments on an optimal set of indices for a relational database. IEEE Transactions on Software Engineering, 18(2):168–171, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=121758.

Fetzer:1999:HAL

[FC99] C. Fetzer and F. Cristian. A highly available local leader election service. IEEE Transactions on Software Engineering, 25(5): 603–618, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=815321.

Flatebo:1994:TSS

[FD94] M. Flatebo and A. K. Datta. Two-state self-stabilizing algorithms for token rings. IEEE Transactions on Software Engineering, 20(6):500–504, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=295897.

Feather:1998:RAL

[Fea98] M. S. Feather. Rapid application of lightweight formal methods for consistency analyses. IEEE Transactions on Software Engineering, 24(11):949–959, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=730544.

Fields:1992:VCS

[FEG92] B. Fields and M. Elvang-Goransson. A VDM case study in mural. IEEE Transactions on Software Engineering, 18(4):279–295, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=129217.

Feldman:1990:NST

[Fel90]  F. A. Feldman.  A new

spectral test for nonrandomness and the DES. IEEE Transactions on Software Engineering, 16(3):261–267, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48934.

Fenton:1994:SMN

[Fen94]  N. Fenton.  Software mea-

surement: a necessary scientific basis. IEEE Transactions on Software Engineering, 20(3):199–206, March 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=268921.

Ferscha:1999:ATW

[Fer99] A. Ferscha. Adaptive time warp simulation of timed Petri nets. IEEE Transactions on Software Engineering, 25(2): 237–257, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761448.

Frakes:1996:QIU

[FF96] W. B. Frakes and C. J. Fox. Quality improvement using a software reuse failure modes model. IEEE Transactions on Software Engineering, 22(4):274–279, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=491652.

Focardi:1997:CSC

[FG97] R. Focardi and R. Gorrieri. The compositional security checker: a tool for the verification of information flow security properties. IEEE Transactions on Software Engineering, 23(9):550–571, September 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=629493.

Finkelstein:1994:IHM

[FGH^{+}94] A. C. W. Finkelstein, D. Gabbay, A. Hunter, J. Kramer, and B. Nuseibeh. Inconsistency handling in multiperspective specifications. IEEE Transactions on Software Engineering, 20(8):569–578, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=310667.

Fischer:1992:SSD

[FGNR92] G. Fischer, A. Girgensohn, K. Nakakoji, and D. Redmiles. Supporting software designers with integrated domain-oriented design environments. IEEE Transactions on Software Engineering, 18(6):511–522, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142873.

Fickas:1992:KRR

[FH92] S. Fickas and B. R. Helm.

Knowledge representation and reasoning in the design of composite systems. IEEE Transactions on Software Engineering, 18(6):470–482, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142870.

Faulk:1997:ISS

[FH97] S. R. Faulk and C. Heitmeyer. Introduction to the special section. IEEE Transactions on Software Engineering, 23(4):201–202, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=590653.

Frankl:1998:ETM

[FHLS98] P. G. Frankl, R. G. Hamlet, B. Littlewood, and L. Strigini.

Evaluating testing methods by delivered reliability [software]. IEEE Transactions on Software Engineering, 24(8):586–601, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707695.

Franki:1999:CP

[FHLS99] P. Franki, D. Hamlet, B. Littlewood, and L. Strigini. Call for papers. IEEE Transactions on Software Engineering, 25(2):286, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761452.

Finney:1996:MNF

[Fin96] K. Finney. Mathematical notation in formal specification: too difficult for the masses? IEEE Transactions on Software Engineering, 22(2):158–159, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=485225.

Fraser:1991:IFR

[FKV91] M. D. Fraser, K. Kumar, and V. K. Vaishnavi. Informal and formal requirements specification languages: bridging the gap. IEEE Transactions on Software Engineering, 17(5):454–466, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90448.

Feng:1997:ASR

[FL97] Wu-Chun Feng and J. W. S.

Liu. Algorithms for scheduling real-time tasks with input error and end-to-end deadlines. IEEE Transactions on Software Engineering, 23(2):93–106, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585499.

Franceschinis:1994:CBP

[FM94] G. Franceschinis and R. R.

Muntz. Computing bounds for the performance indices of quasi-lumpable stochastic wellformed nets. IEEE Transactions on Software Engineering, 20(7):516–525, July 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=297940.

Felder:1994:PPR

[FMM94] M. Felder, D. Mandrioli, and A. Morzenti. Proving properties of real-time systems through logical specifications and Petri net models. IEEE Transactions on Software Engineering, 20(2):127–141, February 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=265634.

Florin:1991:GQN

[FN91] G. Florin and S. Natkin. Generalization of queueing network product form solutions to stochastic Petri nets. IEEE Transactions on Software Engineering, 17(2):99–107, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=67591.

Fenton:1998:SIS

[FN98] N. E. Fenton and M. Neil. A strategy for improving safety related software engineering standards. IEEE Transactions on Software Engineering, 24(11):1002–1013, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=730547.

Fenton:1999:CSD

[FN99] N. E. Fenton and M. Neil. A critique of software defect prediction models. IEEE Transactions on Software Engineering, 25(5): 675–689, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=815326.

Foo:1993:CDS

[Foo93] N. Y. Foo. Comments on “Defining software by continuous smooth functions” by R. A. De Millo and R. J. Lipton. IEEE Transactions on Software Engineering, 19(3):307–309, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221140.

Frakes:1994:ESR

[FP94] W. B. Frakes and T. P. Pole.

An empirical study of representation methods for reusable software components. IEEE Transactions on Software Engineering, 20(8):617–630, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310671.

Fuggetta:1998:UCM

[FPV98] A. Fuggetta, G. P. Picco, and G. Vigna. Understanding code mobility. IEEE Transactions on Software Engineering, 24(5):342–361, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=685258.

Franklin:1996:DIS

[FR96] M. K. Franklin and M. K.

Reiter. The design and implementation of a secure auction service. IEEE Transactions on Software Engineering, 22(5):302–312, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=502223.

France:1992:SED

[Fra92] R. B. France. Semantically extended dataflow diagrams: a formal specification tool. IEEE Transactions on Software Engineering, 18(4):329–346, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=129221.

Freedman:1991:TSC

[Fre91] R. S. Freedman. Testability of software components. IEEE Transactions on Software Engineering, 17(6):553–564, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87281.

Fuggetta:1998:ISS

[FT98] A. Fuggetta and R. N. Taylor. Introduction to the special section. IEEE Transactions on Software Engineering, 24(8):585, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707694.

Ferrucci:1996:SBI

[FTTV96] F. Ferrucci, G. Tortora, M. Tucci, and G. Vitiello. Semanticsbased inference algorithms for adaptive visual environments. IEEE Transactions on Software Engineering, 22 (10):730–750, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=544351.

Fujiwara:1991:TSB

[FvBK^{+}91] S. Fujiwara, G. v. Bochmann, F. Khendek, M. Amalou, and A. Ghedamsi. Test selection based on finite state models. IEEE Transactions on Software Engineering, 17(6):591–603, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87284.

Frankl:1993:ECE

[FW93a] P. G. Frankl and S. N. Weiss.

An experimental comparison of the effectiveness of branch testing and data flow testing. IEEE Transactions on Software Engineering, 19(8):774–787, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238581.

Frankl:1993:FAF

[FW93b] P. G. Frankl and E. J.

Weyuker. A formal analysis of the fault-detecting ability of testing methods. IEEE Transactions on Software Engineering, 19(3):202–213, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221133.

Frankl:1993:PIB

[FW93c] P. G. Frankl and E. J. Weyuker. Provable improvements on branch testing. IEEE Transactions on Software Engineering, 19(10):962–975, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=245738.

Gaeta:1996:EDE

[Gae96] R. Gaeta. Efficient discreteevent simulation of colored Petri nets. IEEE Transactions on Software Engineering, 22(9):629–639, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541434.

Gupta:1991:PAT

[GAF91] A. Gupta, I. F. Akyildiz, and R. M. Fujimoto. Performance analysis of Time Warp with multiple homogeneous processors. IEEE Transactions on Software Engineering, 17(10):1013–1027, October 1991. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99190.

Gait:1991:SAR

[Gai91] J. Gait. Stability, availability, and response in network file service. IEEE Transactions on Software Engineering, 17(2):133–140, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67594.

Garlan:1998:GEI

[Gar98] D. Garlan. Guest editorial. Introduction to the special section. IEEE Transactions on Software Engineering, 24(7):497, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=708565.

Gorla:1990:DEE

[GBB90] N. Gorla, A. C. Benander, and B. A. Benander. Debugging effort estimation using software metrics. IEEE Transactions on Software Engineering, 16(2):223–231, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44385.

Griswold:1998:TSP

[GCB^{+}98] W. G. Griswold, M. I. Chen, R. W. Bowdidge, J. L. Cabaniss, V. B. Nguyen, and J. D. Morgenthaler. Tool support for planning the restructuring of data abstractions in large systems. IEEE Transactions on Software Engineering, 24(7):534–558, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=708568.

Gerrard:1990:FSD

[GCG90] C. P. Gerrard, D. Coleman, and R. M. Gallimore. Formal specification and design time testing. IEEE Transactions on Software Engineering, 16(1):1–12, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44359.

Graefe:1993:EPA

[GD93] G. Graefe and D. L. Davison. Encapsulation of parallelism and architecture-independence in extensible database query execution. IEEE Transactions on Software Engineering, 19(8):749–764, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238579.

Granda:1992:PEP

[GDG92] M. Granda, J. M. Drake, and J. A. Gregorio. Performance evaluation of parallel systems by using unbounded generalized stochastic Petri nets. IEEE Transactions on Software Engineering, 18(1):55–71, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120316.

Grassi:1992:OCC

[GDT92] V. Grassi, L. Donatiello, and S. Tucci. On the optimal checkpointing of critical tasks and transaction-oriented systems. IEEE Transactions on Software Engineering, 18(1):72–77, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120317.

Garland:1990:DLS

[GGH90] S. J. Garland, J. V. Guttag, and J. J. Horning. Debugging larch shared language specifications. IEEE Transactions on Software Engineering, 16(9):1044–1057, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58789.

Gallivan:1990:ECB

[GGJ^{+}90] K. Gallivan, D. Gannon, W. Jalby, A. Malony, and H. Wijshoff. Experimentally characterizing the behavior of multiprocessor memory systems: a case study. IEEE Transactions on Software Engineering, 16(2):216–223, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44384.

Gouda:1991:AP

[GH91] M. G. Gouda and T. Herman. Adaptive programming. IEEE Transactions on Software Engineering, 17(9):911–921, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92911.

Gerber:1995:CRT

[GH95] R. Gerber and Seongsoo Hong. Compiling real-time programs with timing constraint refinement and structural code motion. IEEE Transactions on Software Engineering, 21(5):389–404, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=387469.

Grundy:1998:IMM

[GHM98] J. Grundy, J. Hosking, and W. B. Mugridge. Inconsistency management for multipleview software development environments. IEEE Transactions on Software Engineering, 24(11):960–981, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=730545.

Gerber:1995:GRT

[GHS95] R. Gerber, Seongsoo Hong, and M. Saksena. Guaranteeing real-time requirements with resource-based calibration of periodic processes. IEEE Transactions on Software Engineering, 21(7):579–592, July 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=392979.

Girgis:1993:CCB

[Gir93] M. R. Girgis. Corrigendum for “Constraint-based automatic test data generation” by R. A. DeMillo and A. J. Offutt. IEEE Transactions on Software Engineering, 19(6):640–??, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232028.

See [DO91].

Gupta:1996:FFL

[GJB96] D. Gupta, P. Jalote, and G. Barua. A formal framework for on-line software version change. IEEE Transactions on Software Engineering, 22(2):120–131, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=485222.

Gill:1991:CCD

[GK91] G. K. Gill and C. F. Kemerer. Cyclomatic complexity density and software maintenance productivity. IEEE Transactions on Software Engineering, 17(12):1284–1288, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=106988.

Gansner:1993:TDD

[GKNV93] E. R. Gansner, E. Koutsofios, S. C. North, and K. P. Vo. A technique for drawing directed graphs. IEEE Transactions on Software Engineering, 19(3):214–230, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=221135.

Gallagher:1991:UPS

[GL91] K. B. Gallagher and J. R. Lyle. Using program slicing in software maintenance. IEEE Transactions on Software Engineering, 17(8):751–761, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83912.

Gerber:1992:LAA

[GL92] R. Gerber and I. Lee. A layered approach to automating the verification of realtime systems. IEEE Transactions on Software Engineering, 18(9):768–784, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=159838.

Ghezzi:1991:UHL

[GMMP91] C. Ghezzi, D. Mandrioli, S. Morasca, and M. Pezze. A unified high-level Petri net formalism for time-critical systems. IEEE Transactions on Software Engineering, 17 (2):160–172, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67597.

Guaspari:1990:FVA

[GMP90] D. Guaspari, C. Marceau, and W. Polak. Formal verification of Ada programs. IEEE Transactions on Software Engineering, 16(9):1058–1075, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58790.

Griswold:1995:ATM

[GN95] W. G. Griswold and D. Notkin.

Architectural tradeoffs for a meaning-preserving program restructuring tool. IEEE Transactions on Software Engineering, 21(4):275–287, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=385967.

Gallagher:1997:ATD

[GN97] M. J. Gallagher and V. Lakshmi Narasimhan. ADTEST: a test data generation suite for Ada software systems. IEEE Transactions on Software Engineering, 23(8):473–484, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=624304.

Ghezzi:1999:GEI

[GN99] C. Ghezzi and B. Nuseibeh. Guest editorial: introduction to the special section. IEEE Transactions on Software Engineering, 25(6):782– 783, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824393.

Goldschlag:1990:MVC

[Gol90] D. M. Goldschlag. Mechanically verifying concurrent programs with the Boyer–Moore prover. IEEE Transactions on Software Engineering, 16 (9):1005–1023, September 1990.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58787.

Govindarajan:1993:EHF

[Gov93] R. Govindarajan. Exception handlers in functional programming languages. IEEE Transactions on Software Engineering, 19(8):826–834, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238585.

Godefroid:1996:UPO

[GPS96] P. Godefroid, D. Peled, and M. Staskauskas. Using partialorder methods in the formal validation of industrial concurrent programs. IEEE Transactions on Software Engineering, 22(7):496–507, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=538606.

Gong:1996:CIS

[GQ96] Li Gong and Xiaolei Qian.

Computational issues in secure interoperation. IEEE Transactions on Software Engineering, 22(1):43–52, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=481533.

Graefe:1991:HFM

[Gra91] G. Graefe. Heap-filter merge join: a new algorithm for joining medium-size inputs. IEEE Transactions on Software Engineering, 17(9):979–982, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92919.

Griswold:1994:CLD

[Gri94] W. G. Griswold. Comments on “Language design for program manipulation”. IEEE Transactions on Software Engineering, 20(3):218–219, March 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=268924.

See [MDC92].

Gaudiot:1990:DDP

[GS90a] J. L. Gaudiot and A. Sohn. Data-driven parallel production systems. IEEE Transactions on Software Engineering, 16(3):281–293, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=48936.

Gupta:1990:RSA

[GS90b] R. Gupta and M. L. Soffa. Region scheduling: an approach for detecting and redistributing parallelism. IEEE Transactions on Software Engineering, 16(4):421–431, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54294.

Grandi:1993:BAE

[GS93] F. Grandi and M. R. Scalas.

Block access estimation for clustered data using a finite LRU buffer. IEEE Transactions on Software Engineering, 19(7):641–660, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238566.

Go:1999:DFS

[GS99] Kentaro Go and N. Shiratori. A decomposition of a formal specification: an improved constraint-oriented method. IEEE Transactions on Software Engineering, 25(2):258–273, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761449.

Goldman:1995:PPA

[GSM^{+}95] K. J. Goldman, B. Swaminathan, P. McCartney, M. D. Anderson, and R. Sethuraman. The Programmers’ Playground: I/O abstraction for user-configurable distributed applications. IEEE Transactions on Software Engineering, 21(9):735–746, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=464547.

Ghosal:1991:PWS

[GST91] D. Ghosal, G. Serazzi, and S. K. Tripathi. The processor working set and its use in scheduling multiprocessor systems. IEEE Transactions on Software Engineering, 17(5):443–453, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=90447.

Greif:1992:CSC

[GSW92] I. Greif, A. Seliger, and W. Weihl. A case study of CES: a distributed collaborative editing system implemented in Argus. IEEE Transactions on Software Engineering, 18(9):827–839, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=159831.

Gutjahr:1995:OTD

[Gut95] W. J. Gutjahr. Optimal test distributions for software failure cost estimation. IEEE Transactions on Software Engineering, 21(3):219–228, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372149.

Gutjahr:1999:PTV

[Gut99]  W. J. Gutjahr.  Partition

testing vs. random testing: the influence of uncertainty. IEEE Transactions on Software Engineering, 25(5):661– 674, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=815325.

Ganesan:1992:SMR

[GW92] R. Ganesan and S. Weiss.

Scalar memory references in pipelined multiprocessors: a performance study. IEEE Transactions on Software Engineering, 18(1):78–86, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120318.

Haas:1999:EMN

[Haa99] P. J. Haas. Estimation methods for nonregenerative stochastic Petri nets. IEEE Transactions on Software Engineering, 25(2):218–236, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761447.

Hall:1995:CPR

[Hal95] R. J. Hall. Call path refinement profiles. IEEE Transactions on Software Engineering, 21(6):481–496, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=391375.

Han:1995:PPL

[Han95]  Jia Liang Han.  Program

partition and logic program analysis. IEEE Transactions on Software Engineering, 21(12):959–968, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489072.

Harrison:1992:EBM

[Har92] W. Harrison. An entropybased measure of software complexity. IEEE Transactions on Software Engineering, 18(11):1025–1029, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=177371.

Hsu:1991:IRM

[HBRY91] C. Hsu, M. Bouziane, L. Rattner, and L. Yee. Information resources management in heterogeneous, distributed environments: A metadatabase approach. IEEE Transactions on Software Engineering, 17(6):604–625, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87285.

Hildum:1990:LSP

[HC90] D. Hildum and J. Cohen. A language for specifying program transformations. IEEE Transactions on Software Engineering, 16(6):630–638, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=55091.

Harrison:1998:EMS

[HCN98] R. Harrison, S. J. Counsell, and R. V. Nithi. An evaluation of the MOOD set of objectoriented software metrics. IEEE Transactions on Software Engineering, 24(6):491–496, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=689404.

Halici:1991:OLT

[HD91] U. Halici and A. Dogac. An optimistic locking technique for concurrency control in distributed databases. IEEE Transactions on Software Engineering, 17(7):712–724, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83907.

Hanna:1990:SVU

[HDL90] F. K. Hanna, N. Daeche, and M. Longley. Specification and verification using dependent types. IEEE Transactions on Software Engineering, 16(9):949–964, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58783.

Hellerstein:1993:ASR

[Hel93] J. L. Hellerstein. Achieving service rate objectives with decay usage scheduling. IEEE Transactions on Software Engineering, 19(8):813–825, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238584.

Hemmendinger:1990:SAS

[Hem90] D. Hemmendinger. Specifying Ada server tasks with executable formal grammars. IEEE Transactions on Software Engineering, 16(7):741–754, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=56100.

Heindl:1999:FOA

[HG99] A. Heindl and R. German. A fourth-order algorithm with automatic stepsize control for the transient analysis of DSPNs. IEEE Transactions on Software Engineering, 25(2):194– 206, March/April 1999. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
761445.

Heileman:1992:GFC

[HGR92] G. L. Heileman, M. Georgiopoulos, and W. D. Roome. A general framework for concurrent simulation on neural network models. IEEE Transactions on Software Engineering, 18(7):551–562, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=148474.

Hilderman:1997:NRV

[HH97] R. J. Hilderman and H. J.

Hamilton. A note on regeneration with virtual copies. IEEE Transactions on Software Engineering, 23(1):56–59, January 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=581329.

Hirakawa:1990:CRP

[Hir90] Y. Hirakawa. Comments, with reply, on “On the projection method for protocol verification” by T.-Y. Cheung. IEEE Transactions on Software Engineering, 16(3):370–371, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=48933.

See [Che86].

Hahn:1991:TSK

[HJR91] U. Hahn, M. Jarke, and T. Rose. Teamwork support in a knowledge-based information systems environment. IEEE Transactions on Software Engineering, 17(5):467–482, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=90449.

Harbour:1994:TAF

[HKL94] M. G. Harbour, M. H. Klein, and J. P. Lehoczky. Timing analysis for fixed-priority scheduling of hard real-time systems. IEEE Transactions on Software Engineering, 20(1):13–28, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=263752.

Heitmeyer:1998:UAM

[HKL^{+}98] C. Heitmeyer, J. Kirby, B. Labaw, M. Archer, and R. Bharadwaj. Using abstraction and model checking to detect safety violations in requirements specifications. IEEE Transactions on Software Engineering, 24(11):927–948, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=730543.

Heering:1990:IGP

[HKR90] J. Heering, P. Klint, and J. Rekers. Incremental generation of parsers. IEEE Transactions on Software Engineering, 16(12):1344–1351, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62443.

Han:1991:EEM

[HL91] J. Han and L. Liu. Efficient evaluation of multiple linear recursions. IEEE Transactions on Software Engineering, 17(12):1241–1252, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106985.

Helman:1993:SFA

[HL93] P. Helman and G. Liepins.

Statistical foundations of audit trail analysis for the detection of computer misuse. IEEE Transactions on Software Engineering, 19(9):886–901, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241771.

Heimdahl:1996:CCH

[HL96] M. P. E. Heimdahl and N. G. Leveson. Completeness and consistency in hierarchical statebased requirements. IEEE Transactions on Software Engineering, 22(6):363–377, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=508311.

Harel:1990:SWE

[HLN^{+}90] D. Harel, H. Lachover, A. Naamad, A. Pnueli, M. Politi, R. Sherman, A. Shtull-Trauring, and M. Trakhtenbrot. STATE-MATE: a working environment for the development of complex reactive systems. IEEE Transactions on Software Engineering, 16(4):403–414, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54292.

Halbwachs:1992:PVR

[HLR92] N. Halbwachs, F. Lagnier, and C. Ratel. Programming and verifying real-time systems by means of the synchronous dataflow language LUSTRE. IEEE Transactions on Software Engineering, 18(9):785–793, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=159839.

Harrold:1993:UIP

[HM93] M. J. Harrold and B. Malloy.

A unified interprocedural program representation for a maintenance environment. IEEE Transactions on Software Engineering, 19(6):584–593, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232023.

Hariri:1995:HMA

[HM95] S. Hariri and H. Mutlu. Hierarchical modeling of availability in distributed systems. IEEE Transactions on Software Engineering, 21(1):50–56, January 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=341847.

Hitz:1996:CKM

[HM96] M. Hitz and B. Montazeri. Chidamber and Kemerer’s metrics suite: a measurement theory perspective. IEEE Transactions on Software Engineering, 22(4):267–271, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=491650.

Harrold:1999:GEI

[HM99] M. J. Harrold and H. A.

Muller. Guest editorial: introduction to the special section. IEEE Transactions on Software Engineering, 25(6): 747–748, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824362.

Hurfin:1998:EDD

[HMRS98] M. Hurfin, M. Mizuno, M. Raynal, and M. Singhal. Efficient distributed detection of conjunctions of local predicates. IEEE Transactions on Software Engineering, 24(8):664–677, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707701.

Heydon:1990:MVS

[HMT^{+}90] A. Heydon, M. W. Maimone, J. D. Tygar, J. M. Wing, and A. M. Zaremski. Miro: visual specification of security. IEEE Transactions on Software Engineering, 16(10):1185–1197, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60298.

Helary:1999:CID

[HNR99] J. M. Helary, R. H. B. Netzer, and M. Raynal. Consistency issues in distributed checkpoints. IEEE Transactions on Software Engineering, 25(2):274–281, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761450.

Honiden:1994:AAI

[HNUI94] S. Honiden, K. Nishimura, N. Uchihira, and K. Itoh. An application of artificial intelligence to object-oriented performance design for realtime systems. IEEE Transactions on Software Engineering, 20(11):849–867, November 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=368123.

Hailpern:1990:EOS

[HO90] B. Hailpern and H. Ossher.

Extending objects to support multiple interfaces and access control. IEEE Transactions on Software Engineering, 16(11):1247–1257, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60313.

Hoffman:1990:CMI

[Hof90]  D. Hoffman.  On criteria

for module interfaces. IEEE Transactions on Software Engineering, 16(5):537–542, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=52776.

See comments [Con90].

Holzmann:1997:MCS

[Hol97] G. J. Holzmann. The model checker SPIN. IEEE Transactions on Software Engineering, 23(5):279–295, May 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=588521.

See comments [BCY01].

Houstics:1990:MAR

[Hou90] C. E. Houstics. Module allocation of real-time applications to distributed systems. IEEE Transactions on Software Engineering, 16(7):699–709, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=56096.

Howden:1990:CAP

[How90] W. E. Howden. Comments analysis and programming errors. IEEE Transactions on Software Engineering, 16(1):72–81, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=44365.

Hutchinson:1991:EEK

[HP91] N. C. Hutchinson and L. L. Peterson. The x-Kernel: an architecture for implementing network protocols. IEEE Transactions on Software Engineering, 17(1):64–76, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67579.

Holtzblatt:1997:DRD

[HPR^{+}97] L. J. Holtzblatt, R. L. Piazza, H. B. Reubenstein, S. N. Roberts, and D. R. Harris. Design recovery for distributed systems. IEEE Transactions on Software Engineering, 23(7):461–472, July 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=605763.

Haumer:1998:REV

[HPW98] P. Haumer, K. Pohl, and K. Weidenhaupt. Requirements elicitation and validation with real world scenes. IEEE Transactions on Software Engineering, 24(12):1036–1054, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738338.

Hatton:1994:HAS

[HR94a] L. Hatton and A. Roberts.

How accurate is scientific software? IEEE Transactions on Software Engineering, 20(10):785–797, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=328993.

Huang:1994:DAS

[HR94b] Yen-Min Huang and C. V. Ravishankar. Designing an agent synthesis system for cross-RPC communication. IEEE Transactions on Software Engineering, 20(3):188–198, March 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=268920.

Harrold:1996:SCA

[HR96] M. J. Harrold and G. Rothermel. Separate computation of alias information for reuse. IEEE Transactions on Software Engineering, 22(7):442–460, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=538603.

Huang:1998:CPS

[HR98] Y. M. Huang and C. V. Ravishankar. Constructive protocol specification using Cicero. IEEE Transactions on Software Engineering, 24(4):252–267, April 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=677183.

Hayes-Roth:1995:DSS

[HRPL^{+}95] B. Hayes-Roth, K. Pfleger, P. Lalanda, P. Morignot, and M. Balabanovic. A domainspecific software architecture for adaptive intelligent systems. IEEE Transactions on Software Engineering, 21(4):288–301, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=385968.

Hansen:1998:SAS

[HRS98] K. M. Hansen, A. P. Ravn, and V. Stavridou. From safety analysis to software requirements. IEEE Transactions on Software Engineering, 24(7):573–584, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=708570.

Haban:1990:ART

[HS90] D. Haban and K. G. Shin. Application of real-time monitoring to scheduling tasks with random execution times. IEEE Transactions on Software Engineering, 16(12):1374–1389, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62446.

Hoffman:1991:AMT

[HS91a] D. M. Hoffman and P. Strooper.

Automated module testing in Prolog. IEEE Transactions on Software Engineering, 17 (9):934–943, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92913.

Humphrey:1991:PIS

[HS91b] W. S. Humphrey and N. D. Singpurwalla. Predicting (individual) software productivity. IEEE Transactions on Software Engineering, 17 (2):196–207, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=67600.

Hamidzadeh:1993:SAR

[HS93] B. Hamidzadeh and S. Shekhar. Specification and analysis of real-time problem solvers. IEEE Transactions on Software Engineering, 19(8):788–803, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238582.

Henderson-Sellers:1997:CSS

[HS97] B. Henderson-Sellers. Corrigenda: “Software size estimation of object-oriented systems”. IEEE Transactions on Software Engineering, 23(4):260–261, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=588544.

See [Lar90].

Hamlet:1990:PTD

[HT90] D. Hamlet and R. Taylor. Partition testing does not inspire confidence [program testing]. IEEE Transactions on Software Engineering, 16(12):1402–1411, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62448.

Henderson:1991:EPS

[HT91] W. Henderson and P. G. Taylor. Embedded processes in stochastic Petri nets. IEEE Transactions on Software Engineering, 17(2):108–116, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67592.

Huang:1993:RAP

[HT93] Y. Huang and S. K. Tripathi. Resource allocation for primary-site fault-tolerant systems. IEEE Transactions on Software Engineering, 19(2):108–119, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214829.

Heintze:1996:MSP

[HT96] N. Heintze and J. D. Tygar.

A model for secure protocols and their compositions. IEEE Transactions on Software Engineering, 22(1):16–30, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=481514.

Hirakawa:1990:IPS

[HTI90] M. Hirakawa, M. Tanaka, and T. Ichikawa. An iconic programming system, HI-VISUAL. IEEE Transactions on Software Engineering, 16(10):1178– 1184, October 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
60297.

Hu:1997:EAS

[Hu97] Qing Hu. Evaluating alternative software production functions. IEEE Transactions on Software Engineering, 23(6):379–387, June 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=601078.

See comments [PKJH99].

Huang:1990:SCP

[Hua90]  J. C. Huang.  State con-

straints and pathwise decomposition of programs. IEEE Transactions on Software Engineering, 16(8):880–896, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57625.

Higashino:1994:AAT

[HvB94] T. Higashino and G. v. Bochmann.

Automatic analysis and test case derivation for a restricted class of LOTOS expressions with data parameters. IEEE Transactions on Software Engineering, 20(1):29–42, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=263753.

Howes:1989:MAO

[HW89] N. R. Howes and A. C. Weaver.

Measurements of Ada overhead in OSI-style communications systems. IEEE Transactions on Software Engineering, 15(12):1507–1517, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58763.

See comments [Kar90a].

Haban:1990:HMB

[HW90] D. Haban and D. Wybranietz.

A hybrid monitor for behavior and performance analysis of distributed systems. IEEE Transactions on Software Engineering, 16(2):197–211, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44382.

Harms:1991:CSI

[HW91] D. E. Harms and B. W. Weide. Copying and swapping: influences on the design of reusable software components. IEEE Transactions on Software Engineering, 17(5):424–435, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=90445.

Howden:1994:QMS

[HW94] W. E. Howden and B. Wieand. QDA — a method for systematic informal program analysis. IEEE Transactions on Software Engineering, 20(6):445–462, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=295893.

Hrischuk:1999:TBL

[HWR99] C. E. Hrischuk, C. Murray Woodside, and J. A. Rolia. Trace-based load characterization for generating performance software models. IEEE Transactions on Software Engineering, 25(1):122– 135, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=748921.

Hatono:1991:MOS

[HYT91] I. Hatono, K. Yamagata, and H. Tamura. Modeling and online scheduling of flexible manufacturing systems using stochastic Petri nets. IEEE Transactions on Software Engineering, 17(2):126–132, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67588.

Helander:1998:PMS

[HZO98] M. E. Helander, Ming Zhao, and N. Ohlsson. Planning models for software reliability and cost. IEEE Transactions on Software Engineering, 24(6):420–434, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=689400.

Isakowitz:1996:SSR

[IK96] T. Isakowitz and R. J. Kauffman. Supporting search for reusable software objects. IEEE Transactions on Software Engineering, 22(6):407–423, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=508314.

Ibaraki:1990:MCS

[IKK90] T. Ibaraki, T. Kameda, and N. Katoh. Multiversion cautious schedulers for database concurrency control. IEEE Transactions on Software Engineering, 16(3):302–315, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=48938.

Ilgun:1995:STA

[IKP95] K. Ilgun, R. A. Kemmerer, and P. A. Porras. State transition analysis: a rule-based intrusion detection approach. IEEE Transactions on Software Engineering, 21(3):181–199, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=372146.

Itai:1990:ODR

[IKWZ90] A. Itai, S. Kutten, Y. Wolfstahl, and S. Zaks. Optimal distributed t-resilient election in complete networks. IEEE Transactions on Software Engineering, 16(4):415–420, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=54293.

Iazeolla:1993:LMT

[IM93] G. Iazeolla and F. Marinuzzi. LISPACK — a methodology and tool for the performance analysis of parallel systems and algorithms. IEEE Transactions on Software Engineering, 19(5):486–502, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232014.

Inverardi:1995:FSA

[IW95] P. Inverardi and A. L. Wolf.

Formal specification and analysis of software architectures using the chemical abstract machine model. IEEE Transactions on Software Engineering, 21(4):373–386, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=385973.

Jacob:1991:UPC

[Jac91] J. Jacob. A uniform presentation of confidentiality properties. IEEE Transactions on Software Engineering, 17 (11):1186–1194, November 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106973.

Jacky:1995:SSC

[Jac95]  J. Jacky.  Specifying a

safety-critical control system in z. IEEE Transactions on Software Engineering, 21 (2):99–106, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345826.

Jarzabek:1998:DFS

[Jar98] S. Jarzabek. Design of flexible static program analyzers with PQL. IEEE Transactions on Software Engineering, 24(3):197–215, March 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=667879.

Jazayeri:1998:GEI

[Jaz98] M. Jazayeri. Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 24(9):681, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=713321.

Jaccheri:1993:TPM

[JC93] M. L. Jaccheri and R. Conradi. Techniques for process model evolution in EPOS. IEEE Transactions on Software Engineering, 19(12):1145–1156, December 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=249660.

Jackson:1996:ESA

[JD96] D. Jackson and C. A. Damon. Elements of style: analyzing a software design feature with a counterexample detector. IEEE Transactions on Software Engineering, 22(7):484–495, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=538605.

Johnson:1992:RPR

[JFH92] W. L. Johnson, M. S. Feather, and D. R. Harris. Representation and presentation of requirements knowledge. IEEE Transactions on Software Engineering, 18(10):853–869, October 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=163603.

Jarke:1998:ISI

[JKS98] M. Jarke and R. Kurki-Suonio. Introduction to the special issue. IEEE Transactions on Software Engineering, 24(12):1033–1035, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738337.

Jeffery:1993:CFP

[JLB93] D. R. Jeffery, G. C. Low, and M. Barnes. A comparison of function point counting techniques. IEEE Transactions on Software Engineering, 19(5):529–532, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232016.

Jaffe:1991:SRA

[JLHM91] M. S. Jaffe, N. G. Leveson, M. P. E. Heimdahl, and B. E. Melhart. Software requirements analysis for real-time process-control systems. IEEE Transactions on Software Engineering, 17(3):241–258, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=75414.

Jahanian:1994:MSL

[JM94] F. Jahanian and A. K. Mok. Modechart: a specification language for real-time systems. IEEE Transactions on Software Engineering, 20 (12):933–947, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368134.

Jonsson:1997:QMS

[JO97] E. Jonsson and T. Olovsson. A quantitative model of the security intrusion process based on attacker behavior. IEEE Transactions on Software Engineering, 23(4):235–245, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=588541.

Johnson:1995:AAR

[Joh95] T. Johnson. Approximate analysis of reader/writer queues.

IEEE Transactions on Software Engineering, 21(3):209–218, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372148.

Jorgensen:1995:EAS

[Jor95]  M. Jorgensen.  Experience

with the accuracy of software maintenance task effort prediction models. IEEE Transactions on Software Engineering, 21(8):674–681, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=403791.

Joshi:1999:ARC

[JR99] R. K. Joshi and D. J. Ram.

Anonymous remote computing: a paradigm for parallel programming on interconnected workstations. IEEE Transactions on Software Engineering, 25(1):75–90, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=748919.

Jeffery:1999:ESE

[JV99] D. R. Jeffery and L. G.

Votta. Empirical software engineering. IEEE Transactions on Software Engineering, 25(4):435–437, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=799935.

Jackson:1998:DFC

[JZ98] M. Jackson and P. Zave. Distributed feature composition: a virtual architecture for telecommunications services. IEEE Transactions on Software Engineering, 24(10):831–847, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=729683.

Kailar:1996:AEC

[Kai96] R. Kailar. Accountability in electronic commerce protocols. IEEE Transactions on Software Engineering, 22(5):313–328, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=502224.

Karam:1990:CMA

[Kar90a] G. M. Karam. Comments on “Measurements of Ada overhead in OSI-style communications systems”. IEEE Transactions on Software Engineering, 16(12):1435–1439, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=62440.

See [HW89].

Karam:1990:CIP

[Kar90b] G. M. Karam. Computational improvements in Prolog applications by predicate variable pointers. IEEE Transactions on Software Engineering, 16(5):490–497, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=52772.

Katcher:1993:EAF

[KAS93] D. I. Katcher, H. Arakawa, and J. K. Strosnider. Engineering and analysis of fixed priority schedulers. IEEE Transactions on Software Engineering, 19(9):920–934, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241774.

Kaudel:1990:CAP

[Kau90]  F. Kaudel.  Comments on

“Allocating programs containing branches and loops within a multiple processor system” by D. Towsley. IEEE Transactions on Software Engineering, 16(4):471–??, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54298.

See [Tow86].

Karam:1990:SCR

[KB90] G. M. Karam and R. J. A. Buhr. Starvation and critical race analyzers for Ada. IEEE Transactions on Software Engineering, 16(8):829–843, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=57622.

Karam:1991:TLB

[KB91] G. M. Karam and R. J. A. Buhr. Temporal logic-based deadlock analysis for Ada. IEEE Transactions on Software Engineering, 17(10):1109–1125, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=99197.

See comments [YLT93].

Kunz:1995:UAP

[KB95] T. Kunz and J. P. Black. Using automatic process clustering for design recovery and distributed debugging. IEEE Transactions on Software Engineering, 21(6):515–527, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=391378.

Kanoun:1991:MSR

[KdBMdS91] K. Kanoun, M. R. de Bastos Martini, and J. M. de Souza. A method for software reliability analysis and prediction application to the TROPICO-R switching system. IEEE Transactions on Software Engineering, 17(4):334–344, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90433.

Kemper:1996:NAS

[Kem96] P. Kemper. Numerical analysis of superposed GSPNs. IEEE Transactions on Software Engineering, 22(9):615–628, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541433.

Kemper:1999:TAS

[Kem99] P. Kemper. Transient analysis of superposed GSPNs. IEEE Transactions on Software Engineering, 25(2):182–193, March/ April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761444.

Kinsley:1992:AVM

[KH92a] K. C. Kinsley and C. E.

Hughes. Analysis of a virtual memory model for maintaining database views. IEEE Transactions on Software Engineering, 18(5):402–409, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=135773.

Kleinrock:1992:PPS

[KH92b] L. Kleinrock and J. H. Huang. On parallel processing systems: Amdahl’s law generalized and some results on optimal design. IEEE Transactions on Software Engineering, 18(5):434–447, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=135776.

Kalbarczyk:1999:HSA

[KIR^{+}99] Z. Kalbarczyk, R. K. Iyer, G. L. Ries, J. U. Patel, M. S. Lee, and Y. Xiao. Hierarchical simulation approach to accurate fault modeling for system dependability evaluation. IEEE Transactions on Software Engineering, 25(5): 619–632, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=815322.

Kao:1993:FFI

[KIT93] W. I. Kao, R. K. Iyer, and D. Tang. FINE: A fault injection and monitoring environment for tracing the UNIX system behavior under faults. IEEE Transactions on Software Engineering, 19(11):1105–1118, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256857.

Kitchenham:1998:PAU

[Kit98] B. Kitchenham. A procedure for analyzing unbalanced datasets. IEEE Transactions on Software Engineering, 24(4):278–301, April 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=677185.

Keck:1998:FSI

[KK98] D. O. Keck and P. J. Kuehn.

The feature and service interaction problem in telecommunications systems: a survey. IEEE Transactions on Software Engineering, 24(10):779–796, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=729680.

Krishnan:1999:MPC

[KK99] M. S. Krishnan and M. I.

Kellner. Measuring process consistency: implications for reducing software defects. IEEE Transactions on Software Engineering, 25(6): 800–815, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824401.

Kanoun:1994:SRT

[KL94] K. Kanoun and J. C. Laprie. Software reliability trend analyses from theoretical to practical considerations. IEEE Transactions on Software Engineering, 20(9):740–747, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=317434.

Kramer:1993:CSR

[KLB93] B. Kramer, Luqi, and V. Berzins. Compositional semantics of a real-time prototyping language. IEEE Transactions on Software Engineering, 19(5):453–477, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=232012.

Kelly:1990:ADT

[KM90a] J. P. J. Kelly and S. C. Murphy. Achieving dependability throughout the development process: a distributed software experiment. IEEE Transactions on Software Engineering, 16(2):153–165, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=44379.

Kramer:1990:EPP

[KM90b] J. Kramer and J. Magee. The evolving philosophers problem: dynamic change management. IEEE Transactions on Software Engineering, 16(11):1293–1306, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60317.

Kato:1992:PCI

[KM92] K. Kato and T. Masuda. Persistent caching: An implementation technique for complex objects with object identity. IEEE Transactions on Software Engineering, 18(7):631–645, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=148481.

Kaufmann:1997:IST

[KM97] M. Kaufmann and J. S. Moore.

An industrial strength theorem prover for a logic based on Common Lisp. IEEE Transactions on Software Engineering, 23(4):203–213, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588534.

Karamanolis:1999:CAP

[KM99] C. T. Karamanolis and J. N.

Magee. Client-access protocols for replicated services. IEEE Transactions on Software Engineering, 25(1):3– 21, January/February 1999.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=748915.

Khoshgoftaar:1992:PMT

[KMBR92] T. M. Khoshgoftaar, J. C. Munson, B. B. Bhattacharya, and G. D. Richardson. Predictive modeling techniques of software quality from software measures. IEEE Transactions on Software Engineering, 18(11):979–987, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=177367.

Kang:1996:NP

[KML96] M. H. Kang, I. S. Moskowitz, and D. C. Lee. A network pump. IEEE Transactions on Software Engineering, 22(5):329–338, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=502225.

Krauser:1991:HPS

[KMR91] E. W. Krauser, A. P. Mathur, and V. J. Rego. High performance software testing on SIMD machines. IEEE Transactions on Software Engineering, 17(5):403–423, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90444.

Kozaczynski:1992:PCR

[KNE92] W. Kozaczynski, J. Ning, and A. Engberts. Program concept recognition and transformation. IEEE Transactions on Software Engineering, 18(12):1065–1075, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=184761.

Korel:1990:AST

[Kor90] B. Korel. Automated software test data generation. IEEE Transactions on Software Engineering, 16(8):870–879, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57624.

Korel:1997:CDP

[Kor97] B. Korel. Computation of dynamic program slices for unstructured programs. IEEE Transactions on Software Engineering, 23(1):17–34, January 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=581327.

Kemmerer:1991:CFT

[KP91] R. A. Kemmerer and P. A. Porras. Covert flow trees: a visual approach to analyzing covert storage channels. IEEE Transactions on Software Engineering, 17(11):1166–1185, November 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106972.

Kemerer:1992:IRF

[KP92] C. F. Kemerer and B. S. Porter. Improving the reliability of function point measurement: an empirical study. IEEE Transactions on Software Engineering, 18(11):1011–1024, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177370.

Kitchenham:1995:TFS

[KPF95] B. Kitchenham, S. Lawrence Pfleeger, and N. Fenton. Towards a framework for software measurement validation. IEEE Transactions on Software Engineering, 21(12):929–944, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=489070.

See comments [MBB^{+}97] and reply [KPF97].

Kitchenham:1997:RCT

[KPF97] B. Kitchenham, S. Lawrence Pfleeger, and N. Fenton. Reply to: Comments on “Towards a Framework of Software Measurement Validation”.

IEEE Transactions on Software Engineering, 23(3):189, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585507.

See [MBB^{+}97].

Kay:1993:RGM

[KR93] A. Kay and J. N. Reed. A rely and guarantee method for timed CSP: a specification and design of a telephone exchange. IEEE Transactions on Software Engineering, 19(6):625–639, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232027.

Krishnamurthy:1995:YGP

[KR95] B. Krishnamurthy and D. S.

Rosenblum. Yeast: a general purpose event-action system. IEEE Transactions on Software Engineering, 21 (10):845–857, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=469456.

Kshemkalyani:1991:IBV

[KS91] A. D. Kshemkalyani and M. Singhal. Invariant-based verification of a distributed deadlock detection algorithm. IEEE Transactions on Software Engineering, 17(8):789–799, August 1991. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83914.

Kurki-Suonio:1993:SDR

[KS93] R. Kurki-Suonio. Stepwise design of real-time systems. IEEE Transactions on Software Engineering, 19(1):56–69, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=210307.

Kiskis:1994:SSW

[KS94a] D. L. Kiskis and K. G. Shin. SWSL: a synthetic workload specification language for realtime systems. IEEE Transactions on Software Engineering, 20(10):798–811, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=328992.

Krueger:1994:ALP

[KS94b] P. Krueger and N. G. Shivaratri. Adaptive location policies for global scheduling. IEEE Transactions on Software Engineering, 20(6):432–444, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=295892.

Kshemkalyani:1994:EDR

[KS94c] A. D. Kshemkalyani and M. Singhal. Efficient detection and resolution of generalized distributed deadlocks. IEEE Transactions on Software Engineering, 20(1):43–54, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=263754.

Kemerer:1999:EAS

[KS99a] C. F. Kemerer and S. Slaughter.

An empirical approach to studying software evolution. IEEE Transactions on Software Engineering, 25(4):493–509, July/ August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799945.

Klarlund:1999:DSL

[KS99b] N. Klarlund and M. I. Schwartzbach.

A domain-specific language for regular sets of strings and trees. IEEE Transactions on Software Engineering, 25(3):378–386, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798326.

Karam:1989:CRA

[KSB89] G. M. Karam, C. M. Stanczyk, and G. W. Bond. Critical races in Ada programs. IEEE Transactions on Software Engineering, 15(11):1471–1480, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41338.

See comments [MO90].

Kim:1992:ATN

[KSH^{+}92] S. T. Kim, K. Suwunboriruksa, S. Herath, A. Jayasumana, and J. Herath. Algorithmic transformations for neural computing and performance of supervised learning on a dataflow machine. IEEE Transactions on Software Engineering, 18(7):613–623, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=148479.

Kunz:1991:IDW

[Kun91] T. Kunz. The influence of different workload descriptions on a heuristic load balancing scheme. IEEE Transactions on Software Engineering, 17(7):725–730, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83908.

Karunanithi:1992:PSR

[KWM92] N. Karunanithi, D. Whitley, and Y. K. Malaiya. Prediction of software reliability using connectionist models. IEEE Transactions on Software Engineering, 18(7):563–574, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=148475.

Karger:1991:RVV

[KZB^{+}91] P. A. Karger, M. E. Zurko, D. W. Bonin, A. H. Mason, and C. E. Kahn. A retrospective on the VAX VMM security kernel. IEEE Transactions on Software Engineering, 17(11):1147–1165, November 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106971.

Lamb:1990:SI

[Lam90] D. A. Lamb. Specification of iterators. IEEE Transactions on Software Engineering, 16(12):1352–1360, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62444.

Lamport:1995:TP

[Lam95] L. Lamport. TLA in pictures. IEEE Transactions on Software Engineering, 21 (9):768–775, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=464544.

Lang:1999:EBA

[Lan99] S. D. Lang. An extended banker’s algorithm for deadlock avoidance. IEEE Transactions on Software Engineering, 25(3):428–432, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798330.

Laranjeira:1990:SSE

[Lar90] L. A. Laranjeira. Software size estimation of object-oriented systems. IEEE Transactions on Software Engineering, 16(5):510–522, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=52774.

See corrigenda [HS97].

Lor:1991:ASS

[LB91] K. W. E. Lor and D. M. Berry. Automatic synthesis of SARA design models from system requirements. IEEE Transactions on Software Engineering, 17(12):1229–1240, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=106984.

Lopez-Benitez:1994:DMA

[LB94] N. Lopez-Benitez. Dependability modeling and analysis of distributed programs. IEEE Transactions on Software Engineering, 20(5):345–352, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=286421.

Lim:1995:AWC

[LBJ^{+}95] Sung-Soo Lim, Young Hyun Bae, Gyu Tae Jang, Byung-Do Rhee, Sang Lyul Min, Chang Yun Park, Heonshik Shin, Kunsoo Park, Soo-Mook Moon, and Chong Sang Kim. An accurate worst case timing analysis for RISC processors. IEEE Transactions on Software Engineering, 21(7):593–604, July 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=392980.

Liao:1992:SAH

[LC92] Y. Liao and D. Cohen. A specificational approach to high level program monitoring and measuring. IEEE Transactions on Software Engineering, 18(11):969–978, November 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=177366.

Lee:1998:IAU

[LCK98] Woo Jin Lee, Sung Deok Cha, and Yong Rae Kwon. Integration and analysis of use cases using modular Petri nets in requirements engineering. IEEE Transactions on Software Engineering, 24(12):1115–1130, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=738342.

Leveson:1990:USC

[LCKS90] N. G. Leveson, S. S. Cha, J. C. Knight, and T. J. Shimeall. The use of self checks and voting in software error detection: an empirical study. IEEE Transactions on Software Engineering, 16(4):432–443, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=54295.

Lunt:1990:SSM

[LDS^{+}90] T. F. Lunt, D. E. Denning, R. R. Schell, M. Heckman, and W. R. Shockley. The SeaView security model. IEEE Transactions on Software Engineering, 16(6):593–607, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=55088.

Luo:1994:STB

[LDvB94] G. Luo, A. Das, and G. v. Bochmann. Software testing based on SDL specifications with save. IEEE Transactions on Software Engineering, 20(1):72–87, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=263756.

LeMetayer:1998:DSA

[Le 98] D. Le Metayer. Describing software architecture styles using graph grammars. IEEE Transactions on Software Engineering, 24(7):521–533, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=708567.

Levendel:1990:RAL

[Lev90] Y. Levendel. Reliability analysis of large software systems: defect data modeling. IEEE Transactions on Software Engineering, 16(2):141–152, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44378.

Leite:1991:RVT

[LF91] J. C. S. P. Leite and P. A.

Freeman. Requirements validation through viewpoint resolution. IEEE Transactions on Software Engineering, 17 (12):1253–1269, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106986.

Leveson:1994:RSP

[LHHR94] N. G. Leveson, M. P. E. Heimdahl, H. Hildreth, and J. D. Reese. Requirements specification for process-control systems. IEEE Transactions on Software Engineering, 20(9):684–707, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=317428.

Lee:1995:SDT

[LI95] Inhwan Lee and R. K. Iyer. Software dependability in the Tandem GUARDIAN system. IEEE Transactions on Software Engineering, 21(5):455–467, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=387474.

Liu:1998:PAS

[Liu98] Zhen Liu. Performance analysis of stochastic timed Petri nets using linear programming approach. IEEE Transactions on Software Engineering, 24(11):1014–1030, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=730548.

Low:1990:FPE

[LJ90] G. C. Low and D. R. Jeffery. Function points in the estimation and evaluation of the software process. IEEE Transactions on Software Engineering, 16(1):64–71, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44364.

Lin:1998:MFIa

[LJ98] Yow-Jian Lin and M. Jazayeri.

Managing feature interactions telecommunications software systems — guest editorial. IEEE Transactions on Software Engineering, 24(10):777–778, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=729679.

Leong:1990:TPL

[LJS^{+}90] S. Leong, S. Jodis, K. Sullivan, O. Jiang, and P. A. D. de Maine. A transportable programming language (TPL) system. II. The bifunctional compiler system. IEEE Transactions on Software Engineering, 16(6):639–646, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=55092.

Lakshmanan:1991:PCF

[LJS91] K. B. Lakshmanan, S. Jayaprakash, and P. K. Sinha. Properties of control-flow complexity measures. IEEE Transactions on Software Engineering, 17(12):1289–1295, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=106989.

Laprie:1992:XWR

[LK92] J. C. Laprie and K. Kanoun. X-ware reliability and availability modeling. IEEE Transactions on Software Engineering, 18(2):130–147, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=121755.

Luckham:1995:SAS

[LKA^{+}95] D. C. Luckham, J. J. Kenney, L. M. Augustin, J. Vera, D. Bryan, and W. Mann. Specification and analysis of system architecture using Rapide. IEEE Transactions on Software Engineering, 21(4):336–354, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=385971.

Laprie:1991:KKA

[LKBK91] J. C. Laprie, K. Kanoun, C. Beounes, and M. Kaaniche. The KAT (knowledge–action– transformation) approach to the modeling and evaluation of reliability and availability growth. IEEE Transactions on Software Engineering, 17(4):370–382, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90436.

Lee:1997:ESS

[LL97] Nam-Yong Lee and C. R.

Litecky. An empirical study of software reuse with special attention to Ada. IEEE Transactions on Software Engineering, 23(9):537–549, September 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=629492.

Lin:1998:MFIb

[LLG98] F. J. Lin, Hong Liu, and A. Ghosh. A methodology for feature interaction detection in the AIN 0.1 framework. IEEE Transactions on Software Engineering, 24(10):797–817, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=729681.

Lang:1990:EEC

[LM90] S. D. Lang and Y. Manolopoulos. Efficient expressions for completely and partly unsuccessful batched search of treestructured files. IEEE Transactions on Software Engineering, 16(12):1433–1435, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=62451.

Levy:1995:EAB

[LM95] H. Levy and R. J. T. Morris. Exact analysis of Bernoulli superposition of streams into a least recently used cache. IEEE Transactions on Software Engineering, 21(8):682–688, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403792.

Levy:1996:CAP

[LMM96] H. Levy, T. G. Messinger, and R. J. T. Morris. The cache assignment problem and its application to database buffer management. IEEE Transactions on Software Engineering, 22(11):827–838, November 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=553701.

Lejter:1992:SMO

[LMR92] M. Lejter, S. Meyers, and S. P. Reiss. Support for maintaining object-oriented programs. IEEE Transactions on Software Engineering, 18(12):1045–1052, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=184759.

Liu:1998:CSF

[LOHS98a] S. Liu, A. J. Offutt, and C. Ho-Stuart. Correction to “SOFL: A formal engineering methodology for industrial applications”. IEEE Transactions on Software Engineering, 24(5):390, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=685261.

See [LOHS^{+}98b].

Liu:1998:SFE

[LOHS^{+}98b] Shaoying Liu, A. J. Offutt, C. Ho-Stuart, Y. Sun, and M. Ohba. SOFL: a formal engineering methodology for industrial applications. IEEE Transactions on Software Engineering, 24(1):24–45, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=663996.

See correction [LOHS98a].

Lecerof:1998:ASU

[LP98a] A. Lecerof and F. Paterno.

Automatic support for usability evaluation. IEEE Transactions on Software Engineering, 24(10):863–888, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=729686.

Lederer:1998:CMS

[LP98b] A. L. Lederer and J. Prasad. A causal model for software cost estimating error. IEEE Transactions on Software Engineering, 24(2):137–148, February 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=666827.

Lin:1992:DLB

[LR92] H. C. Lin and C. S. Raghavendra. A dynamic load-balancing policy with a central job dispatcher (LBC). IEEE Transactions on Software Engineering, 18(2):148–158, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=121756.

Ladd:1995:LIL

[LR95] D. A. Ladd and J. C. Ramming. A*: a language for implementing language processors. IEEE Transactions on Software Engineering, 21 (11):894–901, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=473218.

Lowe:1997:UCD

[LR97] G. Lowe and B. Roscoe. Using CSP to detect errors in the TMN protocol. IEEE Transactions on Software Engineering, 23(10):659–669, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637148.

Lee:1995:RAP

[LRF95] Yong-Fong Lee, B. G. Ryder, and M. E. Fiuczynski. Region analysis: a parallel elimination method for data flow analysis. IEEE Transactions on Software Engineering, 21(11):913–926, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=473220.

Litwin:1991:THC

[LRLH91] W. A. Litwin, N. Roussopoulos, G. Levy, and W. Hong. Trie hashing with controlled load. IEEE Transactions on Software Engineering, 17(7):678–691, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83904.

Lam:1990:RNS

[LS90a] S. S. Lam and A. U. Shankar.

A relational notation for state transition systems. IEEE Transactions on Software Engineering, 16(7):755–775, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=56101.

Liu:1990:SAR

[LS90b] L. Y. Liu and R. K. Shyamasundar. Static analysis of realtime distributed systems. IEEE Transactions on Software Engineering, 16(4):373–388, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=54290.

Lu:1990:MMS

[LS90c] W. P. Lu and M. K. Sundareshan. A model for multilevel security in computer networks. IEEE Transactions on Software Engineering, 16(6):647–659, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=55093.

Lee:1991:ELM

[LS91] S. Lee and S. Sluizer. An executable language for modeling simple behavior. IEEE Transactions on Software Engineering, 17(6):527–543, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=87279.

Lam:1994:TIM

[LS94] S. S. Lam and A. U. Shankar. A theory of interfaces and modules I-composition theorem. IEEE Transactions on Software Engineering, 20(1):55–71, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=263755.

Lortz:1995:SQP

[LS95] V. B. Lortz and K. G. Shin.

Semaphore queue priority assignment for real-time multiprocessor synchronization. IEEE Transactions on Software Engineering, 21(10):834–844, October 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=469457.

Lupu:1999:CPB

[LS99] E. C. Lupu and M. Sloman. Conflicts in policy-based distributed systems management. IEEE Transactions on Software Engineering, 25(6): 852–869, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824414.

Liu:1994:DAD

[LSB94] Ying Liu, A. K. Singh, and R. L. Bagrodia. A decompositional approach to the design of parallel programs. IEEE Transactions on Software Engineering, 20(12):914–932, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368135.

Lichter:1994:PIS

[LSHZ94] H. Lichter, M. Schneider-Hufschmidt, and H. Zullighoven. Prototyping in industrial software projects-bridging the gap between theory and practice. IEEE Transactions on Software Engineering, 20(11):825–832, November 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368126.

Li:1994:NSI

[LSY94] Jianan Li, I. Suzuki, and M. Yamashita. A new structural induction theorem for rings of temporal Petri nets. IEEE Transactions on Software Engineering, 20(2):115–126, February 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=265633.

Luqi:1990:GMS

[Luq90]  Luqi.  A graph model for

software evolution. IEEE Transactions on Software Engineering, 16(8):917–927, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=57627.

Lustman:1994:STB

[Lus94]  F. Lustman.  Specifying

transaction-based information systems with regular expressions. IEEE Transactions on Software Engineering, 20(3): 207–217, March 1994. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
268922.

Luckham:1995:EBA

[LV95] D. C. Luckham and J. Vera.

An event-based architecture definition language. IEEE Transactions on Software Engineering, 21(9):717–734, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=464548.

Lanubile:1997:ERF

[LV97] F. Lanubile and G. Visaggio. Extracting reusable functions by flow graph based program slicing. IEEE Transactions on Software Engineering, 23(4):246–259, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588543.

Luo:1994:TSB

[LvBP94] Gang Luo, G. von Bochmann, and A. Petrenko. Test selection based on communicating nondeterministic finite-state machines using a generalized wp-method. IEEE Transactions on Software Engineering, 20(2):149–162, February 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=265636.

Li:1990:CEP

[LW90] Guo-Jie Li and B. W. Wah. Computational efficiency of parallel combinatorial OR-tree searches. IEEE Transactions on Software Engineering, 16(1):13–31, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44360.

Lutz:1992:DUE

[LW92] R. R. Lutz and J. S. K. Wong. Detecting unsafe error recovery schedules. IEEE Transactions on Software Engineering, 18(8):749–760, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153384.

Littlewood:1997:SCS

[LW97] B. Littlewood and D. Wright. Some conservative stopping rules for the operational testing of safety critical software. IEEE Transactions on Software Engineering, 23 (11):673–683, November 1997.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637384.

Lieberherr:1993:OOS

[LX93] K. J. Lieberherr and C. Xiao.

Object-oriented software evolution. IEEE Transactions on Software Engineering, 19 (4):313–343, April 1993. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
223802.

Lee:1995:RM

[LZ95] A. H. Lee and J. L. Zachary.

Reflections on metaprogramming. IEEE Transactions on Software Engineering, 21 (11):883–893, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=473217.

Miyoshi:1993:ESE

[MA93] T. Miyoshi and M. Azuma.

An empirical study of evaluating software development environment quality. IEEE Transactions on Software Engineering, 19(5):425–435, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232010.

Madhavji:1992:EEP

[Mad92]  N. H. Madhavji.  Environ-

ment evolution: the Prism model of changes. IEEE Transactions on Software Engineering, 18(5):380–392, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=135771.

Morasca:1997:CTF

[MBB^{+}97] S. Morasca, L. C. Briand, V. R. Basili, E. J. Weyuker, M. V. Zelkowitz, B. Kitchenham, S. Lawrence Pfleeger, and N. Fenton. Comments on “Towards a framework for software measurement validation”. IEEE Transactions on Software Engineering, 23(3):187–189, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585506.

See [KPF95] and reply [KPF97].

Maarek:1991:IRA

[MBK91] Y. S. Maarek, D. M. Berry, and G. E. Kaiser. An information retrieval approach for automatically constructing software libraries. IEEE Transactions on Software Engineering, 17(8):800–813, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83915.

Matson:1994:SDC

[MBM94] J. E. Matson, B. E. Barrett, and J. M. Mellichamp. Software development cost estimation using function points. IEEE Transactions on Software Engineering, 20(4):275–287, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277575.

Mark:1992:GR

[MC92] L. Mark and R. J. Cochrane.

Grammars and relations. IEEE Transactions on Software Engineering, 18(9):840–849, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=159832.

Mok:1997:MMR

[MC97] A. K. Mok and D. Chen.

A multiframe model for realtime tasks. IEEE Transactions on Software Engineering, 23(10):635–645, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637146.

Mannino:1990:OOF

[MCB90] M. V. Mannino, J. J. Choi, and D. S. Batory. The object-oriented functional data language. IEEE Transactions on Software Engineering, 16(11):1258–1272, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=60314.

McCullough:1990:HTM

[McC90] D. McCullough. A hookup theorem for multilevel security. IEEE Transactions on Software Engineering, 16(6):563–568, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=55085.

McHugh:1996:ISS

[McH96] J. McHugh. Introduction to the special section. IEEE Transactions on Software Engineering, 22(1):3–??, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=481512.

McLean:1996:GTC

[McL96] J. McLean. A general theory of composition for a class of “possibilistic” properties. IEEE Transactions on Software Engineering, 22(1):53–67, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=481534.

Mylopoulos:1992:RUN

[MCN92] J. Mylopoulos, L. Chung, and B. Nixon. Representing and using nonfunctional requirements: a process-oriented approach. IEEE Transactions on Software Engineering, 18(6):483–497, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142871.

Merks:1992:LDP

[MDC92] E. A. T. Merks, J. M. Dyck, and R. D. Cameron. Language design for program manipulation. IEEE Transactions on Software Engineering, 18(1):19–32, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120313.

See comments [Gri94].

Meadows:1996:GEI

[Mea96] C. Meadows. Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 22(5):281–??, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=502220.

Miriyala:1991:ADF

[MH91] K. Miriyala and M. T. Harandi.

Automatic derivation of formal software specifications from informal descriptions. IEEE Transactions on Software Engineering, 17(10):1126–1142, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99198.

Mahony:1992:CST

[MH92] B. P. Mahony and I. J. Hayes. A case-study in timed refinement: a mine pump. IEEE Transactions on Software Engineering, 18(9):817–826, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=159841.

Milewski:1990:FDS

[Mil90] J. Milewski. Functional data structures as updatable objects. IEEE Transactions on Software Engineering, 16(12): 1427–1432, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62450.

Minsky:1991:IPO

[Min91] N. H. Minsky. The imposition of protocols over open distributed systems. IEEE Transactions on Software Engineering, 17(2):183–195, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67599.

Mehra:1997:SCS

[MIS97] A. Mehra, A. Indiresan, and K. G. Shin. Structuring communication software for qualityof-service guarantees. IEEE Transactions on Software Engineering, 23(10):616–634, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637145.

Mukhopadhyay:1992:SEM

[MK92a] T. Mukhopadhyay and S. Kekre.

Software effort models for early estimation of process control applications. IEEE Transactions on Software Engineering, 18(10):915–924, October 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=163607.

Munson:1992:DFP

[MK92b] J. C. Munson and T. M. Khoshgoftaar. The detection of fault-prone programs. IEEE Transactions on Software Engineering, 18(5):423–433, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=135775.

Micallef:1993:SAI

[MK93] J. Micallef and G. E. Kaiser. Support algorithms for incremental attribute evaluation of asynchronous subtree replacements. IEEE Transactions on Software Engineering, 19(3):231–252, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221136.

Matsuura:1997:EFP

[MKH97] S. Matsuura, H. Kuruma, and S. Honiden. EVA: a flexible programming method for evolving systems. IEEE Transactions on Software Engineering, 23(5):296–313, May 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588522.

May:1995:MCS

[ML95] J. H. R. May and A. D. Lunn. A model of code sharing for estimating software failure on demand probabilities. IEEE Transactions on Software Engineering, 21(9):747–753, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=464546.

Markowitz:1990:IEE

[MM90] V. M. Markowitz and J. A. Makowsky. Identifying extended entity-relationship object structures in relational schemas. IEEE Transactions on Software Engineering, 16(8):777–790, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57618.

Mili:1995:RSI

[MMM95] H. Mili, F. Mili, and A. Mili.

Reusing software: issues and research directions. IEEE Transactions on Software Engineering, 21(6):528–562, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=391379.

Mili:1997:SRS

[MMM97a] R. Mili, A. Mili, and R. T. Mittermeir. Storing and retrieving software components: a refinement based system. IEEE Transactions on Software Engineering, 23(7):445–460, July 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=605762.

Myers:1997:AEN

[MMM^{+}97b] B. A. Myers, R. G. Mc-Daniel, R. C. Miller, A. S. Ferrency, A. Faulring, B. D. Kyle, A. Mickish, A. Klimovitski, and P. Doane. The Amulet environment: new models for effective user interface software development. IEEE Transactions on Software Engineering, 23(6):347–365, June 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=601073.

Miller:1992:EPF

[MMN^{+}92] K. W. Miller, L. J. Morell, R. E. Noonan, S. K. Park, D. M. Nicol, B. W. Murrill, and M. Voas. Estimating the probability of failure when testing reveals no failures. IEEE Transactions on Software Engineering, 18(1):33–43, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=120314.

McNamee:1990:CCR

[MO90] C. M. McNamee and R. A. Olsson. Comments on “Critical races in Ada programs”. IEEE Transactions on Software Engineering, 16(12):1439–??, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=62452.

See [KSB89].

Moore:1990:SVD

[Moo90] A. P. Moore. The specification and verified decomposition of system requirements using CSP. IEEE Transactions on Software Engineering, 16(9):932–948, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58782.

Morell:1990:TFB

[Mor90a] L. J. Morell. A theory of fault-based testing. IEEE Transactions on Software Engineering, 16(8):844–857, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57623.

Mortazavi:1990:PMR

[Mor90b] B. Mortazavi. Performance of MAP in the remote operation of a CNC. IEEE Transactions on Software Engineering, 16(2):231–237, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44386.

Moser:1990:DDG

[Mos90] L. E. Moser. Data dependency graphs for Ada programs. IEEE Transactions on Software Engineering, 16(5):498–509, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=52773.

Moss:1992:WPO

[Mos92] J. E. B. Moss. Working with persistent objects: to swizzle or not to swizzle. IEEE Transactions on Software Engineering, 18(8):657–673, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=153378.

Moriconi:1995:CAR

[MQR95] M. Moriconi, X. Qian, and R. A. Riemenschneider. Correct architecture refinement. IEEE Transactions on Software Engineering, 21(4):356–372, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=385972.

McCann:1998:CPA

[MR98] P. J. McCann and G. C. Roman. Compositional programming abstractions for mobile computing. IEEE Transactions on Software Engineering, 24(2):97–110, February 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=666824.

McLellan:1998:EUW

[MRF^{+}98] S. McLellan, A. Roesler, Z. Fei, S. Chandran, and C. Spinuzzi. Experience using Webbased shotgun measures for large-system characterization and improvement. IEEE Transactions on Software Engineering, 24(4):268–277, April 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=677184.

Mishra:1991:EAS

[MRT91] S. K. Mishra, V. V. Raghavan, and N. F. Tzeng. Efficient algorithms for selection of recovery points in tree task models. IEEE Transactions on Software Engineering, 17(7):731–734, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83909.

Madhavji:1991:PMP

[MS91] N. H. Madhavji and W. Schafer.

Prism-methodology and processoriented environment. IEEE Transactions on Software Engineering, 17(12):1270–1283, December 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=106987.

Mesrobian:1992:SES

[MS92] E. Mesrobian and J. Skrzypek.

A software environment for studying computational neural systems. IEEE Transactions on Software Engineering, 18(7):575–589, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=148476.

Mahajan:1997:VTC

[MS97a] R. Mahajan and B. Shneiderman. Visual and textual consistency checking tools for graphical user interfaces. IEEE Transactions on Software Engineering, 23(11):722–735, November 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=637386.

Mukherjee:1997:MSD

[MS97b] A. Mukherjee and D. P. Siewiorek. Measuring software dependability by robustness benchmarking. IEEE Transactions on Software Engineering, 23(6):366–378, June 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=601075.

Myrtveit:1999:CEA

[MS99] I. Myrtveit and E. Stensrud. A controlled experiment to assess the benefits of estimating with analogy and regression models. IEEE Transactions on Software Engineering, 25(4):510–525, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799947.

Mainkar:1996:SCE

[MT96] V. Mainkar and K. S. Trivedi.

Sufficient conditions for existence of a fixed point in stochastic reward net-based iterative models. IEEE Transactions on Software Engineering, 22(9):640–653, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541435.

Mark:1992:CBS

[MTMS92] W. Mark, S. Tyler, J. McGuire, and J. Schlossberg. Commitmentbased software development. IEEE Transactions on Software Engineering, 18(10):870– 885, October 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
163604.

Muhanna:1991:CPH

[Muh91]  W. A. Muhanna.  Com-

posite programs: hierarchical construction, circularity, and deadlocks. IEEE Transactions on Software Engineering, 17(4):320–333, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90432.

Mullin:1990:OSD

[Mul90] J. K. Mullin. Optimal semijoins for distributed database systems. IEEE Transactions on Software Engineering, 16(5):558–560, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=52778.

Mutka:1992:ECS

[Mut92] M. W. Mutka. Estimating capacity for sharing in a privately owned workstation environment. IEEE Transactions on Software Engineering, 18(4):319–328, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=129220.

Maxwell:1996:SDP

[MVD96] K. D. Maxwell, L. Van Wassenhove, and S. Dutta. Software development productivity of European space, military, and industrial applications. IEEE Transactions on Software Engineering, 22(10):706–718, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=544349.

Malaiya:1993:EFE

[MvMS93] Y. K. Malaiya, A. von Mayrhauser, and P. K. Srimani. An examination of fault exposure ratio. IEEE Transactions on Software Engineering, 19(11):1087–1094, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256855.

Moriconi:1990:ARA

[MW90] M. Moriconi and T. C. Winkler. Approximate reasoning about the semantic effects of program changes. IEEE Transactions on Software Engineering, 16(9):980–992, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58785.

Manna:1992:FDP

[MW92] Z. Manna and R. Waldinger.

Fundamentals of deductive program synthesis. IEEE Transactions on Software Engineering, 18(8):674–704, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153379.

Murphy:1999:EES

[MWB99] G. C. Murphy, R. J. Walker, and E. L. A. Banlassad. Evaluating emerging software development technologies: lessons learned from assessing aspectoriented programming. IEEE Transactions on Software Engineering, 25(4):438–455, July/ August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=799936.

Menasce:1994:UFE

[MYK94] D. A. Menasce, Y. Yesha, and K. Kalpakis. On a unified framework for the evaluation of distributed quorum attainment protocols. IEEE Transactions on Software Engineering, 20(11):868–884, November 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368122.

Nakajo:1993:CHD

[NAT93] T. Nakajo, I. Azuma, and M. Tada. A case history development of a foolproofing interface documentation system. IEEE Transactions on Software Engineering, 19(8):765–773, August 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238580.

Ng:1995:ITS

[NBK95] F. Ng, G. Butler, and Judy Kay. An intelligent tutoring system for the Dijkstra– Gries methodology. IEEE Transactions on Software Engineering, 21(5):415–428, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=387471.

Narayana:1990:FSL

[ND90] K. T. Narayana and S. Dharap.

Formal specification of a look manager. IEEE Transactions on Software Engineering, 16(9):1089–1103, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58792.

Ness:1993:TCE

[Nes93] L. Ness. L.0: a truly concurrent executable temporal logic language for protocols. IEEE Transactions on Software Engineering, 19(4):410–423, April 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=223807.

Nicola:1990:MCF

[NG90] V. F. Nicola and A. Goyal.

Modeling of correlated failures and community error recovery in multiversion software. IEEE Transactions on Software Engineering, 16(3):350–359, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48942.

Novak:1992:NIS

[NHWS92] G. S. Novak, F. N. Hill, M. L. Wan, and B. G. Sayrs. Negotiated interfaces for software reuse. IEEE Transactions on Software Engineering, 18(7):646–653, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=148482.

Nicholl:1990:USM

[Nic90]  R. A. Nicholl.  Unreach-

able states in model-oriented specifications. IEEE Transactions on Software Engineering, 16(4):472–477, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=54300.

Niemi:1991:PBM

[NJ91] T. Niemi and K. Jarvelin. Prolog-based meta-rules for relational database representation and manipulation. IEEE Transactions on Software Engineering, 17(8):762–788, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83913.

Notkin:1996:GEI

[NJ96] D. Notkin and D. R. Jeffery. Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 22(6):361–??, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=508310.

Nakajo:1991:CHA

[NK91] T. Nakajo and H. Kume. A case history analysis of software error cause-effect relationships. IEEE Transactions on Software Engineering, 17(8):830–838, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83917.

Nuseibeh:1994:FER

[NKF94] B. Nuseibeh, J. Kramer, and A. Finkelstein. A framework for expressing the relationships between multiple views in requirements specification. IEEE Transactions on Software Engineering, 20(10):760–773, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=328995.

Notomi:1994:HRG

[NM94] M. Notomi and T. Murata.

Hierarchical reachability graph of bounded Petri nets for concurrent-software analysis. IEEE Transactions on Software Engineering, 20(5):325–336, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=286423.

Neil:1998:LUZ

[NOTS98] M. Neil, G. Ostrolenk, M. Tobin, and M. Southworth. Lessons from using Z to specify a software tool. IEEE Transactions on Software Engineering, 24(1):15–23, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=663995.

Novak:1995:CUM

[Nov95a] G. S. Novak. Conversion of units of measurement. IEEE Transactions on Software Engineering, 21(8):651–661, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403789.

Novak:1995:CVR

[Nov95b] G. S. Novak. Creation of views for reuse of software with different data representations. IEEE Transactions on Software Engineering, 21(12):993–1005, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489074.

Novak:1997:SRS

[Nov97] G. S. Novak. Software reuse by specialization of generic procedures through views. IEEE Transactions on Software Engineering, 23(7):401–417, July 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=605759.

Nota:1992:QES

[NP92] G. Nota and G. Pacini. Querying of executable software specifications. IEEE Transactions on Software Engineering, 18(8):705–716, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153380.

Nakata:1991:PSP

[NS91] I. Nakata and M. Sassa. Programming with streams in a Pascal-like language. IEEE Transactions on Software Engineering, 17(1):1–9, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67573.

Nicollin:1992:CRT

[NSY92] X. Nicollin, J. Sifakis, and S. Yovine. Compiling realtime specifications into extended automata. IEEE Transactions on Software Engineering, 18(9):794–804, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=159837.

Nishida:1991:SAP

[NTFT91] F. Nishida, S. Takamatsu, Y. Fujita, and T. Tani. Semiautomatic program construction from specifications using library modules. IEEE Transactions on Software Engineering, 17(9):853–871, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92909.

Nicola:1990:CAD

[NvS90] V. F. Nicola and J. M. van Spanje. Comparative analysis of different models of checkpointing and recovery. IEEE Transactions on Software Engineering, 16(8):807–821, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=57620.

Neilson:1995:SBC

[NWPM95] J. E. Neilson, C. M. Woodside, D. C. Petriu, and S. Majumdar. Software bottlenecking in client-server systems and rendezvous networks. IEEE Transactions on Software Engineering, 21(9):776–782, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=464543.

Neufeld:1990:DIA

[NY90] G. N. Neufeld and Y. Yang. The design and implementation of an ASN.1-C compiler. IEEE Transactions on Software Engineering, 16(10):1209–1220, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60300.

Ohlsson:1996:PFP

[OA96] N. Ohlsson and H. Alberg.

Predicting fault-prone software modules in telephone switches. IEEE Transactions on Software Engineering, 22(12):886–894, December 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553637.

Oivo:1992:RSE

[OB92] M. Oivo and V. R. Basili. Representing software engineering models: the TAME goal oriented approach. IEEE Transactions on Software Engineering, 18(10):886–898, October 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=163605.

Ortalo:1999:EQE

[ODK99] R. Ortalo, Y. Deswarte, and M. Kaaniche. Experimenting with quantitative evaluation tools for monitoring operational security. IEEE Transactions on Software Engineering, 25(5): 633–650, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=815323.

Offutt:1994:EEW

[OL94] A. J. Offutt and S. D. Lee. An empirical evaluation of weak mutation. IEEE Transactions on Software Engineering, 20(5):337–344, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=286422.

Olender:1990:CSC

[OO90] K. M. Olender and L. J. Osterweil. Cecil: a sequencing constraint language for automatic static analysis generation. IEEE Transactions on Software Engineering, 16(3):268–280, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48935.

Owre:1995:FVF

[ORSvH95] S. Owre, J. Rushby, N. Shankar, and F. von Henke. Formal verification for fault-tolerant architectures: prolegomena to the design of PVS. IEEE Transactions on Software Engineering, 21(2):107–125, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345827.

Osterbye:1995:LSP

[Ost95] K. Osterbye. Literate Smalltalk programming using hypertext. IEEE Transactions on Software Engineering, 21 (2):138–145, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345829.

Plateau:1991:SAN

[PA91] B. Plateau and K. Atif. Stochastic automata network of modeling parallel systems. IEEE Transactions on Software Engineering, 17(10):1093–1108, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99196.

Parnas:1993:PLS

[Par93] D. L. Parnas. Predicate logic for software engineering. IEEE Transactions on Software Engineering, 19(9):856–862, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241769.

Podgurski:1990:FMP

[PC90] A. Podgurski and L. A. Clarke. A formal model of program dependences and its implications for software testing, debugging, and maintenance. IEEE Transactions on Software Engineering, 16(9):965–979, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58784.

Papelis:1992:SAP

[PC92] Y. E. Papelis and T. L. Casavant. Specification and analysis of parallel/distributed software and systems by Petri nets with transition enabling functions. IEEE Transactions on Software Engineering, 18(3):252–261, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126774.

Poels:1997:CPB

[PDB^{+}97] G. Poels, G. Dedene, L. C. Briand, S. Morasca, and V. R. Basili. Comments on “Propertybased software engineering measurement: refining the additivity properties”. IEEE Transactions on Software Engineering, 23(3):190–197, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585508.

See [BMB96, BMB97].

Pekergin:1992:PCO

[Pek92]  M. F. Pekergin.  Parallel

computing optimization in the Apollo Domain network. IEEE Transactions on Software Engineering, 18(4):296–303, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=129218.

Pitt:1990:DCT

[PF90] D. H. Pitt and D. E. Freestone. The derivation of conformance tests from LOTOS specifications. IEEE Transactions on Software Engineering, 16(12):1337–1343, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62442.

Pike:2006:NIA

[Pik06]  L. Pike.  A note on in-

consistent axioms in Rushby’s “Systematic formal verification for fault-tolerant timetriggered algorithms”. IEEE Transactions on Software Engineering, 32(5):347–348, May 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1642681.

Purtilo:1991:EDF

[PJ91] J. M. Purtilo and P. Jalote.

An environment for developing fault-tolerant software. IEEE Transactions on Software Engineering, 17(2):153–159, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67596.

Porter:1997:ASR

[PJ97] A. A. Porter and P. M. Johnson. Assessing software review meetings: results of a comparative analysis of two experimental studies. IEEE Transactions on Software Engineering, 23(3):129–145, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585501.

Perry:1991:MSD

[PK91] D. E. Perry and G. E. Kaiser. Models of software development environments. IEEE Transactions on Software Engineering, 17(3):283–295, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=75417.

Pickard:1999:CEA

[PKJH99] L. Pickard, B. Kitchenham, P. Jones, and Qing Hu. Comments on: “Evaluating alternative software production functions”. IEEE Transactions on Software Engineering, 25(2):282–285, March/April 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=761451.

See [Hu97].

Pande:1994:IDU

[PLR94] H. D. Pande, W. A. Landi, and B. G. Ryder. Interprocedural def-use associations for C systems with single level pointers. IEEE Transactions on Software Engineering, 20(5):385–403, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=286418.

Paulo:1999:HES

[PMdO99] F. Borges Paulo, P. C. Masiero, and M. C. Ferreira de Oliveira. Hypercharts: extended statecharts to support hypermedia specification. IEEE Transactions on Software Engineering, 25(1):33–49, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=748917.

Parnas:1994:PDW

[PMI94] D. Lorge Parnas, J. Madey, and M. Iglewski. Precise documentation of well-structured programs. IEEE Transactions on Software Engineering, 20(12):948–976, December 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=368133.

Pillai:1997:MSD

[PN97] K. Pillai and V. S. Sukumaran Nair. A model for software development effort and cost estimation. IEEE Transactions on Software Engineering, 23(8):485–497, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=624305.

Paul:1994:FSC

[PP94] S. Paul and A. Prakash. A framework for source code search using program patterns. IEEE Transactions on Software Engineering, 20(6):463–475, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=295894.

See comments [Dev95].

Paul:1996:QAP

[PP96] S. Paul and A. Prakash. A query algebra for program databases. IEEE Transactions on Software Engineering, 22(3):202–217, March 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489080.

Peters:1998:UTO

[PP98] D. K. Peters and D. L. Parnas. Using test oracles generated from program documentation. IEEE Transactions on Software Engineering, 24(3):161–173, March 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=667877.

Peng:1993:OSC

[PS93] D. T. Peng and K. G. Shin. Optimal scheduling of cooperative tasks in a distributed system using an enumerative method. IEEE Transactions on Software Engineering, 19(3):253–267, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221134.

Peng:1997:ASC

[PSA97] Dar-Tzen Peng, K. G. Shin, and T. F. Abdelzaher. Assignment and scheduling communicating periodic tasks in distributed real-time systems. IEEE Transactions on Software Engineering, 23(12):745–758, December 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637388.

Protsko:1991:TAG

[PSTS91] L. B. Protsko, P. G. Sorenson, J. P. Tremblay, and D. A. Schaefer. Towards the automatic generation of software diagrams. IEEE Transactions on Software Engineering, 17(1):10–21, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67575.

Porter:1997:EAC

[PSTV97] A. A. Porter, H. P. Siy, C. A. Toman, and L. G. Votta. An experiment to assess the cost-benefits of code inspections in large scale software development. IEEE Transactions on Software Engineering, 23(6):329–346, June 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=601071.

Prechelt:1998:CEA

[PT98] L. Prechelt and W. F. Tichy. A controlled experiment to assess the benefits of procedure argument type checking. IEEE Transactions on Software Engineering, 24(4):302–312, April 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=677186.

Pucci:1992:NAM

[Puc92] G. Pucci. A new approach to the modeling of recovery block structures. IEEE Transactions on Software Engineering, 18(2):159–167, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=121757.

Porter:1995:CDM

[PVB95] A. A. Porter, L. G. Votta, and V. R. Basili. Comparing detection methods for software requirements inspections: a replicated experiment. IEEE Transactions on Software Engineering, 21(6):563–575, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=391380.

Paulson:1992:AAI

[PW92] D. Paulson and Y. Wand. An automated approach to information systems decomposition. IEEE Transactions on Software Engineering, 18(3):174–189, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126767.

Plaice:1993:NAV

[PW93] J. Plaice and W. W. Wadge. A new approach to version control. IEEE Transactions on Software Engineering, 19(3):268–276, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221137.

Parrish:1991:ARS

[PZ91] A. Parrish and S. H. Zweben. Analysis and refinement of software test data adequacy properties. IEEE Transactions on Software Engineering, 17(6):565–581, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=87282.

Parrish:1993:CSF

[PZ93] A. S. Parrish and S. H. Zweben. Clarifying some fundamental concepts in software testing. IEEE Transactions on Software Engineering, 19(7):742–746, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238573.

Parrish:1995:RAA

[PZ95] A. S. Parrish and S. H. Zweben. On the relationships among the all-uses, all-DU-paths, and alledges testing criteria. IEEE Transactions on Software Engineering, 21(12):1006–1009, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489075.

Puketza:1996:MTI

[PZC^{+}96] N. J. Puketza, K. Zhang, M. Chung, B. Mukherjee, and R. A. Olsson. A methodology for testing intrusion detection systems. IEEE Transactions on Software Engineering, 22(10):719–729, October 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=544350.

Qadah:1991:EAI

[QHK91] G. Z. Qadah, L. J. Henschen, and J. J. Kim. Efficient algorithms for the instantiated transitive closure queries. IEEE Transactions on Software Engineering, 17(3):296–309, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=75418.

Qin:1993:MRT

[Qin93] Bin Qin. Meet real-time requirements of parallel programs and maximally utilize system resources. IEEE Transactions on Software Engineering, 19(10):976–981, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=245739.

Qureshi:1996:AGS

[QSvMG96] M. A. Qureshi, W. H. Sanders, A. P. A. van Moorsel, and R. German. Algorithms for the generation of state-level representations of stochastic activity networks with general reward structures. IEEE Transactions on Software Engineering, 22(9):603–614, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=541432.

Rajlich:1992:GEI

[Raj92] V. Rajlich. Guest Editor’s introduction: special issue on software maintenance. IEEE Transactions on Software Engineering, 18(12):1037, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1263032.

Rangan:1993:TRP

[Ran93] P. V. Rangan. Trust requirements and performance of a fast subtransport-level protocol for secure communication. IEEE Transactions on Software Engineering, 19(2):181–186, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=214834.

Roman:1990:MPM

[RC90] G. C. Roman and H. C. Cunningham. Mixed programming metaphors in a shared dataspace model of concurrency. IEEE Transactions on Software Engineering, 16(12):1361–1373, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62445.

Ripoll:1997:OAS

[RCGF97] I. Ripoll, A. Crespo, and A. Garcia-Fornes. An optimal algorithm for scheduling soft aperiodic tasks in dynamicpriority preemptive systems. IEEE Transactions on Software Engineering, 23(6):388–400, June 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=601081.

Ramesh:1992:SSD

[RD92] B. Ramesh and V. Dhar. Supporting systems development by capturing deliberations during requirements engineering. IEEE Transactions on Software Engineering, 18(6):498–510, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142872.

Reiter:1996:SGM

[Rei96] M. K. Reiter. A secure group membership protocol. IEEE Transactions on Software Engineering, 22(1):31–42, January 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=481515.

Rich:1992:SLK

[RF92] C. Rich and Y. A. Feldman. Seven layers of knowledge representation and reasoning in support of software development. IEEE Transactions on Software Engineering, 18(6):451–469, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=142869.

Roman:1998:ISI

[RG98] G. C. Roman and C. Ghezzi. Introduction to the special issue. IEEE Transactions on Software Engineering, 24(5):313–314, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=685255.

Roma:1993:FDR

[RGB93] G. C. Roma, R. F. Gamble, and W. E. Ball. Formal derivation of rule-based programs. IEEE Transactions on Software Engineering, 19(3):277–296, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=221138.

Roberts:1998:FII

[RGFR98] T. L. Roberts, M. L. Gibson, K. T. Fields, and R. K. Rainer. Factors that impact implementing a system development methodology. IEEE Transactions on Software Engineering, 24(8):640–649, August 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=707699.

See comments [YSW^{+}01].

Roberts:2001:RCF

[RGRF01] T. L. Roberts, M. L. Gibson, R. K. Rainer, and K. T. Fields. Response to “Comments on factors that impact the implementation of a systems development methodology”. IEEE Transactions on Software Engineering, 27(3):282–286, March 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=910864.

See [RGFR98, YSW^{+}01].

Radivojevic:1991:EDA

[RH91] I. P. Radivojevic and J. Herath.

Executing DSP applications in a fine-grained dataflow environment. IEEE Transactions on Software Engineering, 17(10):1028–1041, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99191.

Rothermel:1996:ART

[RH96] G. Rothermel and M. J. Harrold. Analyzing regression test selection techniques. IEEE Transactions on Software Engineering, 22(8):529–551, August 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=536955.

Rothermel:1998:ESS

[RH98] G. Rothermel and M. J. Harrold. Empirical studies of a safe regression test selection technique. IEEE Transactions on Software Engineering, 24(6):401–419, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=689399.

Rangarajan:1993:CVS

[RJT93] S. Rangarajan, P. Jalote, and S. K. Tripathi. Capacity of voting systems. IEEE Transactions on Software Engineering, 19(7):698–706, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238570.

Rask:1993:SCA

[RLL93] R. Rask, P. Laamanen, and K. Lyyttinen. Simulation and comparison of Albrecht’s function point and DeMarco’s function bang metrics in a CASE environment. IEEE Transactions on Software Engineering, 19(7):661–671, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238567.

Ryder:1990:PID

[RLP90] B. G. Ryder, W. Landi, and H. D. Pande. Profiling an incremental data flow analysis algorithm. IEEE Transactions on Software Engineering, 16(2):129–140, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44377.

Richter:1993:TRI

[RM93] G. Richter and B. Maffeo. Toward a rigorous interpretation of ESML-extended systems modeling language. IEEE Transactions on Software Engineering, 19(2):165–180, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214833.

Roussopoulos:1991:AHP

[RMSF91] N. Roussopoulos, L. Mark, T. Sellis, and C. Faloutsos. An architecture for high performance engineering information systems. IEEE Transactions on Software Engineering, 17(1):22–33, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67576.

Rommen:1991:PLB

[Rom91] C. G. Rommen. The probability of load balancing success in a homogeneous network. IEEE Transactions on Software Engineering, 17(9):922–933, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92912.

Rosenblum:1995:PAP

[Ros95] D. S. Rosenblum. A practical approach to programming with assertions. IEEE Transactions on Software Engineering, 21(1):19–31, January 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=341844.

Rushby:1998:SSP

[ROS98] J. Rushby, S. Owre, and N. Shankar. Subtypes for specifications: predicate subtyping in PVS. IEEE Transactions on Software Engineering, 24(9):709–720, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=713327.

Rao:1993:AFI

[RP93] H. C. Rao and L. L. Peterson. Accessing files in an Internet: the Jade file system. IEEE Transactions on Software Engineering, 19(6):613–624, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232026.

Robinson:1999:MRI

[RP99] W. N. Robinson and S. D.

Pawlowski. Managing requirements inconsistency with development goal monitors. IEEE Transactions on Software Engineering, 25(6):816– 835, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824411.

Ravn:1993:SVR

[RRH93] A. P. Ravn, H. Rischel, and K. M. Hansen. Specifying and verifying requirements of real-time systems. IEEE Transactions on Software Engineering, 19(1):41–55, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=210306.

Ramanathan:1993:UCT

[RS93] P. Ramanathan and K. G. Shin.

Use of common time base for checkpointing and rollback recovery in a distributed system. IEEE Transactions on Software Engineering, 19(6):571–583, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=232022.

Rice:1994:FMM

[RS94] M. D. Rice and S. B. Seidman. A formal model for module interconnection languages. IEEE Transactions on Software Engineering, 20(1):88–101, January 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=263757.

Rolia:1995:ML

[RS95] J. A. Rolia and K. C. Sevcik. The method of layers. IEEE Transactions on Software Engineering, 21(8):689–700, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403785.

Rajlich:1996:ERO

[RS96] V. Rajlich and J. H. Silva. Evolution and reuse of orthogonal architecture. IEEE Transactions on Software Engineering, 22(2):153–157, February 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=485224.

Rolland:1998:GGM

[RSA98] C. Rolland, C. Souveyet, and C. B. Achour. Guiding goal modeling using scenarios. IEEE Transactions on Software Engineering, 24(12):1055–1071, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738339.

Ryu:1990:PAD

[RT90] I. K. Ryu and A. Thomasian.

Performance analysis of dynamic locking with the nowaiting policy. IEEE Transactions on Software Engineering, 16(7):684–698, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=56095.

Richardson:1993:ATD

[RT93] D. J. Richardson and M. C.

Thompson. An analysis of test data selection criteria using the RELAY model of fault detection. IEEE Transactions on Software Engineering, 19(6):533–553, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232020.

Ramamoorthy:1990:ESE

[RUPT90] C. V. Ramamoorthy, Y. Usuda, A. Prakash, and W. T. Tsai. The evolution support environment system. IEEE Transactions on Software Engineering, 16(11):1225–1234, November 1990. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=60311.

Rushby:1999:SFV

[Rus99] J. Rushby. Systematic formal verification for faulttolerant time-triggered algorithms. IEEE Transactions on Software Engineering, 25(5): 651–660, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=815324.

See note [Pik06].

Rushby:1993:FVA

[RvH93] J. M. Rushby and F. von Henke. Formal verification of algorithms for critical systems. IEEE Transactions on Software Engineering, 19(1):13–23, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=210304.

Reubenstein:1991:RAA

[RW91] H. B. Reubenstein and R. C. Waters. The Requirements Apprentice: automated assistance for requirements acquisition. IEEE Transactions on Software Engineering, 17(3):226–240, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=75413.

Roman:1994:ADR

[RW94] G. C. Roman and C. D.

Wilcox. Architecture-directed refinement. IEEE Transactions on Software Engineering, 20(4):239–258, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277573.

Rosenblum:1997:UCI

[RW97] D. S. Rosenblum and E. J.

Weyuker. Using coverage information to predict the costeffectiveness of regression testing strategies. IEEE Transactions on Software Engineering, 23(3):146–156, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585502.

Ramming:1999:GEI

[RW99] J. C. Ramming and D. S. Wile.

Guest editorial — introduction to the special section. IEEE Transactions on Software Engineering, 25(3):289–290, May/ June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798319.

Rubin:1990:EEV

[RWG90] R. V. Rubin, J. Walker, and E. J. Golin. Early experience with the Visual Programmer’s WorkBench. IEEE Transactions on Software Engineering, 16(10):1107–1121, October 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60292.

Sahinoglu:1992:CPS

[Sah92]  M. Sahinoglu.  Compound-

Poisson software reliability model. IEEE Transactions on Software Engineering, 18 (7):624–630, July 1992. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
148480.

Sarper:1995:NSS

[Sar95] H. Sarper. No special schemes are needed for solving software reliability optimization models. IEEE Transactions on Software Engineering, 21(8):701–702, August 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=403793.

Satyanarayanan:1992:ISD

[Sat92] M. Satyanarayanan. The influence of scale on distributed file system design. IEEE Transactions on Software Engineering, 18(1):1–9, January 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=120311.

Selby:1991:AEP

[SB91] R. W. Selby and V. R. Basili.

Analyzing error-prone system structure. IEEE Transactions on Software Engineering, 17(2):141–152, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67595.

Seaman:1998:COE

[SB98] C. B. Seaman and V. R. Basili.

Communication and organization: an empirical study of discussion in inspection meetings. IEEE Transactions on Software Engineering, 24(7):559–572, July 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=708569.

Simpson:1999:SEE

[SB99] D. J. Simpson and F. W. Burton. Space efficient execution of deterministic parallel programs. IEEE Transactions on Software Engineering, 25(6): 870–882, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824415.

Stocks:1996:FSB

[SC96] P. Stocks and D. Carrington.

A framework for specificationbased testing. IEEE Transactions on Software Engineering, 22(11):777–793, November 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553698.

Stiemerling:1998:UCS

[SC98] O. Stiemerling and A. B. Cremers. The use of cooperation scenarios in the design and evaluation of a CSCW system. IEEE Transactions on Software Engineering, 24(12):1171–1181, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738345.

Schneidewind:1992:MVS

[Sch92] N. F. Schneidewind. Methodology for validating software metrics. IEEE Transactions on Software Engineering, 18(5):410–422, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=135774.

Schneidewind:1993:SRM

[Sch93]  N. F. Schneidewind.  Soft-

ware reliability model with optimal selection of failure data. IEEE Transactions on Software Engineering, 19(11): 1095–1104, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=256856.

Schneider:1998:VAP

[Sch98] S. Schneider. Verifying authentication protocols in CSP. IEEE Transactions on Software Engineering, 24(9):741–758, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=713329.

Schneidewind:1999:MEM

[Sch99] N. F. Schneidewind. Measuring and evaluating maintenance process using reliability, risk, and test metrics. IEEE Transactions on Software Engineering, 25(6): 769–781, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824387.

Sarkar:1990:ESP

[SD90] D. Sarkar and N. Deo. Estimating the speedup in parallel parsing. IEEE Transactions on Software Engineering, 16(7):677–683, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=56094.

Snoeck:1998:EDK

[SD98] M. Snoeck and G. Dedene. Existence dependency: The key to semantic integrity between structural and behavioral aspects of object types. IEEE Transactions on Software Engineering, 24(4):233–251, April 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=677182.

Shaw:1995:ASA

[SDK^{+}95] Mary Shaw, Robert DeLine, Daniel V. Klein, Theodore L. Ross, David M. Young, and Gregory Zelesnik. Abstractions for software architecture and tools to support them. IEEE Transactions on Software Engineering, 21(4):314–335, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=385970.

Sears:1993:LAM

[Sea93] A. Sears. Layout appropriateness: a metric for evaluating user interface widget layout. IEEE Transactions on Software Engineering, 19(7):707–719, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238571.

Seaman:1999:QME

[Sea99] C. B. Seaman. Qualitative methods in empirical studies of software engineering. IEEE Transactions on Software Engineering, 25(4):557–572, July/ August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799955.

Sen:1997:ROS

[Sen97] A. Sen. The role of opportunism in the software design reuse process. IEEE Transactions on Software Engineering, 23(7):418–436, July 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=605760.

Sereno:1996:AMV

[Ser96] M. Sereno. Approximate mean value analysis for stochastic marked graphs. IEEE Transactions on Software Engineering, 22(9):654–664, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541436.

Srinivasan:1995:MLA

[SF95] K. Srinivasan and D. Fisher.

Machine learning approaches to estimating software development effort. IEEE Transactions on Software Engineering, 21(2):126–137, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=345828.

Stamos:1990:IRE

[SG90] J. W. Stamos and D. K. Gifford. Implementing remote evaluation. IEEE Transactions on Software Engineering, 16(7):710–722, July 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=56097.

Shepard:1991:PRT

[SG91] T. Shepard and J. A. M. Gagne. A pre-run-time scheduling algorithm for hard realtime systems. IEEE Transactions on Software Engineering, 17(7):669–677, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83903.

See comment [AS97].

Spezialetti:1995:LMS

[SG95] M. Spezialetti and R. Gupta. Loop monotonic statements. IEEE Transactions on Software Engineering, 21(6):497–505, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=391376.

Shieh:1990:MHD

[SGCT90] Y. B. Shieh, D. Ghosal, P. R. Chintamaneni, and S. K. Tripathi. Modeling of hierarchical distributed systems with fault-tolerance. IEEE Transactions on Software Engineering, 16(4):444–457, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54296.

Sun:1997:BCT

[SGL97] J. Sun, M. K. Gardner, and J. W. S. Liu. Bounding completion times of jobs with arbitrary release times, variable execution times and resource sharing. IEEE Transactions on Software Engineering, 23(10):603–615, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637144.

Shaw:1992:CRT

[Sha92] A. C. Shaw. Communicating real-time state machines. IEEE Transactions on Software Engineering, 18(9):805–816, September 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=159840.

Stoyenko:1991:AHR

[SHH91] A. D. Stoyenko, V. C. Hamacher, and R. C. Holt. Analyzing hardreal-time programs for guaranteed schedulability. IEEE Transactions on Software Engineering, 17(8):737–750, August 1991. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83911.

Shilling:1993:ILP

[Shi93] J. J. Shilling. Incremental LL(1) parsing in languagebased editors. IEEE Transactions on Software Engineering, 19(9):935–940, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=241775.

Shin:1991:FRH

[SI91] D. G. Shin and K. B. Irani. Fragmenting relations horizontally using a knowledge-based approach. IEEE Transactions on Software Engineering, 17(9):872–883, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92906.

Singhal:1990:UTN

[Sin90] M. Singhal. Update transport: a new technique for update synchronization in replicated database systems. IEEE Transactions on Software Engineering, 16(12):1325–1336, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=62441.

Singpurwalla:1991:DOT

[Sin91] N. D. Singpurwalla. Determining an optimal time interval for testing and debugging software. IEEE Transactions on Software Engineering, 17(4):313–319, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90431.

Sztrik:1991:AAH

[SK91] J. Sztrik and D. Kouvatsos.

Asymptotic analysis of a heterogeneous multiprocessor system in a randomly changing environment. IEEE Transactions on Software Engineering, 17(10):1069–1075, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99194.

Sullivan:1996:EMM

[SKN96] K. J. Sullivan, I. J. Kalet, and D. Notkin. Evaluating the mediator method: Prism as a case study. IEEE Transactions on Software Engineering, 22(8):563–579, August 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=536957.

Shimeall:1991:ECS

[SL91] T. J. Shimeall and N. G. Leveson. An empirical comparison of software fault tolerance and fault elimination. IEEE Transactions on Software Engineering, 17(2):173–182, February 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67598.

Stafylopatis:1992:PIR

[SL92] A. Stafylopatis and A. Likas. Pictorial information retrieval using the random neural network. IEEE Transactions on Software Engineering, 18(7): 590–600, July 1992. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
148477.

Shih:1993:MRM

[SLL93] Wei Kuan Shih, J. W. S. Liu, and C. L. Liu. Modified rate-monotonic algorithm for scheduling periodic jobs with deferred deadlines. IEEE Transactions on Software Engineering, 19(12):1171–1179, December 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=249662.

Steegmans:1992:GIP

[SLvH92] E. Steegmans, J. Lewi, and I. van Horebeek. Generation of interactive parsers with error handling. IEEE Transactions on Software Engineering, 18(5):357–367, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=135769.

Saxena:1994:LCA

[SM94] N. R. Saxena and E. J. Mc-Cluskey. Linear complexity assertions for sorting. IEEE Transactions on Software Engineering, 20(6):424–431, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=295891.

Sutcliffe:1998:DTR

[SM98] A. Sutcliffe and N. Maiden.

The domain theory for requirements engineering. IEEE Transactions on Software Engineering, 24(3):174–196, March 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=667878.

Smith:1990:KSP

[Smi90] D. R. Smith. KIDS: a semiautomatic program development system. IEEE Transactions on Software Engineering, 16(9):1024–1043, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58788.

Smith:1991:MSR

[Smi91]  G. W. Smith.  Modeling

security-relevant data semantics. IEEE Transactions on Software Engineering, 17(11): 1195–1203, November 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=106974.

Smith:1992:FNN

[Smi92] L. S. Smith. A framework for neural net specification. IEEE Transactions on Software Engineering, 18(7):601–612, July 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=148478.

Sutcliffe:1998:SSB

[SMMM98] A. G. Sutcliffe, N. A. M. Maiden, S. Minocha, and D. Manuel. Supporting scenariobased requirements engineering. IEEE Transactions on Software Engineering, 24(12): 1072–1088, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=738340.

Sullivan:1999:ACB

[SMS99] K. J. Sullivan, M. Marchukov, and J. Socha. Analysis of a conflict between aggregation and interface negotiation in Microsoft’s Component Object Model. IEEE Transactions on Software Engineering, 25(4):584–599, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799960.

Stoyen:1999:DEC

[SMYP99] A. D. Stoyen, T. J. Marlowe, M. F. Younis, and P. V. Petrov. A development environment for complex distributed real-time applications. IEEE Transactions on Software Engineering, 25(1):50–74, January/February 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=748918.

Sorensen:1993:CMM

[SNH93] E. V. Sorensen, J. Nordahl, and N. H. Hansen. From CSP models to Markov models. IEEE Transactions on Software Engineering, 19(6):554–570, June 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=232021.

Song:1994:EAC

[SO94] Xiping Song and L. J. Osterweil. Experience with an approach to comparing software design methodologies. IEEE Transactions on Software Engineering, 20(5):364–384, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=286419.

Song:1998:ESD

[SO98] Xiping Song and L. J. Osterweil. Engineering software design processes to guide process execution. IEEE Transactions on Software Engineering, 24(9):759–775, September 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=713330.

Son:1997:GEI

[Son97] Sang H. Son. Guest editorial: Introduction to the special section. IEEE Transactions on Software Engineering, 23(10):601–602, October 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=637143.

Sibert:1996:AIS

[SPL96] O. Sibert, P. A. Porras, and R. Lindell. An analysis of the Intel 80x86 security architecture and implementations. IEEE Transactions on Software Engineering, 22(5):283–293, May 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=502221.

Seiter:1998:EOB

[SPL98] L. M. Seiter, J. Palsberg, and K. J. Lieberherr. Evolution of object behavior using context relations. IEEE Transactions on Software Engineering, 24(1):79–92, January 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=663999.

Setliff:1992:KRR

[SR92] D. E. Setliff and R. A. Rutenbar. Knowledge representation and reasoning in a software synthesis architecture. IEEE Transactions on Software Engineering, 18(6):523–533, June 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=142874.

Solheim:1993:EST

[SR93] J. A. Solheim and J. H. Rowland. An empirical study of testing and integration strategies using artificial software systems. IEEE Transactions on Software Engineering, 19(10):941–949, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=245736.

Sowmya:1998:EST

[SR98] A. Sowmya and S. Ramesh. Extending statecharts with temporal logic. IEEE Transactions on Software Engineering, 24(3):216–231, March 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=667880.

Siff:1999:IMC

[SR99] M. Siff and T. Reps. Identifying modules via concept analysis. IEEE Transactions on Software Engineering, 25(6): 749–768, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=824377.

Sellis:1991:QON

[SS91] T. K. Sellis and L. Shapiro. Query optimization for nontraditional database applications. IEEE Transactions on Software Engineering, 17(1):77–86, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67580.

Saavedra:1995:PCO

[SS95a] R. H. Saavedra and A. J. Smith. Performance characterization of optimizing compilers. IEEE Transactions on Software Engineering, 21(7):615–628, July 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=392982.

Sriram:1995:MPL

[SS95b] M. G. Sriram and M. Singhal. Measures of the potential for load sharing in distributed computing systems. IEEE Transactions on Software Engineering, 21(5):468–475, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=387476.

Shepperd:1997:ESP

[SS97] M. Shepperd and C. Schofield.

Estimating software project effort using analogies. IEEE Transactions on Software Engineering, 23(11):736–743, November 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637387.

Suzuki:1990:PMV

[SSM90] T. Suzuki, S. M. Shatz, and T. Murata. A protocol modeling and verification approach based on a specification language and Petri nets. IEEE Transactions on Software Engineering, 16(5):523–536, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=52775.

Sommerville:1999:MPI

[SSV99] I. Sommerville, P. Sawyer, and S. Viller. Managing process inconsistency using viewpoints. IEEE Transactions on Software Engineering, 25(6): 784–799, November/December 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=824395.

Staskauskas:1993:FDC

[Sta93]  M. G. Staskauskas.  For-

mal derivation of concurrent programs: an example from industry. IEEE Transactions on Software Engineering, 19(5):503–528, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=232015.

Stavely:1995:VDI

[Sta95] A. M. Stavely. Verifying definite iteration over data structures. IEEE Transactions on Software Engineering, 21(6):506–514, June 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=391377.

Stoller:1994:APR

[Sto94] S. Stoller. Addendum to “Proof rules for flush channels”. IEEE Transactions on Software Engineering, 20(8):664–??, August 1994. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=310675.

See [CKA93].

Suzuki:1990:FAA

[Suz90] I. Suzuki. Formal analysis of the alternating bit protocol by temporal Petri nets. IEEE Transactions on Software Engineering, 16(11):1273–1281, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=60315.

Sinha:1992:CFE

[SV92] A. P. Sinha and I. Vessey. Cognitive fit: an empirical study of recursion and iteration. IEEE Transactions on Software Engineering, 18(5):368–379, May 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=135770.

Stewart:1997:DDR

[SVK97] D. B. Stewart, R. A. Volpe, and P. K. Khosla. Design of dynamically reconfigurable real-time software using portbased objects. IEEE Transactions on Software Engineering, 23(12):759–776, December 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637390.

Smith:1993:SPE

[SW93] C. U. Smith and L. G. Williams. Software performance engineering: a case study including performance comparison with design alternatives. IEEE Transactions on Software Engineering, 19(7):720–741, July 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=238572.

Sitaraman:1997:PNA

[SWO97] M. Sitaraman, B. W. Weide, and W. F. Ogden. On the practical need for abstraction relations to verify abstract data type representations. IEEE Transactions on Software Engineering, 23(3):157–170, March 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=585503.

Strom:1993:ETC

[SY93] R. E. Strom and D. M. Yellin. Extending typestate checking using conditional liveness analysis. IEEE Transactions on Software Engineering, 19(5):478–485, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=232013.

Schwan:1992:DSH

[SZ92] K. Schwan and H. Zhou.

Dynamic scheduling of hard real-time tasks and real-time threads. IEEE Transactions on Software Engineering, 18(8):736–748, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153383.

Tai:1996:TFB

[Tai96] Kuo-Chung Tai. Theory of fault-based predicate testing for computer programs. IEEE Transactions on Software Engineering, 22(8):552–562, August 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=536956.

Taylor:1992:LFC

[Tay92] H. Taylor. A lingua franca for concurrent logic programming. IEEE Transactions on Software Engineering, 18(3):225–236, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126771.

Taff:1991:EDE

[TBH91] L. M. Taff, J. W. Borchering, and W. R. Hudgins. Estimeetings: development estimates and a front-end process for a large project. IEEE Transactions on Software Engineering, 17(8):839–849, August 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83918.

Tyrrell:1995:CMI

[TC95] A. M. Tyrrell and G. F. Carpenter. CSP methods for identifying atomic actions in the design of fault tolerant concurrent systems. IEEE Transactions on Software Engineering, 21(7):629–639, July 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=392983.

Tai:1991:DCA

[TCO91] K. C. Tai, R. H. Carver, and E. E. Obaid. Debugging concurrent Ada programs by deterministic execution. IEEE Transactions on Software Engineering, 17(1):45–63, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67578.

Tsoukalas:1993:SRE

[TDN93] M. Z. Tsoukalas, J. W. Duran, and S. C. Ntafos. On some reliability estimation problems in random and partition testing. IEEE Transactions on Software Engineering, 19(7):687–697, July 1993. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=238569.

Tsai:1990:NMR

[TFCB90] J. J. P. Tsai, K. Y. Fang, H. Y.

Chen, and Y. D. Bi. A noninterference monitoring and replay mechanism for real-time software testing and debugging. IEEE Transactions on Software Engineering, 16(8):897–916, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57626.

Tsai:1990:ICS

[TGC90] C. R. Tsai, V. D. Gligor, and C. S. Chandersekaran. On the identification of covert storage channels in secure systems. IEEE Transactions on Software Engineering, 16(6):569–580, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=55086.

Tian:1995:ITD

[Tia95] J. Tian. Integrating time domain and input domain analyses of software reliability using tree-based models. IEEE Transactions on Software Engineering, 21(12):945–958, December 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=489071.

Taylor:1992:STC

[TLK92] R. N. Taylor, D. L. Levine, and C. D. Kelly. Structural testing of concurrent programs. IEEE Transactions on Software Engineering, 18(3):206–215, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=126769.

Tian:1995:TEB

[TLP95] J. Tian, Peng Lu, and J. Palma. Test-execution-based reliability measurement and modeling for large commercial software. IEEE Transactions on Software Engineering, 21(5):405–414, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=387470.

Tsang:1998:LDA

[TM98] S. Tsang and E. H. Magill. Learning to detect and avoid run-time feature interactions in intelligent networks. IEEE Transactions on Software Engineering, 24(10):818–830, October 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=729682.

Taylor:1996:CMB

[TMA^{+}96] R. N. Taylor, N. Medvidovic, K. M. Anderson, E. J. Whitehead, J. E. Robbins, K. A. Nies, P. Oreizy, and D. L. Dubrow. A component- and messagebased architectural style for GUI software. IEEE Transactions on Software Engineering, 22(6):390–406, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=508313.

Thibault:1999:DSL

[TMC99] S. A. Thibault, R. Marlet, and C. Consel. Domain-specific languages: from design to implementation application to video device drivers generation. IEEE Transactions on Software Engineering, 25(3):363–377, May/ June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=798325.

Torii:1999:GEC

[TMN^{+}99] K. Torii, K. Matsumoto, K. Nakakoji, Y. Takada, S. Takada, and K. Shima. Ginger2: an environment for computeraided empirical software engineering. IEEE Transactions on Software Engineering, 25(4):474–492, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=799942.

Tomek:1993:MCS

[TMT93] L. A. Tomek, J. K. Muppala, and K. S. Trivedi. Modeling correlation in software recovery blocks. IEEE Transactions on Software Engineering, 19(11):1071–1086, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=256854.

Triantafillou:1997:ASC

[TN97] P. Triantafillou and C. Neilson. Achieving strong consistency in a distributed file system. IEEE Transactions on Software Engineering, 23(1):35–55, January 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=581328.

Towsley:1986:APC

[Tow86] D. Towsley. Allocating programs containing branches and loops within a multiple processor system. IEEE Transactions on Software Engineering, SE-12(10):1018–1024, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313018.

See comments [Kau90].

Towsley:1990:CAP

[Tow90]  D. Towsley.  Correction to

“Allocating programs containing branches and loops within a multiple processor system”. IEEE Transactions on Software Engineering, 16(4):472–??, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54299.

Thomasian:1991:PAT

[TR91] A. Thomasian and I. K. Ryu.

Performance analysis of twophase locking. IEEE Transactions on Software Engineering, 17(5):386–402, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90443.

Triantafillou:1996:IRL

[Tri96]  P. Triantafillou.  Indepen-

dent recovery in large-scale distributed systems. IEEE Transactions on Software Engineering, 22(11):812–826, November 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=553700.

Triantafillou:1995:LBP

[TT95] P. Triantafillou and D. J. Taylor.

The location-based paradigm for replication: Achieving efficiency and availability in distributed systems. IEEE Transactions on Software Engineering, 21(1):1–18, January 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=341843.

Tucci:1994:PNL

[TVC94] M. Tucci, G. Vitiello, and G. Costagliola. Parsing nonlinear languages. IEEE Transactions on Software Engineering, 20(9):720–739, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=317427.

Tsai:1990:ATC

[TVK90] W. T. Tsai, D. Volovik, and T. F. Keefe. Automated test case generation for programs specified by relational algebra queries. IEEE Transactions on Software Engineering, 16(3):316–324, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48939.

Tsai:1992:HKR

[TWJ92] J. J. P. Tsai, T. Weigert, and H. C. Jang. A hybrid knowledge representation as a basis of requirement specification and specification analysis. IEEE Transactions on Software Engineering, 18(12):1076–1100, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=184762.

Tsai:1995:TCP

[TYC95] J. J. P. Tsai, S. Jennhwa Yang, and Yao-Hsiung Chang. Timing constraint Petri nets and their application to schedulability analysis of real-time system specifications. IEEE Transactions on Software Engineering, 21(1):32–49, January 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=341845.

Tohma:1991:EPH

[TYOJ91] Y. Tohma, H. Yamano, M. Ohba, and R. Jacoby. The estimation of parameters of the hypergeometric distribution and its application to the software reliability growth model. IEEE Transactions on Software Engineering, 17(5):483–489, May 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90450.

Tian:1995:CME

[TZ95] J. Tian and M. V. Zelkowitz.

Complexity measure evaluation and selection. IEEE Transactions on Software Engineering, 21(8):641–650, August 1995. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403788.

Ukelson:1993:UNC

[UGB93] J. P. Ukelson, J. D. Gould, and S. J. Boies. User navigation in computer applications. IEEE Transactions on Software Engineering, 19(3):297–306, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=221139.

Varvarigou:1999:RMA

[VAA99] T. A. Varvarigou, M. E. Anagnostou, and S. R. Ahuja. Reconfiguration models and algorithms for stateful interactive processes. IEEE Transactions on Software Engineering, 25(3):401–415, May/June 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=798328.

Venkatesan:1995:TDD

[VD95] S. Venkatesan and B. Dathan. Testing and debugging distributed programs using global predicates. IEEE Transactions on Software Engineering, 21(2):163–177, February 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=345831.

Vestal:1994:FPS

[Ves94] S. Vestal. Fixed-priority sensitivity analysis for linear compute time models. IEEE Transactions on Software Engineering, 20(4):308–317, April 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=277577.

Venkatesh:1992:SDE

[VF92] G. A. Venkatesh and C. N.

Fischer. SPARE: a development environment for program analysis algorithms. IEEE Transactions on Software Engineering, 18(4):304–318, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=129219.

vanGenuchten:1991:WSL

[vG91] M. van Genuchten. Why is software late? An empirical study of reasons for delay in software development. IEEE Transactions on Software Engineering, 17(6):582–590, June 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=87283.

vanLamsweerde:1998:MCG

[vLDL98] A. van Lamsweerde, R. Darimont, and E. Letier. Managing conflicts in goal-driven requirements engineering. IEEE Transactions on Software Engineering, 24(11):908–926, November 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=730542.

vanLamsweerde:1998:IDR

[vLW98] A. van Lamsweerde and L. Willemet. Inferring declarative requirements specifications from operational scenarios. IEEE Transactions on Software Engineering, 24(12):1089–1114, December 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=738341.

Valenzano:1990:SPT

[VMC90] A. Valenzano, P. Montuschi, and L. Ciminiera. Some properties of timed token medium access protocols. IEEE Transactions on Software Engineering, 16(8):858–869, August 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=57628.

vonMayrhauser:1999:CSS

[vML99] A. von Mayrhauser and S. Lang. A coding scheme to support systematic analysis of software comprehension. IEEE Transactions on Software Engineering, 25(4):526–540, July/August 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=799950.

Voas:1992:PDF

[Voa92] J. M. Voas. PIE: a dynamic failure-based technique. IEEE Transactions on Software Engineering, 18(8):717–727, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153381.

Vose:1991:LAG

[Vos91] Michael D. Vose. A linear algorithm for generating random numbers with a given distribution. IEEE Transactions on Software Engineering, 17(9):972–975, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=92917.

Vranes:1995:IMP

[VS95] S. Vranes and M. Stanojevic. Integrating multiple paradigms within the blackboard framework. IEEE Transactions on Software Engineering, 21(3):244–262, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372151.

Verner:1992:SSM

[VT92] J. Verner and G. Tate. A software size model. IEEE Transactions on Software Engineering, 18(4):265–278, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=129216.

VanderWiel:1993:ASD

[VV93] S. A. Vander Wiel and L. G. Votta. Assessing software designs using capture-recapture methods. IEEE Transactions on Software Engineering, 19(11):1045–1054, November 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=256852.

VonMayrhauser:1996:IDC

[VV96] A. Von Mayrhauser and A. M. Vans. Identification of dynamic comprehension processes during large scale maintenance. IEEE Transactions on Software Engineering, 22(6):424–437, June 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=508315.

Wang:1996:TII

[Wan96] Lin Lin Wang. Thorough investigation into “An improved algorithm based on subset closures for synthesizing a relational database scheme”. IEEE Transactions on Software Engineering, 22(4):271–274, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=491651.

See [YLN88].

Ward:1996:DDI

[War96]  M. Ward.  Derivation of

data intensive algorithms by formal transformation: the Schnorr–Waite graph marking algorithm. IEEE Transactions on Software Engineering, 22(9):665–686, September 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=541437.

Wellings:1997:IAA

[WB97] A. Wellings and A. Burns.

Implementing atomic actions in Ada 95. IEEE Transactions on Software Engineering, 23(2):107–123, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=585500.

Welling:1998:AEE

[WB98] G. Welling and B. R. Badrinath.

An architecture for exporting environment awareness to mobile computing applications. IEEE Transactions on Software Engineering, 24(5):391–400, May 1998. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=685262.

White:1990:AAS

[WC90] S. R. White and L. Comerford. ABYSS: an architecture for software protection. IEEE Transactions on Software Engineering, 16(6):619–629, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=55090.

Weide:1994:DSI

[WEHL94] B. W. Weide, S. H. Edwards, D. E. Harms, and D. A. Lamb. Design and specification of iterators using the swapping paradigm. IEEE Transactions on Software Engineering, 20(8):631–643, August 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=310672.

Weyuker:1990:CDF

[Wey90] E. J. Weyuker. The cost of data flow testing: an empirical study. IEEE Transactions on Software Engineering, 16(2):121–128, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=44376.

Weyuker:1993:MED

[Wey93] E. J. Weyuker. More experience with data flow testing. IEEE Transactions on Software Engineering, 19(9):912–919, September 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=241773.

Wu:1993:RTU

[WF93] K. L. Wu and W. K. Fuchs.

Rapid transaction-undo recovery using twin-page storage management. IEEE Transactions on Software Engineering, 19(2):155–164, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214832.

Wermelinger:1998:CMP

[WF98] M. Wermelinger and J. L. Fiadeiro. Connectors for mobile programs. IEEE Transactions on Software Engineering, 24(5):331–341, May 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=685257.

Weyuker:1994:AGT

[WGS94] E. Weyuker, T. Goradia, and A. Singh. Automatically generating test data from a Boolean specification. IEEE Transactions on Software Engineering, 20(5):353–363, May 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=286420.

Wilde:1992:MSO

[WH92] N. Wilde and R. Huitt. Maintenance support for objectoriented programs. IEEE Transactions on Software Engineering, 18(12):1038–1044, December 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1263033.

Weerahandi:1994:SQM

[WH94] S. Weerahandi and R. E. Hausman. Software quality measurement based on faultdetection data. IEEE Transactions on Software Engineering, 20(9):665–676, September 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=317425.

Waldspurger:1992:SDC

[WHH^{+}92] C. A. Waldspurger, T. Hogg, B. A. Huberman, J. O. Kephart, and W. S. Stornetta. Spawn: a distributed computational economy. IEEE Transactions on Software Engineering, 18(2):103–117, February 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=121753.

Wing:1990:ULS

[Win90] J. M. Wing. Using Larch to specify Avalon/C++ objects. IEEE Transactions on Software Engineering, 16(9): 1076–1088, September 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58791.

Weyuker:1991:APT

[WJ91] E. J. Weyuker and B. Jeng.

Analyzing partition testing strategies. IEEE Transactions on Software Engineering, 17(7):703–711, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=83906.

Willson:1990:PNT

[WK90] R. G. Willson and B. H.

Krogh. Petri net tools for the specification and analysis of discrete controllers. IEEE Transactions on Software Engineering, 16(1):39–50, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44362.

Walter:1997:FVL

[WLS97] C. J. Walter, P. Lincoln, and N. Suri. Formally verified online diagnosis. IEEE Transactions on Software Engineering, 23(11):684–721, November 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=637385.

Wang:1995:RTB

[WM95] Rwo-Hsi Wang and A. K. Mok. Response-time bounds of EQL rule-based programs under rule priority structure. IEEE Transactions on Software Engineering, 21(7):605–614, July 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=392981.

Wang:1997:DVC

[WM97] Changqing Wang and D. R. Musser. Dynamic verification of C++ generic algorithms. IEEE Transactions on Software Engineering, 23(5):314–323, May 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588523.

Whang:1992:TDS

[WMS^{+}92] K. Y. Whang, A. Malhotra, G. H. Sockut, L. Burns, and K. S. Choi. Two-dimensional specification of universal quantification in a graphical database query language. IEEE Transactions on Software Engineering, 18(3):216–224, March 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=126770.

Woodside:1995:TVM

[Woo95] C. M. Woodside. A threeview model for performance engineering of concurrent software. IEEE Transactions on Software Engineering, 21 (9):754–767, September 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=464545.

Wang:1994:SBS

[WP94] Yabo Wang and D. L. Parnas. Simulating the behavior of software modules by trace rewriting. IEEE Transactions on Software Engineering, 20(10):750–759, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=328996.

Wohlin:1994:CSC

[WR94a] C. Wohlin and P. Runeson. Certification of software components. IEEE Transactions on Software Engineering, 20(6):494–499, June 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=295896.

Wohlwend:1994:SSI

[WR94b] H. Wohlwend and S. Rosenbaum. Schlumberger’s software improvement program. IEEE Transactions on Software Engineering, 20(11):833– 839, November 1994. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        368125.

Wilcox:1996:RAP

[WR96] C. D. Wilcox and G. C. Roman. Reasoning about places, times, and actions in the presence of mobility. IEEE Transactions on Software Engineering, 22(4):225–247, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=491647.

Waheed:1998:MED

[WRH98] A. Waheed, D. T. Rover, and J. K. Hollingsworth. Modeling and evaluating design alternatives for an on-line instrumentation system: a case study. IEEE Transactions on Software Engineering, 24(6):451–470, June 1998. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=689402.

Woodbury:1990:MAW

[WS90] M. H. Woodbury and K. G. Shin. Measurement and analysis of workload effects on fault latency in real-time systems. IEEE Transactions on Software Engineering, 16 (2):212–216, February 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44383.

Wang:1992:CAA

[WS92] D. Wang and B. Schurmann.

Computer aided analysis and derivation for artificial neural systems. IEEE Transactions on Software Engineering, 18(8):728–735, August 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=153382.

Wu:1997:PSS

[WS97] M. Y. Wu and W. Shu. On parallelization of static scheduling algorithms. IEEE Transactions on Software Engineering, 23(8):517–528, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=624307.

Wolfson:1991:MCN

[WSY91] O. Wolfson, S. Sengupta, and Y. Yemini. Managing communication networks by monitoring databases. IEEE Transactions on Software Engineering, 17(9):944–953, September 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=92914.

Whittaker:1994:MCM

[WT94] J. A. Whittaker and M. G. Thomason. A Markov chain model for statistical software testing. IEEE Transactions on Software Engineering, 20 (10):812–824, October 1994. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=328991.

Wand:1990:OMI

[WW90] Y. Wand and R. Weber. An ontological model of an information system. IEEE Transactions on Software Engineering, 16(11):1282–1292, November 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=60316.

Wojcik:1991:RGE

[WW91] Z. M. Wojcik and B. E. Wojcik. Rough grammar for efficient and fault-tolerant computing on a distributed system. IEEE Transactions on Software Engineering, 17(7):652–668, July 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=83902.

Xu:1990:SPR

[XP90] J. Xu and D. L. Parnas.

Scheduling processes with release times, deadlines, precedence and exclusion relations. IEEE Transactions on Software Engineering, 16(3):360–369, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=48943.

Xu:1993:CST

[XP93a] J. Xu and D. L. Parnas. Correction to ‘On satisfying timing constraints in hard-real-time systems’ by J. Xu and D. L. Parnas. IEEE Transactions on Software Engineering, 19(3):310–??, March 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=221141.

Xu:1993:STC

[XP93b] J. Xu and D. L. Parnas. On satisfying timing constraints in hard-real-time systems. IEEE Transactions on Software Engineering, 19(1):70–84, January 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=210308.

Xu:1993:MSP

[Xu93] J. Xu. Multiprocessor scheduling of processes with release times, deadlines, precedence, and exclusion relations. IEEE Transactions on Software Engineering, 19(2):139–154, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214831.

Yadav:1990:CDM

[Yad90] S. B. Yadav. Control and definition modularization: an improved software design technique for organizing programs. IEEE Transactions on Software Engineering, 16(1):92–99, January 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=44367.

Yu:1992:WCR

[YCHL92] P. S. Yu, M. S. Chen, H. U. Heiss, and S. Lee. On workload characterization of relational database environments. IEEE Transactions on Software Engineering, 18(4):347–355, April 1992. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=129222.

Yu:1992:AHC

[YD92] P. S. Yu and D. M. Dias. Analysis of hybrid concurrency control schemes for a high data contention environment. IEEE Transactions on Software Engineering, 18(2):118–129, February 1992. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=121754.

Yu:1993:PAC

[YD93] P. S. Yu and D. M. Dias.

Performance analysis of concurrency control using locking with deferred blocking. IEEE Transactions on Software Engineering, 19(10):982–996, October 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=245740.

Yasumatsu:1995:SST

[YD95] K. Yasumatsu and N. Doi.

SPiCE: a system for translating Smalltalk programs into a C environment. IEEE Transactions on Software Engineering, 21(11):902–912, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=473219.

Yu:1990:SVM

[YG90] C. F. Yu and V. D. Gligor.

A specification and verification method for preventing denial of service. IEEE Transactions on Software Engineering, 16(6):581–592, June 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=55087.

Yang:1988:IAB

[YLN88] C. C. Yang, G. Li, and P. A. B. Ng. An improved algorithm based on subset closures for synthesizing a relational database scheme. IEEE Transactions on Software Engineering, 14(11):1731–1738, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9058. See

investigation [Wan96].

Young:1993:CTL

[YLT93] M. Young, D. L. Levine, and R. N. Taylor. Comments on “Temporal logic-based deadlock analysis for Ada” by G. M. Karam and R. J. A. Burh. IEEE Transactions on Software Engineering, 19(2):198–199, February 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=214836.

See [KB91].

Younis:1999:SSS

[YMST99] M. F. Younis, T. J. Marlowe, A. D. Stoyen, and G. Tsai. Statically safe speculative execution for real-time systems. IEEE Transactions on Software Engineering, 25(5): 701–721, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=815328.

Young:1997:CVS

[You97] W. D. Young. Comparing verification systems: interactive consistency in ACL2. IEEE Transactions on Software Engineering, 23(4):214–223, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=588536.

Yazia-Pekergin:1991:SBE

[YPV91] N. Yazia-Pekergin and J. M.

Vincent. Stochastic bounds on execution times of parallel programs. IEEE Transactions on Software Engineering, 17(10):1005–1012, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=99189.

Yen:1993:HFE

[YS93] M. Y. M. Yen and R. W.

Scamell. A human factors experimental comparison of SQL and QBE. IEEE Transactions on Software Engineering, 19(4):390–409, April 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=223806.

Yadav:2001:CFI

[YSW^{+}01] S. B. Yadav, N. G. Shaw, L. Webb, C. Sutcu, T. L. Roberts, M. L. Gibson, R. K.

Rainer, and K. T. Fields. Comments on “Factors that impact implementing a system development methodology” [with reply]. IEEE Transactions on Software Engineering, 27(3):279–286, March 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=910863.

See [RGFR98] and response [RGRF01].

Yan:1997:SSM

[YZM97] Yong Yan, Xiaodong Zhang, and Qian Ma. Software support for multiprocessor latency measurement and evaluation. IEEE Transactions on Software Engineering, 23(1):4–16, January 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=581326.

Zahedi:1991:SRA

[ZA91] F. Zahedi and N. Ashrafi. Software reliability allocation based on structure, utility, price, and cost. IEEE Transactions on Software Engineering, 17(4):345–356, April 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=90434.

Zave:1991:IEP

[Zav91] P. Zave. An insider’s evaluation of PAISLey. IEEE Transactions on Software Engineering, 17(3):212–225, March 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=75412.

Zweben:1995:ELE

[ZEWH95] S. H. Zweben, S. H. Edwards, B. W. Weide, and J. E. Hollingsworth. The effects of layering and encapsulation on software development cost and quality. IEEE Transactions on Software Engineering, 21(3):200–208, March 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=372147.

Zhang:1991:PMM

[Zha91] X. Zhang. Performance measurement and modeling to evaluate various effects on a shared memory multiprocessor. IEEE Transactions on Software Engineering, 17(1):87–93, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=67581.

Zhu:1996:FAS

[Zhu96] Hong Zhu. A formal analysis of the subsume relation between software test adequacy criteria. IEEE Transactions on Software Engineering, 22(4):248–255, April 1996. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=491648.

Zave:1996:WDO

[ZJ96] P. Zave and M. Jackson. Where do operations come from? A multiparadigm specification technique. IEEE Transactions on Software Engineering, 22(7):508–528, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=538607.

Zhang:1991:PPE

[ZQ91] X. Zhang and X. Qin. Performance prediction and evaluation of parallel processing on a NUMA multiprocessor. IEEE Transactions on Software Engineering, 17(10):1059–1068, October 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=99193.

Zuse:1997:RPB

[Zus97] H. Zuse. Reply to: “Propertybased software engineering measurement”. IEEE Transactions on Software Engineering, 23(8):533–??, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=624309.

See [BMB96].